### v1

# On importance sampling and independent Metropolis–Hastings with an unbounded weight function

Abstract
Importance sampling and independent Metropolis–Hastings (IMH) are among the fundamental building blocks of Monte Carlo methods. Both require a proposal distribution that globally approximates the target distribution. The Radon–Nikodym derivative of the target distribution relative to the proposal is called the weight function. Under the weak assumption that the weight is unbounded but has a number of finite moments under the proposal distribution, we obtain new results on the approximation error of importance sampling and of the particle independent Metropolis–Hastings algorithm (PIMH), which includes IMH as a special case. For IMH and PIMH, we show that the common random numbers coupling is maximal. Using that coupling we derive bounds on the total variation distance of a PIMH chain to the target distribution. The bounds are sharp with respect to the number of particles and the number of iterations. Our results allow a formal comparison of the finite-time biases of importance sampling and IMH. We further consider bias removal techniques using couplings of PIMH, and provide conditions under which the resulting unbiased estimators have finite moments. We compare the asymptotic efficiency of regular and unbiased importance sampling estimators as the number of particles goes to infinity.
1 Introduction
1.1 Context and contributions
Monte Carlo with global proposals.
Monte Carlo methods aim to approximate a target distribution $\pi$ on a measurable space $(\mathbb{X},\mathcal{X})$, for example $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$. These techniques are crucial when analytical computation of expectations under $\pi$ is infeasible due to high dimensionality or complex distribution forms. The goal is to evaluate integrals of functions $f:\mathbb{X}\longrightarrow\mathbb{R}$ with respect to $\pi$:
[…]
Two primary approaches are Markov Chain Monte Carlo (MCMC) methods, that construct a Markov chain with $\pi$ as its stationary distribution, and Importance Sampling (IS) methods, where the target distribution is approximated by weighted samples. Among MCMC methods, the independent Metropolis–Hastings (IMH) algorithm is a specialized form of the Metropolis–Rosenbluth–Teller–Hastings (MRTH) algorithm (Metropolis et al., 1953; Hastings, 1970), in which proposals are drawn from a distribution $q$ independently of the current state of the chain. The same proposal $q$ can be employed to generate draws in importance sampling, a procedure that can be traced back to Kahn (1949), as noted in Andral (2022). We assume that we can draw from $q$ and that its density, also denoted by $q$, can be evaluated pointwise. Therefore, IMH and IS propose two ways of correcting for the discrepancy between a proposal $q$ and a target $\pi$, and their comparison is a natural and fundamental question.
Our contributions
concern the performance of importance sampling and independent Metropolis–Hastings, and their comparison. Our key assumption is that the Radon–Nykodym derivative $\omega$ of $\pi$ with respect to $q$, termed the weight, has $p$ finite moments under $q$. We first show that the bias of self-normalized importance sampling is of order $N^{-1}$, and we obtain new bounds on the moments of the error in importance sampling in Section 2. We then consider IMH, and show that the common random numbers coupling is optimal in Section 3. Using this coupling, in Section 4 we show that the total variation distance between IMH at iteration $t$ and $\pi$ decays as $t^{p-1}$. We obtain matching lower bounds in an example. We also obtain explicit dependencies in $N$ for the particle IMH algorithm (Andrieu et al., 2010). In Section 5 we consider the bias removal technique of Glynn & Rhee (2014) applied by Middleton et al. (2019) to the particle IMH algorithm. This yields an unbiased estimator that can be implemented whenever self-normalized importance sampling or IMH can be implemented. We provide conditions under which the estimators have finite moments, and conditions under which their efficiency is asymptotically equivalent to that of importance sampling.
1.2 Importance sampling
Self-normalized importance sampling (IS) is described in Algorithm 1, see also Chapter 9.2 in Owen (2013). Central to importance sampling is the weight function defined as
$\omega:x\mapsto\frac{\pi(x)}{q(x)},\quad\text{so that}\quad q(\omega)=1.$ (2)
Since multiplicative constants in $\omega$ have no effect on the IS estimator (4), it can be computed as long as the user can evaluate a function proportional to $\omega$ in (2). Unless specified otherwise, by IS we will refer to the self-normalized procedure in Algorithm 1; and not to the more basic estimator $N^{-1}\sum_{n=1}^{N}\omega(x_{n})f(x_{n})$ that depends on the multiplicative constant in $\omega$.
  1. 1.
    Sample $N$ particles independently $x_{1},\ldots,x_{N}$ from $q$.
  2. 2.
    Compute the importance weights $\omega(x_{n})=\pi(x_{n})/q(x_{n})$ for $n\in[N]=\{1,\cdots,N\}$.
  3. 3.
    Compute
    $\hat{Z}(x_{1},\ldots,x_{N})=N^{-1}\sum_{n=1}^{N}\omega(x_{n}).$ (3)
  4. 4.
    For any test function $f$, compute the IS estimator
    $\hat{F}(x_{1},\ldots,x_{N})=\frac{\sum_{n=1}^{N}\omega(x_{n})f(x_{n})}{\sum_{n=1}^{N}\omega(x_{n})}.$ (4)
  5. 5.
    Return $\hat{F}(x_{1},\ldots,x_{N})$ and $\hat{Z}(x_{1},\ldots,x_{N})$.
Algorithm 1 Self-normalized importance sampling.
We make the following assumption throughout.
Assumption 1.
For any measurable set $A\in\mathcal{X}$, if $q(A)=0$, then $\pi(A)=0$, in other words $\pi$ is absolutely continuous with respect to $q$. Furthermore, $\omega(x)$ with $x\sim q$ is almost surely positive, and $q(\omega)=1$.
Under Assumption 1, if $\pi(f)$ exists then $\hat{F}(x_{1},\ldots,x_{N})\to\pi(f)$ as $N\to\infty$ almost surely. The asymptotic variance of IS is directly computed from the delta method (Owen, 2013; Robert & Casella, 2004; Liu, 2008), assuming $q(\omega^{2}\cdot f^{2})<\infty$ and $q(\omega^{2})<\infty$,
$\lim_{n\to\infty}\mathbb{V}\left[\sqrt{N}(\hat{F}(x_{1},\ldots,x_{N})-\pi(f))\right]=q(\omega^{2}\cdot(f-\pi(f))^{2}).$ (5)
Agapiou et al. (2017) provide non-asymptotic bounds on the mean squared error and on the bias of importance sampling, which are both inversely proportional to the number $N$ of draws from $q$; see Theorem 2.2. The exact form of the asymptotic bias of IS is well-known (e.g. Skare et al., 2003; Liu, 2008), and we provide a formal statement in Section 2.
1.3 Independent Metropolis–Hastings
Independent Metropolis-Hastings (IMH) is an instance of the Metropolis–Rosenbluth–Teller–Hastings algorithm, where the draws are proposed from $q$ independently of the current state of the chain (Hastings, 1970, Section 2.5); see Algorithm 2. Thus IMH is implementable in the same settings as importance sampling. Under Assumption 1, the IMH chain is $\pi$-irreducible, on top of being aperiodic and $\pi$-invariant by design, thus for $\pi$-almost every $x$, $|P^{t}(x,\cdot)-\pi|_{\mathrm{TV}}\to 0$ as $t\to\infty$ (Theorem 4 in Roberts & Rosenthal, 2004), where $P$ denotes the transition kernel of IMH, $P^{t}$ denotes the $t$-steps transition kernel, and $|\mu-\nu|_{\text{TV}}=\sup_{A\in\mathcal{X}}\mu(A)-\nu(A)$.
The asymptotic variance of the ergodic average $t^{-1}\sum_{s=0}^{t-1}f(x_{t})$ generated by IMH is finite if $\pi(f^{2})<\infty$ and $q(\omega^{2}\cdot f^{2})$ (Theorem 2 in Deligiannidis & Lee, 2018), and under these conditions its variance is greater than or equal to the asymptotic variance of importance sampling in (5) (Proposition 2 in Deligiannidis & Lee, 2018). Thus, in terms of asymptotic variance, the comparison is clear: IS outperforms IMH. Since IMH defines a Markov transition, it can directly be used as a step within an encompassing Gibbs sampler (Skare et al., 2003), and it is commonly used within sequential Monte Carlo samplers (Chopin, 2002; South et al., 2019), and thus has its specific uses irrespective of the performance comparison with importance sampling.
  1. 1.
    Draw $x^{\star}\sim q$.
  2. 2.
    Compute the acceptance probability:
    $\alpha_{\text{RH}}(x,x^{\star})=\min\left\{1,\frac{\omega(x^{\star})}{\omega(x)}\right\}.$ (6)
  3. 3.
    Draw $u$ from a $\text{Uniform}(0,1)$ distribution.
  4. 4.
    If $u<\alpha_{\text{RH}}(x,x^{\star})$, set $x^{\prime}=x^{\star}$, otherwise $x^{\prime}=x$.
  5. 5.
    Return $x^{\prime}$.
Algorithm 2 IMH algorithm describing one step starting from $x$.
When it comes to non-asymptotic behavior, for IMH there is an important distinction between two cases (Mengersen & Tweedie, 1996): either the weight is bounded, in which case the chain is geometrically ergodic and exact rates are obtained in Wang (2022), or the weight is unbounded and the convergence is polynomial at best; in the latter case, various results are provided e.g. in Jarner & Roberts (2002); Douc et al. (2007); Roberts & Rosenthal (2011); Andrieu et al. (2022) and Douc et al. (2018, Chapter 17); see Section 4.3. In Section 4 we provide polynomial bounds on the total variation distance to stationarity for IMH under moment conditions on $\omega$ under $q$. Our results enable a comparison of the biases of IS and IMH in Section 4.4, which turns out in favor of IMH.
In the following we consider the particle IMH (PIMH) generalization of IMH, where $N$ proposals are drawn at each iteration (Andrieu et al., 2010, Section 4.2); see Algorithm 3. We define the algorithm on the state space $\mathbb{X}^{N}$, use boldface to denote its elements, e.g. $\mathbf{x}=(x_{1},\ldots,x_{N})\in\mathbb{X}^{N}$, and denote the transition kernel by $P$. If $N=1$ the algorithm corresponds to IMH, and our results apply for all $N\geq 1$. To view Algorithm 3 as a special case of IMH, define for any $N\geq 1$
$\bar{\pi}(x_{1},\ldots,x_{N})$ $=\sum_{k=1}^{N}\frac{\pi(x_{k})}{N}\prod_{n\neq k}q(x_{n})=\left(\frac{1}{N}\sum_{k=1}^{N}\omega(x_{k})\right)\prod_{n=1}^{N}q(x_{n}),$ (7)
$\bar{q}(x_{1},\ldots,x_{N})$ $=\prod_{n=1}^{N}q(x_{n}),$ (8)
and, in the case $N=1$, $\bar{\pi}(x_{1})=\pi(x_{1})$. From (3) and the above definitions, we can write:
$\bar{\omega}(\mathbf{x})=\frac{\bar{\pi}(\mathbf{x})}{\bar{q}(\mathbf{x})}=\frac{1}{N}\sum_{n=1}^{N}\omega(x_{n})=\hat{Z}(\mathbf{x}).$ (9)
Hence, IMH as in Algorithm 2, with proposal $\bar{q}$ and target $\bar{\pi}$, is equivalent to Algorithm 3.
  1. 1.
    Draw $\mathbf{x}^{\star}=(x^{\star}_{1},\ldots,x_{N}^{\star})\sim\bar{q}$.
  2. 2.
    Compute the acceptance probability:
    $\alpha_{\text{RH}}(\mathbf{x},\mathbf{x}^{\star})=\min\left\{1,\frac{\hat{Z}(\mathbf{x}^{\star})}{\hat{Z}(\mathbf{x})}\right\},$ (10)
    where $\hat{Z}:\mathbf{x}\mapsto N^{-1}\sum_{n=1}^{N}\omega(x_{n})$.
  3. 3.
    Draw $u$ from a $\text{Uniform}(0,1)$ distribution.
  4. 4.
    If $u<\alpha_{\text{RH}}(\mathbf{x},\mathbf{x}^{\star})$, set $\mathbf{x}^{\prime}=\mathbf{x}^{\star}$, otherwise $\mathbf{x}^{\prime}=\mathbf{x}$.
  5. 5.
    Return $\mathbf{x}^{\prime}$.
Algorithm 3 PIMH algorithm describing one step starting from $\mathbf{x}=(x_{1},\ldots,x_{N})$.
In order to estimate an expectation $\pi(f)$ from the PIMH output, a Rao–Blackwellization argument explained in Andrieu et al. (2010, Appendix B.5) leads to
$\mathbb{E}_{\mathbf{x}\sim\bar{\pi}}\left[\hat{F}(\mathbf{x})\right]=\mathbb{E}_{\bar{\pi}}\left[\frac{\sum_{n=1}^{N}\omega(x_{n})f(x_{n})}{\sum_{n=1}^{N}\omega(x_{n})}\right]=\pi(f),$ (11)
with $\hat{F}$ as in (4). For completeness we restate the arguments of Andrieu et al. (2010) justifying (11) in Appendix A.1. Thus, we can compute $\hat{F}(\mathbf{x})=\sum_{n=1}^{N}\omega(x_{n})f(x_{n})/\sum_{n=1}^{N}\omega(x_{n})$ at each iteration of the chain $(\mathbf{x}_{t})_{t\geq 0}$, and the ergodic average $T^{-1}\sum_{t=0}^{T-1}\hat{F}(\mathbf{x}_{t})$ may converge to $\pi(f)$. This allows connections between PIMH and IS, the latter being equal to $\hat{F}(\mathbf{x})$ with $\mathbf{x}\sim\bar{q}$.
1.4 Moment conditions on the weight
We introduce the assumption under which most of our results are derived.
Assumption 2.
The weights have a finite $p$-th moment for $p\geq 2$: $q(\omega^{p})<\infty$.
This is a weak and natural assumption in the context of both self-normalized importance sampling and IMH. For bounded test functions Assumption 2 is necessary for the asymptotic variance of both self-normalized importance sampling and IMH to be finite.
Example 1 (Exponential distributions).
Let $\pi$ be the Exponential(1) distribution and let $q$ be the Exponential($k$) distribution with $q(x)=ke^{-kx}$, both on $\mathbb{R}_{+}$. If $k\leq 1$, the weight $\omega(x)$ is upper bounded by $k^{-1}$, and Assumption 2 holds for all $p\geq 2$. If $k>1$, then $q(\omega^{p})<\infty$ holds with any $p<k/(k-1)$, and the requirement $p\geq 2$ translates into $k<2$. The example is considered in Jarner & Roberts (2007); Roberts & Rosenthal (2011); Andrieu et al. (2022).
Example 2 (Normal distributions).
Let $\pi$ be the Normal(0,1) distribution and let $q$ be the Normal$(0,\sigma^{2})$ distribution, both on $\mathbb{R}$. If $\sigma^{2}\geq 1$, the weight $\omega(x)$ is upper bounded by $\sigma$, and Assumption 2 holds for all $p\geq 2$. If $\sigma^{2}<1$, then $q(\omega^{p})<\infty$ holds for $p<\sigma^{-2}/(\sigma^{-2}-1)$. The requirement $p\geq 2$ in Assumption 2 translates into $\sigma^{2}>1/2$. The example is considered in Roberts & Rosenthal (2011); Owen (2013).
Assumption 2 implies the following well-known behavior of the average of $N$ independent weights. The proofs of the results below are in Appendix A.1.
Proposition 1.1.
Let $\mathbf{x}=(x_{n})_{n=1}^{N}$ be $N$ i.i.d. random variables from $q$. Under Assumptions 1-2, with $p\geq 2$, $\hat{Z}(\mathbf{x})=N^{-1}\sum_{n=1}^{N}\omega(x_{n})$ satisfies, for all $N\geq 1$,
$\mathbb{E}_{\bar{q}}[\hat{Z}(\mathbf{x})^{p}]$ $\leq\left(1+\frac{2^{1-1/p}(p-1)(1+q(\omega^{p}))^{1/p}}{\sqrt{N}}\right)^{p},$ (12)
$\mathbb{E}_{\bar{q}}[|\hat{Z}(\mathbf{x})-1|^{p}]$ $\leq\left(\frac{2^{1-1/p}(p-1)(1+q(\omega^{p}))^{1/p}}{\sqrt{N}}\right)^{p}=:M(p)N^{-p/2}.$ (13)
We will repeatedly use the following consequence of Markov’s inequality and Proposition 1.1.
Lemma 1.1.
Under Assumptions 1-2, with $p\geq 2$, there exists a constant $M(p)>0$ such that, for any $z>0$ and $N\geq 1$:
$\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})\geq 1+z\right)\leq\frac{M(p)}{N^{p/2}z^{p}}.$ (14)
Remark 1.1.
Most of our proofs require that $\hat{Z}(\mathbf{x})$ is a non-negative random variable, with $\mathbb{E}_{\bar{q}}[\hat{Z}(\mathbf{x})]=1$, $\mathbb{E}_{\bar{q}}[|\hat{Z}(\mathbf{x})-1|^{p}]\leq M(p)N^{-p/2}$ for some $M(p)$ independent of $N$, but not directly that $\hat{Z}(\mathbf{x})$ is an average of i.i.d. weights. Thus $\hat{Z}(\mathbf{x})$ could for example be the normalizing constant estimator generated by a sequential Monte Carlo sampler. Results on moments of sequential Monte Carlo normalizing constant estimators can be found in e.g. Del Moral (2013, Section 16.5).
2 Bias and moments of importance sampling
Our first contribution is a clean statement on the asymptotic bias of self-normalized importance sampling. Introductory material on importance sampling often makes the point that the basic importance sampling estimator $N^{-1}\sum_{n=1}^{N}f(x_{n})\omega(x_{n})$ is unbiased, but since $\omega$ can only be evaluated up to a multiplicative constant, users may need to resort to the self-normalized estimator in (4), which is biased: $\mathbb{E}[\hat{F}(\mathbf{x})]\neq\pi(f)$. The form of the asymptotic bias is well known, e.g. Section 2.5. in Liu (2008). However, somewhat surprisingly, formal results appear to be lacking. The closest may be Theorem 2 in Skare et al. (2003), but their emphasis is on the pointwise relative error of the density of a particle selected from the IS approximation. Their Remark 1 translates this into a bound on the bias for bounded functions under the assumption of bounded weights. We provide Theorem 2.1, with a proof in Appendix A.2, which gives the leading term in the bias of IS under more general conditions on the weights.
Theorem 2.1.
Let $f$ with $\pi(|f|)<\infty$. Assume that $x_{1},\ldots,x_{n}$ are i.i.d. from $q$, let $\omega:x\mapsto\pi(x)/q(x)$, and let $\hat{F}(\mathbf{x})=\sum_{n=1}^{N}\omega(x_{n})f(x_{n})/\sum_{m=1}^{N}\omega(x_{m})$. Assume that $q(|f|\cdot\omega^{2})<\infty$, $q(|f|\cdot\omega^{3})<\infty$ and $q(\omega^{-\eta})<\infty$ for some $\eta>0$. Then
$\lim_{N\to\infty}N\times\mathbb{E}_{\mathbf{x}\sim\bar{q}}\left[\hat{F}(\mathbf{x})-\pi(f)\right]$ $=-\int\left(f(x)-\pi(f)\right)\omega^{2}(x)q(dx).$ (15)
Theorem 2.1 assumes a finite inverse moment of the weight, and for bounded $f$ the theorem requires $q(\omega^{3})<\infty$. The inverse moment assumption may be removed at the cost of higher positive moments. Agapiou et al. (2017) provide an upper bound on the bias under weaker assumptions, which we restate below.
Theorem 2.2 (Bias part of Theorem 2.1 in Agapiou et al. (2017)).
Suppose that $q(\omega^{2})<\infty$ and that $|f|_{\infty}\leq 1$. Then, for all $N\geq 1$,
$\mathbb{E}_{\mathbf{x}\sim\bar{q}}[\hat{F}(\mathbf{x})-\pi(f)]\leq\frac{12}{N}q(\omega^{2}).$
Theorem 2.3 in Agapiou et al. (2017) provides upper bounds of order $N^{-1}$ also for unbounded test functions, under moment conditions on $f$ and on $f\cdot\omega$. Our Theorem 2.1 establishes that $N^{-1}$ is the exact order of the asymptotic bias as a function of $N$, but requires additional conditions. We next provide a result on the $s$-th moments of the error in importance sampling for unbounded test functions. Theorem 2.3 generalizes the MSE part of Theorem 2.3 in Agapiou et al. (2017) to arbitrary orders $s\geq 2$, and its assumptions are weaker in the case $s=2$, as discussed below. The proof is in Appendix A.2. The bounds are central to the results of Section 5.
Theorem 2.3.
Assume that there exist $p\in[2,\infty)$ and $r\in[2,\infty]$ such that $q(\omega^{p})<\infty$ and $q(|f|^{r})<\infty$, and $q(f^{2}\cdot\omega^{2})<\infty$, then for any $2\leq s\leq pr/(p+r+2)$ and any $N\geq 1$, we have:
$\mathbb{E}_{\bar{q}}\left[\left\lvert\hat{F}(\mathbf{x})-\pi(f)\right\rvert^{s}\right]\leq CN^{-s/2},$
where the constant $C$ depends on $r,p,s,q(|f|^{r}),q(\omega^{p}),q(f^{2}\cdot\omega^{2})$. When $r=\infty$, the statement holds for $f$ such that $|f|_{\infty}<\infty$ and all $s\leq p$.
A few remarks are in order:
  • The condition $s\leq pr/(p+r+2)$ implies $s\leq\min\{p,r\}$.
  • We have $q((f\omega)^{pr/p+r})<\infty$ if $q(\omega^{p})<\infty$ and $q(f^{r})<\infty$. Indeed, when $r<\infty$, $q((f\omega)^{pr/p+r})\leq q(f^{r})^{p/p+r}q(\omega^{p})^{r/p+r}<\infty$. When $r=\infty$, the claim remains correct (by understanding $pr/(p+r)$ as $p$), since $q((f\omega)^{p})\leq\lVert f\rVert_{\infty}^{p}q(\omega^{p})$. This observation leads to two facts: 1) If $pr/(p+r)\geq 2$ (e.g. $p=r=4$ or $p=2,r=\infty$), the assumption $q(f^{2}\cdot\omega^{2})<\infty$ in Theorem 2.3 can be derived from the assumptions $q(\omega^{p})<\infty$ and $q(f^{r})<\infty$. 2) The basic importance sampling estimator $N^{-1}\sum_{n=1}^{N}f(x_{n})\omega(x_{n})$ has a finite $s$-th moment under the same conditions, as it has a finite $pr/(p+r)$-th moment, and $s\leq pr/(p+r+2)\leq pr/(p+r)$.
  • We may be particularly interested in the mean-squared error (MSE) of IS, corresponding to $s=2$. Theorem 2.3 implies that the MSE is of order $1/N$ as long as $2\leq pr/(p+r+2)$. This condition holds, for example, if $\min\{p,r\}\geq 2(1+\sqrt{2})\approx 4.828$, or if $p\geq 3$ and $r\geq 10$, or if $p=2$ and $r=\infty$. The case $s=2$ can be compared to the MSE part of Theorem 2.3 in Agapiou et al. (2017). In our notation, they require $q(|f\cdot\omega|^{2d})<\infty$, $q(\omega^{2e})<\infty$, $q(|f|^{2a})<\infty$, $q(\omega^{2b(1+a^{-1})})<\infty$, for $a,b,d,e>1$ such that $a^{-1}+b^{-1}=1,d^{-1}+e^{-1}=1$. Their assumption implies ours, as can be seen by setting $r=2a$ and $p=2b(1+a^{-1})$, since then
    $\frac{pr}{(p+r+2)}=\frac{4b(a+1)}{2a+2b+2ba^{-1}+2}=\frac{2(a+1)}{a(1-a^{-1})+1+a^{-1}+(1-a^{-1})}=\frac{2(a+1)}{a+1}=2,$
    i.e. our theorem holds with $s=2$ under their assumptions.
3 Optimality of coupling IMH with common draws
With a view toward deriving upper bounds on the total variation distance of IMH to stationarity, we consider the common draws (or common random numbers) coupling of a generic IMH algorithm, described in Algorithm 4, and PIMH is retrieved as a special case. The coupling is very simple and was considered in Liu (1996); Roberts & Rosenthal (2011). The pseudocode describes the transition kernel $\bar{P}((\mathbf{x},\mathbf{y}),\cdot)$ of the coupled chains, and we denote the transition of IMH by $P$. It was remarked around Lemma 1 in Wang et al. (2021) that this coupling is “one-step maximal”, in the sense that the probability $\bar{P}((\mathbf{x},\mathbf{y}),D)$ where $D=\{(\mathbf{x},\mathbf{y}):\mathbf{x}=\mathbf{y}\}$ is maximal over all couplings, and is equal to one minus
$|P(\mathbf{x},\cdot)-P(\mathbf{y},\cdot)|_{\text{TV}}=\int\min\left\{\frac{\hat{Z}(\mathbf{x}^{\star})}{\hat{Z}(\mathbf{x})},\;\frac{\hat{Z}(\mathbf{x}^{\star})}{\hat{Z}(\mathbf{y})},\;1\right\}\bar{q}(\mathrm{d}\mathbf{x}^{\star}).$ (16)
  1. 1.
    Draw $\mathbf{x}^{\star}\sim\bar{q}$.
  2. 2.
    Draw $u$ from a Uniform(0,1) distribution.
  3. 3.
    If $u<{\hat{Z}(\mathbf{x}^{\star})}/{\hat{Z}(\mathbf{x})}$, set $\mathbf{x}^{\prime}=\mathbf{x}^{\star}$, otherwise set $\mathbf{x}^{\prime}=\mathbf{x}$.
  4. 4.
    If $u<{\hat{Z}(\mathbf{x}^{\star})}/{\hat{Z}(\mathbf{y})}$, set $\mathbf{y}^{\prime}=\mathbf{x}^{\star}$, otherwise set $\mathbf{y}^{\prime}=\mathbf{y}$.
  5. 5.
    Return $(\mathbf{x}^{\prime},\mathbf{y}^{\prime})$
Algorithm 4 Common draws coupling of IMH, denoted by $\bar{P}$, for chains currently at $(\mathbf{x},\mathbf{y})$.
Let $(\mathbf{x}_{t},\mathbf{y}_{t})$ be a coupled chain started from $(\mathbf{x},\mathbf{y})$ and evolving according to $\bar{P}$. Denoting the meeting time by
$\tau=\inf\{t\geq 1:\mathbf{x}_{t}=\mathbf{y}_{t}\},$ (17)
the coupling inequality states that, for $t\geq 1$,
$|P^{t}(\mathbf{x},\cdot)-P^{t}(\mathbf{y},\cdot)|_{\text{TV}}\leq\mathbb{P}_{\mathbf{x},\mathbf{y}}(\tau>t),$ (18)
where the probability $\mathbb{P}_{\mathbf{x},\mathbf{y}}$ is under the law of $(\mathbf{x}_{t},\mathbf{y}_{t})$ started from $(\mathbf{x},\mathbf{y})$ at time zero. We will relate the probability $\mathbb{P}_{\mathbf{x},\mathbf{y}}(\tau>t)$ to the rejection probabilities of IMH from $\mathbf{x}$ and $\mathbf{y}$, and we define
$r:\mathbf{x}\mapsto\int_{\mathbf{x}^{\star}\neq\mathbf{x}}\left(1-\alpha_{\text{RH}}(\mathbf{x},\mathbf{x}^{\star})\right)\bar{q}(\mathrm{d}\mathbf{x}^{\star}),$ (19)
where $\alpha_{\text{RH}}(\mathbf{x},\mathbf{x}^{\star})$ is defined in (10).
The meeting time $\tau$ is the first time at which both chains accept the proposal simultaneously, which corresponds to the first time at which the chain with the highest weight accepts the proposal. Indeed, if $\hat{Z}(\mathbf{x})\geq\hat{Z}(\mathbf{y})$, then $\alpha_{\text{RH}}(\mathbf{x},\mathbf{x}^{\star})\leq\alpha_{\text{RH}}(\mathbf{y},\mathbf{x}^{\star})$ for all $\mathbf{x}^{\star}$, and thus $u<\alpha_{\text{RH}}(\mathbf{x},\mathbf{x}^{\star})$ implies that $u<\alpha_{\text{RH}}(\mathbf{y},\mathbf{x}^{\star})$. Thus, conditionally on $\mathbf{x}_{0}=\mathbf{x},\mathbf{y}_{0}=\mathbf{y}$, the meeting time $\tau$ follows a Geometric distribution with parameter $1-r(\mathbf{x})$, where $r(\mathbf{x})$ is defined in (19). Recall that the survival function of a Geometric variable $T$ with parameter $\gamma$ is given by: $\mathbb{P}(T>t)=(1-\gamma)^{k}$ for $t\in\mathbb{N}$. Still assuming $\hat{Z}(\mathbf{x})\geq\hat{Z}(\mathbf{y})$, we obtain, for $t\geq 1$,
$|P^{t}(\mathbf{x},\cdot)-P^{t}(\mathbf{y},\cdot)|_{\text{TV}}\leq\mathbb{P}_{\mathbf{x},\mathbf{y}}(\tau>t)=(r(\mathbf{x}))^{t}.$ (20)
The above upper bound is given in Roberts & Rosenthal (2011). In their remark following Theorem 5, they state that this is also a lower bound without providing a proof. We do so below, for both discrete and continuous state spaces; Roberts & Rosenthal (2011) focus on non-atomic spaces. First, we express
$|P^{t}(\mathbf{x},\cdot)-P^{t}(\mathbf{y},\cdot)|_{\text{TV}}=\sup_{A\in\mathcal{B}(\mathbb{R}^{d})}|P^{t}(\mathbf{x},A)-P^{t}(\mathbf{y},A)|,$ (21)
and we select the set $A=\mathbb{R}^{d}\setminus\{\mathbf{x}\}$ to obtain a lower bound. Then $P^{t}(\mathbf{y},A)$ = 1 since $\mathbf{x}\neq\mathbf{y}$ and assuming that $q(\{\mathbf{x}\})=0$, while $P^{t}(\mathbf{x},A)=1-\left(r(\mathbf{x})\right)^{t}$, i.e. the chain is in $A$ at step $t$ except if $t$ proposals have been rejected. The situation is slightly more complicated if the proposal has non-zero mass on $\{\mathbf{x}\}$ and $\{\mathbf{y}\}$, i.e. in discrete state spaces, but the following result still holds. The proof is in Appendix A.3.
Theorem 3.1.
Let $(\mathbf{x}_{t},\mathbf{y}_{t})$ be a Markov chain evolving according to $\bar{P}$ in Algorithm 4 and starting from $\mathbf{x}_{0}=\mathbf{x}$ and $\mathbf{y}_{0}=\mathbf{y}$. Let $\tau=\inf\{t\geq 1:\mathbf{x}_{t}=\mathbf{y}_{t}\}$, and let $r(\mathbf{x})$ be defined as in (19). Then, for all $t\geq 1$,
$|P^{t}(\mathbf{x},\cdot)-P^{t}(\mathbf{y},\cdot)|_{\text{TV}}=\mathbb{P}_{\mathbf{x},\mathbf{y}}(\tau>t)=\max(r(\mathbf{x}),r(\mathbf{y}))^{t}.$ (22)
Thus, the chain $(\mathbf{x}_{t},\mathbf{y}_{t})$ generated by the common draws coupling follows a maximal coupling, as in Pitman (1976): the coupling inequality is an equality. To the best of our knowledge, this is the only known case of “all time maximal” couplings of an MCMC algorithm. Note also that the upper bound in (22) decreases geometrically in $t$. The polynomial rates come later, when we integrate over $\mathbf{x}$ or $\mathbf{y}$.
4 Meeting times and polynomial convergence
4.1 Meeting times of lagged chains
We consider coupled IMH chains with a lag, as in Middleton et al. (2019). The construction is described in Algorithm 5. The generated chains $(\mathbf{x}_{t})_{t\geq 0}$ and $(\mathbf{y}_{t})_{t\geq 0}$ have the same marginal distribution, that of an IMH chain started from $\bar{q}$.
  1. 1.
    Set $\tau=+\infty$ and $t=1$.
  2. 2.
    Draw $\mathbf{x}_{0}\sim\bar{q}$ and $\mathbf{y}_{0}\sim\bar{q}$ independently.
  3. 3.
    Draw $u$ from a $\text{Uniform}(0,1)$ distribution.
  4. 4.
    If $u<\hat{Z}(\mathbf{y}_{0})/\hat{Z}(\mathbf{x}_{0})$, set $\mathbf{x}_{1}=\mathbf{y}_{0},\tau=1$. Otherwise, set $\mathbf{x}_{1}=\mathbf{x}_{0}$.
  5. 5.
    While $\tau=+\infty$
    1. (a)
      Set $t=t+1$.
    2. (b)
      Sample $(\mathbf{x}_{1+t},\mathbf{y}_{t})\sim\bar{P}((\mathbf{x}_{t},\mathbf{y}_{t-1}),\cdot)$, the common draws coupling of PIMH in Algorithm 4.
    3. (c)
      If $\mathbf{x}_{1+t}=\mathbf{y}_{t}$, set $\tau=1+t$.
  6. 6.
    Return $\tau,\mathbf{x}_{0},\mathbf{y}_{0},\mathbf{x}_{1},\mathbf{y}_{1},\ldots,\mathbf{y}_{\tau-1},\mathbf{x}_{\tau}$.
Algorithm 5 Coupled PIMH with a lag.
We relate the distribution of the meeting times generated by Algorithm 5 to the expected rejection probability in the following result.
Proposition 4.1.
Consider $\tau$ generated by Algorithm 5. Under Assumption 1, for all $t\geq 1$, we have
$\mathbb{P}(\tau>t)\leq\mathbb{E}_{\bar{q}}\left[\left(r(\mathbf{x})\right)^{t}\right].$ (23)
This connection between meeting times and expected rejection probability motivates our subsequent analysis of the expected rejection probability, which appears central in the study of IMH (e.g. Theorem 6 in Roberts & Rosenthal (2011)). Our bounds are explicit functions of $t$ and $N$.
Proposition 4.2.
Fix $p\geq 2$ and let
$\beta_{p}:=1-\frac{1}{2^{\frac{3p-2}{p-1}}q(\omega^{p})^{\frac{1}{p-1}}}.$ (24)
Under Assumptions 1-2, there exist constants $A_{p},C_{p}>0$, depending only on $p$ and $q(\omega^{p})$, such that for all $N\geq 1$, for all $t\geq 1$, the following holds:
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\right]\leq\frac{A_{p}}{N^{(t\wedge p)/2}}\beta_{p}^{t}+\frac{C_{p}}{t^{p}N^{p/2}}.$ (25)
Proposition 4.2 holds for all $t\geq 1$ and all $N\geq 1$. The bounds decay to 00 as either $N$ or $t$ approaches infinity, polynomially with rate at most $N^{-1/2}$ w.r.t. $N$, and, for fixed $N$, polynomially with rate $t^{-p}$ where $p$ is the number of assumed moments of $\omega$ under $q$. A direct consequence of the previous two propositions is a bound on the tails of the meeting times.
Proposition 4.3.
Consider $\tau$ generated by Algorithm 5. Under Assumptions 1-2, there exists $C$ such that for all $N\geq 1$ and all $t\geq 1$, if $p\geq 2$ in Assumption 2,
$\mathbb{P}(\tau>t)\leq\frac{C}{\sqrt{N}t^{p}}.$ (26)
That bound retains the slowest rates in $N$ and $t$ from the previous result. Proposition 4.3 is consistent with Proposition 8 in Middleton et al. (2019), which showed that $\mathbb{P}(\tau=1)$ approaches one as $N\to\infty$ under the assumption of bounded weights. However, our present assumptions are considerably weaker, and we provide explicit dependencies on both $N$ and $t$.
Remark 4.1.
We comment on the sharpness of the dependency on $N$ in Proposition 4.3. For $t=1$, the result reads $\mathbb{P}(\tau>1)\leq C/\sqrt{N}$. The event $\{\tau>1\}$ corresponds to the rejection of $\mathbf{x}^{\star}$ from a state $\mathbf{x}$, both $\mathbf{x},\mathbf{x}^{\star}$ being independent draws from $\bar{q}$. Here we show that we cannot improve upon the rate $N^{-1/2}$ as a function of $N$. The central limit theorem implies $\sqrt{N}(\hat{Z}_{N}(\mathbf{x})-1)\to\text{Normal}(0,q(\omega^{2})-1)$ in distribution. Therefore, $\mathbb{P}(\hat{Z}_{N}(\mathbf{x})\geq 1+N^{-1/2})\rightarrow p_{0}$ as $N\to\infty$, with $p_{0}$ depending on $q(\omega^{2})$. The same argument shows $\mathbb{P}(\hat{Z}_{N}(\mathbf{x}^{\star})\leq 1-N^{-1/2})\rightarrow p_{1}$ as $N\to\infty$, with $p_{1}$ depending on $q(\omega^{2})$. Therefore, we can choose a large enough $N$ that depends on $q(\omega^{2})$ such that $\mathbb{P}(\hat{Z}_{N}(\mathbf{x})\geq 1+N^{-1/2})\geq p_{0}/2$ and $\mathbb{P}(\hat{Z}_{N}(\mathbf{x}^{\star})\leq 1-N^{-1/2})\geq p_{1}/2$. Thus, with a constant probability $c$, $Z_{N}(\mathbf{x}^{\star})\leq 1-N^{-1/2}$ and $\hat{Z}_{N}(\mathbf{x})\geq 1+N^{-1/2}$ occur simultaneously, and thus the acceptance probability is at most $(1-N^{-1/2})/(1+N^{1/2})\leq 1-N^{-1/2}$. In turn this means that the rejection probability is at least $cN^{-1/2}$.
4.2 Polynomial convergence rates
As discussed in Section 6 of Jacob et al. (2020) and in Biswas et al. (2019), lagged chains such as those generated by Algorithm 5 can be employed to bound the total variation distance between the chain at time $t$ and its stationary distribution. We aim for bounds on $\left|P^{t}(\mathbf{x},\cdot)-\pi\right|_{\mathrm{TV}}$ that are explicit in their dependency on the iteration $t$ and the number of particles $N$. We present the following result.
Theorem 4.1.
Consider $\tau$ generated by Algorithm 5. Let $P$ be the transition kernel of the PIMH chain as in Algorithm 3. Under Assumption 1, we have that for all $t\geq 0$,
$\left|\bar{q}P^{t}-\bar{\pi}\right|_{\text{TV}}\leq\mathbb{E}\left[\max\left(0,\tau-1-t\right)\right].$ (27)
Furthermore, if Assumption 2 also holds, there exists a constant $C$, independent of $t$ and $N$, such that for all $N\geq 1$ and $t\geq 0$,
$\left|\bar{q}P^{t}-\bar{\pi}\right|_{\text{TV}}\leq\frac{C}{\sqrt{N}(1+t)^{p-1}}.$ (28)
Remark 4.2.
The case $t=0$ states that $|\bar{q}-\bar{\pi}|_{\text{TV}}\leq CN^{-1/2}$, which may seem strange as both $\bar{\pi}$ and $\bar{q}$ defined in (7)-(8) are defined on spaces growing with $N$. With the density representation of the total variation distance and $\bar{\pi}(\mathbf{x})=\hat{Z}(\mathbf{x})\bar{q}(\mathbf{x})$, we can directly compute
$|\bar{q}-\bar{\pi}|_{\text{TV}}$ $=\frac{1}{2}\int|1-\hat{Z}(\mathbf{x})|\bar{q}(\mathrm{d}\mathbf{x})=\frac{1}{2}\mathbb{E}_{\bar{q}}\left[|1-\hat{Z}(\mathbf{x})|\right]\leq\frac{1}{2}\mathbb{E}_{\bar{q}}\left[|1-\hat{Z}(\mathbf{x})|^{2}\right]^{1/2}\leq\frac{1}{2}M(2)^{1/2}N^{-1/2},$ (29)
where the first inequality is Cauchy–Schwarz and the second uses Proposition 1.1 under Assumption 2. Furthermore, in the large $N$ asymptotics we expect $1-\hat{Z}(\mathbf{x})$ to behave as a Normal distribution with mean zero and standard deviation $\sqrt{q(\omega^{2})-1}/\sqrt{N}$, so that its expectation should indeed behave as $\sqrt{(2/\pi)(q(\omega^{2})-1)}/\sqrt{N}$.
We can also state a bound for the convergence of the chain started at any initial point $\mathbf{x}\in\mathbb{X}$.
Corollary 4.1.
Under Assumptions 1-2, there exists a constant $\tilde{C}$, independent of $t$ and $N$, such that for all $N\geq 1$, $t\geq 1$, and any starting point $\mathbf{x}\in\mathbb{X}^{N}$,
$\left|P^{t}(\mathbf{x},\cdot)-\bar{\pi}\right|_{\text{TV}}\leq(r(\mathbf{x}))^{t}+\frac{\tilde{C}}{\sqrt{N}(1+t)^{p-1}}.$ (30)
Theorem 4.1 and Corollary 4.1 provide explicit bounds on the convergence rate of the PIMH algorithm. Both results are interpretable in terms of the number of iterations $t$ and the number of particles $N$, and apply to IMH as a special case when $N=1$. The difference between these results lies in the starting distribution. Practitioners would typically start the algorithm from the proposal distribution, as it is the best available approximation of the target. Corollary 4.1 reveals two phases in the convergence: an initial phase where the distance decays exponentially in $t$ but not arbitrarily with $N$, followed by a polynomial decay in both $t$ and $N$.
We add a result for the case $N=1$ i.e. standard IMH, which holds under the assumption that $q(\omega^{p})<\infty$ for $p>1$, whereas Corollary 4.1 requires $p\geq 2$ in Assumption 2. The proof is in Appendix A.4.3. As discussed in Section 4.3 the result in Proposition 4.4 is similar to existing results in the literature, although we have not found statements expressed as simply, and our proof appears to be original.
Proposition 4.4.
Consider IMH under Assumption 1, and assume $q(\omega^{p})<\infty$ for $p>1$. There exists a constant $D$ independent of $t$ such that for all $t\geq 1$, and any starting point $x\in\mathbb{X}$,
$\left|P^{t}(x,\cdot)-\pi\right|_{\text{TV}}\leq(r(x))^{t}+\frac{D}{(1+t)^{p-1}}.$ (31)
The purpose of the following example is to demonstrate that the rate $t^{-(p-1)}$ in Corollary 4.1 and Proposition 4.4 cannot be improved beyond polylogarithmic factors. The proof is provided in Appendix A.4.4.
Example 3.
Consider the IMH algorithm targeting $\pi(x):=Z_{\pi}x^{-p}$ on $[2,\infty)$, with proposal distribution $q(x):=Z_{q}\log^{2}(x)/x^{-(p+1)}$ on $[2,\infty)$, started from $x_{0}=3$. If $p\geq 2$, Assumption 2 holds with that $p$, and there exist $C<\infty$ and $t_{0}\in\mathbb{N}$ such that, for all $t\geq t_{0}$,
$\left|P^{t}(x_{0},\cdot)-\pi\right|_{\text{TV}}$ $\geq\frac{C}{t^{p-1}(\log t)^{3(p-1)}}.$
4.3 Related results on IMH
The convergence of IMH has garnered significant interest over decades, and in particular the sub-geometric rates have been studied in several works including Jarner & Roberts (2002); Douc et al. (2007); Roberts & Rosenthal (2011); Andrieu et al. (2022). One approach utilizes drift and minorization techniques (Jarner & Roberts, 2002).
Theorem 4.2 (Theorem 5.3 in Jarner & Roberts (2002)).
Let $P$ be the transition kernel of the IMH chain as in Algorithm 2. Assume that for some $r>0$,
$\displaystyle\pi(A_{\epsilon})=\mathcal{O}\left(\epsilon^{1/r}\right)\quad\text{for}\quad\epsilon\to 0,$ (32)
where $A_{\epsilon}=\left\{x\in\mathbb{X}:\omega(x)>1/\epsilon\right\}$, for any $\epsilon>0$. Then, for any $x\in\mathbb{X}$, and any $t\geq 1$, we have that
$\displaystyle\lim_{t\to\infty}(1+t)^{\beta}\left|P^{t}(x,\cdot)-\pi\right|_{\mathrm{TV}}=0,$ (33)
for any $0\leq\beta\leq\frac{s-r}{r}$, with $r<s<r+1$.
The $\mathcal{O}$ notation here is such that if $f(x)=\mathcal{O}(g(x))$ then there exists a constant $M$ such that $|f(x)|\leq M|g(x)|$ for all $x$ in the domain of $f$. Theorem 4.2 provides a polynomial rate of convergence for the IMH chain in total variation of order $o\left(t^{-1/r+\kappa}\right)$ for any $\kappa>0$ under the assumption that the tail weights satisfy the condition specified in equation (32). Notably, under the assumption $q(\omega^{p})<\infty$, the condition in (32) is satisfied for $r=1/(p-1)$, using Markov’s inequality. Our Proposition 4.4 differs slightly as our bounds are in $t^{-(p-1)}$ instead of $t^{-(p-1)+\kappa}$ for some arbitrarily small $\kappa>0$. Similar results can be obtained using weak Poincaré inequalities as described in Remark 29 of Andrieu et al. (2022), under $\pi(\omega^{p})$ with $p>1$ which amounts to our Assumption 2 with $p>2$.
Our bounds in Theorem 4.1 and Corollary 4.1 have the advantage of providing an explicit dependency on $N$ in the case of PIMH, which is critical for the results on bias removal in Section 5.
4.4 Comparing the biases of IS and IMH
We can now compare the bias of IMH and IS under the assumption $q(\omega^{p})<\infty$ for $p\geq 2$. Suppose that the goal is to obtain a single sample close to $\pi$ in total variation distance, with a budget of $N$ samples from $q$ and $N$ evaluations of $\omega$.
  • One approach is sampling-importance resampling (SIR), which refers to the following procedure. First, run Algorithm 1. Then, draw $k\sim\text{Categorical}(\omega(x_{1}),\ldots,\omega(x_{N}))$ and return $x_{k}$. For a test function $f$ with $|f|_{\infty}\leq 1$, under the conditions of Theorem 2.1 the marginal distribution $\mu_{N}^{\text{SIR}}$ of $x_{k}$ satifies
    $\mu_{N}^{\text{SIR}}(f)-\pi(f)\sim_{N\to\infty}-q(\omega^{2}\cdot(f-\pi(f)))N^{-1}.$ (34)
    Skare et al. (2003) provide a similar result in the case of bounded weights, and propose a modification of SIR to reduce this bias to $N^{-2}$.
  • On the other hand, Theorem 4.1 suggests that IMH (with one particle) after $N$ iterations provides a sample from a distribution $qP^{N-1}$, for which, under the condition $q(\omega^{p})<\infty$,
    $\sup_{f:|f|_{\infty}\leq 1}\left\{qP^{N-1}(f)-\pi(f)\right\}\leq CN^{-(p-1)}.$ (35)
    Thus, the bias may be reduced faster when increasing the computing budget in IMH rather than in SIR, as soon as $p>2$; for the modified SIR of Skare et al. (2003), IMH is faster as soon as $p>3$.
In the case of bounded weights, MCMC methods such as PIMH or particle Gibbs (Andrieu et al., 2010) are geometrically ergodic (e.g. Lee et al., 2020) and the bias comparison is clearly at the advantage of MCMC algorithms, as discussed in Cardoso et al. (2022).
5 Bias removal for self-normalized importance sampling
5.1 Construction
The bias of importance sampling was described in Section 2, and that of IMH in Section 4. Here we consider the removal of the bias, and the associated cost. For this purpose, Middleton et al. (2019) employ common random numbers couplings of PIMH and the approach of Glynn & Rhee (2014). We pursue this strategy.
Upon running Algorithm 5 with $N\geq 1$, with $\tau=\inf\{t\geq 1:\mathbf{x}_{t}=\mathbf{y}_{t-1}\}$, one can compute the following unbiased estimator:
$\hat{F}_{u}=\hat{F}(\mathbf{x}_{0})+\sum_{t=1}^{\tau-1}\{\hat{F}(\mathbf{x}_{t})-\hat{F}(\mathbf{y}_{t-1})\},$ (36)
where $\hat{F}:\mathbf{x}\mapsto\sum_{n=1}^{N}\omega(x_{n})f(x_{n})/(\sum_{n=1}^{N}\omega(x_{n}))$ and $f$ is a test function. By convention the sum in (36) is zero in the event $\{\tau=1\}$, and it is also equal to the infinite sum $\sum_{t=1}^{\infty}\{\hat{F}(\mathbf{x}_{t})-\hat{F}(\mathbf{y}_{t-1})\}$ since $\hat{F}(\mathbf{x}_{t})=\hat{F}(\mathbf{y}_{t-1})$ from time $\tau$ onward. The lack of bias can be seen via a telescopic sum argument, since $\mathbf{x}_{t}$ and $\mathbf{y}_{t}$ have the same marginal distribution for all $t$, and provided that limit and expectation can be swapped. Since $\mathbf{x}_{0}\sim\bar{q}$, $\hat{F}(\mathbf{x}_{0})$ is the (biased) IS estimator. In contrast, $\hat{F}_{u}$ in (36) is unbiased, under some conditions. Middleton et al. (2019) consider the case where $\omega$ is uniformly upper bounded, and they show that (36) can have a finite variance. Below we work under the weaker Assumptions 1-2, and we derive results on the moments of unbiased IS and on its comparison with regular IS.
Remark 5.1.
(36) is an instance of unbiased MCMC (Jacob et al., 2020; Atchadé & Jacob, 2024), where here MCMC is PIMH, and various generic improvements could be considered, such as increasing the lag between the chains, or introducing a burn-in parameter. However, in the particular case of PIMH, the number of particles $N$ is a key parameter and here we focus on the regime $N\to\infty$, in which case $\hat{F}_{u}$ naturally compares with $\hat{F}(\mathbf{x}_{0})$, which is the regular IS estimator. Hence we call (36) the unbiased self-normalized importance sampling estimator (UIS), but it could also be called unbiased PIMH.
5.2 Moments of unbiased self-normalized importance sampling
We subtract $\pi(f)$ from all terms in (36) to obtain
$\hat{F}_{u}-\pi(f)=\hat{F}(\mathbf{x}_{0})-\pi(f)+\sum_{t=1}^{\infty}\{\hat{F}(\mathbf{x}_{t})-\hat{F}(\mathbf{y}_{t-1})\}\mathds{1}(\tau>t).$ (37)
We introduce the notation
$\Delta_{t}$ $=\hat{F}(\mathbf{x}_{t})-\hat{F}(\mathbf{y}_{t-1}),\quad\text{BC}=\sum_{t=1}^{\infty}\Delta_{t}\mathds{1}(\tau>t),$ (38)
where BC stands for the bias cancellation term. Using Minkowski’s inequality, the moments of the error of $\hat{F}_{u}$ can be bounded by the moments of the error of the IS estimator $\hat{F}(\mathbf{x}_{0})$, as in Theorem 2.3, and the moments of BC.
A first result is that, for bounded test functions $f$, $\hat{F}_{u}$ has as many moments as the meeting time $\tau$, which is nearly $p$ under Assumption 2. The proof is in Appendix A.5.1.
Proposition 5.1.
Assume that $|f|_{\infty}\leq 1$ and let $s\geq 1$. If Assumptions 1-2 hold with $p\geq 2$ and $p>s$, then the meeting time $\tau$ has $s$ finite moments, and the unbiased self-normalized importance sampling (UIS) estimator $\hat{F}_{u}$ in (36) has $s$ finite moments.
To deal with unbounded test functions, we need to control the moments of the terms $\Delta_{t}$. For this we derive the following result about PIMH at any iteration $t$, under the same conditions as Theorem 2.3. The proof is in Appendix A.5.2.
Proposition 5.2.
Assume that there exist $p\in[2,\infty)$ and $r\in[2,\infty]$ such that $q(\omega^{p})<\infty$ and $q(f^{r})<\infty$, and $q(f^{2}\cdot\omega^{2})<\infty$. Let $(\mathbf{x}_{t})$ be the PIMH chain started from $\mathbf{x}_{0}\sim\bar{q}$. Then, for any $2\leq s\leq pr/(p+r+2)$ and any $N\geq 1$, there exists $C$ such that for all $t\geq 0$:
$\mathbb{E}_{\mathbf{x}_{0}\sim\bar{q}}\left[\left\lvert\hat{F}(\mathbf{x}_{t})-\pi(f)\right\rvert^{s}\right]\leq CN^{-s/2},$
where the constant $C$ depends on $r,p,s,q(f^{r}),q(\omega^{p}),q(f^{2}\cdot\omega^{2})$. When $r=\infty$, the statement holds for $f$ such that $|f|_{\infty}<\infty$ and all $s\leq p$.
By Minkowski’s inequality, under the conditions of Proposition 5.2, the moments of $\Delta_{t}$ have similar bounds. This can be used to obtain the following result, proven in Appendix A.5.3.
Proposition 5.3.
Assume that there exist $p\in(2,\infty)$ and $r\in[2,\infty]$ such that $q(\omega^{p})<\infty$ and $q(|f|^{r})<\infty$, and $q(f^{2}\cdot\omega^{2})<\infty$, then for any $2\leq s<p$ such that $pr/(p+r+2)>ps/(p-s)$, and for any $N\geq 1$, the unbiased importance sampling (UIS) estimator satisfies:
$\mathbb{E}\left[\left\lvert\hat{F}_{u}-\pi(f)\right\rvert^{s}\right]\leq CN^{-s/2},$
where the constant $C$ depends on $r,p,s,q(|f|^{r}),q(\omega^{p}),q(f^{2}\cdot\omega^{2})$. When $r=\infty$, the statement holds for $f$ such that $|f|_{\infty}<\infty$ and all $s<p$ such that $p>sp/(p-s)$.
Two finite moments ($s=2$ in the above results) are particularly useful. Indeed, a confidence interval for $\pi(f)$ can be constructed from independent copies of $\hat{F}_{u}$ using the Central Limit Theorem as long as $\mathbb{V}[\hat{F}_{u}]<\infty$. Beyond the empirical average, robust mean estimation strategies could be envisioned. The typical assumption in robust mean estimation is that users have access to a random variable $X$ with mean $\mu$ of interest, and a finite variance $\sigma^{2}$ that may be unknown. The goal is to aggregate independent copies of $X$ into an estimator of $\mu$ with lighter tails than the empirical average. Such results can be found in Devroye et al. (2016); Lugosi & Mendelson (2019b, a); Lecué & Lerasle (2020); Minsker & Ndaoud (2021). These techniques are directly applicable with $X=\hat{F}_{u}$ and $\mu=\pi(f)$, whereas the application of robust mean estimation to self-normalized importance sampling is not a priori straightforward, and was explored in Dau (2022) in the case of median-of-means.
5.3 The asymptotic price of bias removal
We consider the price of debiasing self-normalized importance sampling in terms of inefficiency, defined by the variance multiplied by the average cost. We start by comparing the mean squared errors of unbiased and regular IS. From (37), we take the square and use Cauchy–Schwarz to obtain
$\left|\mathbb{E}\left[(\hat{F}_{u}-\pi(f))^{2}\right]-\mathbb{E}\left[(\hat{F}(\mathbf{x}_{0})-\pi(f))^{2}\right]\right|$ $\leq\sqrt{\mathbb{E}\left[(\hat{F}(\mathbf{x}_{0})-\pi(f))^{2}\right]\cdot\mathbb{E}[\text{BC}^{2}]}+\mathbb{E}[\text{BC}^{2}],$ (39)
with $\text{BC}=\sum_{t=1}^{\tau-1}\{\hat{F}(\mathbf{x}_{t})-\hat{F}(\mathbf{y}_{t-1})\}$.
The mean squared error (MSE) of IS, which is the term $\mathbb{E}[(\hat{F}(\mathbf{x}_{0})-\pi(f))^{2}]$, is of order $N^{-1}$ under conditions stated in Theorem 2.3. If we can bound $\mathbb{E}[\text{BC}^{2}]$ by a term that decreases faster than $N^{-1}$, then the MSE of UIS would be asymptotically equivalent to that of IS. Intuitively, the bias cancellation term goes to zero for two reasons: first because $\tau$ goes to one as $N\to\infty$, and the bias cancellation term equals zero in the event $\{\tau=1\}$. Secondly, each term $\hat{F}(\mathbf{x}_{t})-\hat{F}(\mathbf{y}_{t-1})$ goes to 00 as $N\to\infty$, under the conditions of Proposition 5.2. We obtain the following result, proven in Appendix A.5.4.
Proposition 5.4.
Let $\hat{F}_{u}$ be the UIS estimator defined as (36) and $\hat{F}(\mathbf{x})$ with $\mathbf{x}\sim\bar{q}$ be the IS estimator. Suppose that the assumptions of Proposition 5.3 are satisfied with $s=2$, that is: $p>2$ and $r>2$ such that $q(\omega^{p})<\infty$ and $q(|f|^{r})<\infty$, and $q(f^{2}\cdot\omega^{2})<\infty$, with $2p+4r+4<r\cdot p$. Then the mean squared error of $\hat{F}_{u}$ and that of $\hat{F}(\mathbf{x})$ are asymptotically equivalent:
$\lim_{N\to\infty}N\cdot\mathbb{E}\left[(\hat{F}_{u}-\pi(f))^{2}\right]=\lim_{N\to\infty}N\cdot\mathbb{E}_{\mathbf{x}\sim\bar{q}}\left[(\hat{F}(\mathbf{x})-\pi(f))^{2}\right].$
The assumption $2p+4r+4<r\cdot p$ is for example satisfied if $r=\infty$ and $p=4+\epsilon$ with an arbitrary $\epsilon>0$, or if $p=5$ and $r=15$. However it cannot be satisfied with $p\leq 4$.
The cost of UIS is that of running Algorithm 5. It starts with two draws from $\bar{q}$, i.e. $2N$ draws from $q$, and as many evaluations of the weight function $\omega$. Then either $\tau=1$ or the algorithm enters its while loop up to the meeting time $\tau$, drawing $N$ new particles at each iterate of the loop. Counting the cost in terms of the number of evaluations of $\pi$, UIS has an overall cost of $\mathcal{C}=2N+N(\tau-1)$. If Assumption 2 holds with $p\geq 2$, using Proposition 4.3 then $\mathbb{E}[\tau]=\sum_{t\geq 0}\mathbb{P}(\tau>t)$ is finite, and furthermore, $\mathbb{E}[\mathcal{C}]\sim_{N\to\infty}2N$. We summarise these first observations in the next statement.
Proposition 5.5.
The cost of the UIS estimator $\hat{F}_{u}$ in (36) is $\mathcal{C}=2N+N(\tau-1)$, and, under Assumption 2 for $p\geq 2$, there exists a constant $C$ such that $\mathbb{E}[\tau]\leq 1+C/\sqrt{N}$.
Combining Proposition 5.4 with the above result, we see that under some conditions, the inefficiency $\mathbb{E}[\mathcal{C}]\mathbb{V}[\hat{F}_{u}]$ is equivalent to twice that of self-normalized importance sampling as $N\to\infty$. Indeed, the mean squared errors are equivalent but the cost of UIS behaves as $2N$ instead of $N$ for IS. In Algorithm 5 the $N$ particles in $\mathbf{y}_{0}$ are required to determine $\tau$ but in the event $\{\tau=1\}$ they do not participate directly in the estimator $\hat{F}_{u}$.
5.4 An improved unbiased estimator
A trick provides a remedy, and cuts the asymptotic inefficiency by a half. We can view $\hat{F}_{u}$ as a deterministic function of initial states $\mathbf{x}_{0}$ and $\mathbf{y}_{0}$ drawn from $\bar{q}$ independently, as well as extra variables (a sequence of proposals from $\bar{q}$, a sequence of uniform random variables) that we collectively label $\zeta$; we denote $\hat{F}_{u}$ by $\hat{F}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)$. Then we define the symmetrized UIS estimator:
$\tilde{F}_{u}=\frac{1}{2}\left(\hat{F}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)+\hat{F}_{u}(\mathbf{y}_{0},\mathbf{x}_{0},\zeta)\right).$ (40)
Computing (40) only requires simple modifications of Algorithm 5. Indeed, either $\hat{Z}(\mathbf{x}_{0})\geq\hat{Z}(\mathbf{y}_{0})$ or $\hat{Z}(\mathbf{x}_{0})<\hat{Z}(\mathbf{y}_{0})$. In the first case, we always have $\hat{F}_{u}(\mathbf{y}_{0},\mathbf{x}_{0},\zeta)=\hat{F}(\mathbf{y}_{0})$, and $\hat{F}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)$ can be computed following Algorithm 5 and (36). In the second case, we always have $\hat{F}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)=\hat{F}(\mathbf{x}_{0})$, and $\hat{F}_{u}(\mathbf{y}_{0},\mathbf{x}_{0},\zeta)$ can be computed following Algorithm 5 and (36) with the role of $\mathbf{x}_{0}$ and $\mathbf{y}_{0}$ swapped. That trick amounts to a Rao–Blackwellization over the arbitrary specification of which draws from $\bar{q}$ are used as $\mathbf{x}_{0}$ and as $\mathbf{y}_{0}$. The following statement is a mild variation of the previous results and is stated without a proof.
Proposition 5.6.
Consider the symmetrized UIS estimator $\tilde{F}_{u}$ in (40). Suppose that the conditions of Proposition 5.4 are satisfied. Then $\tilde{F}_{u}$ is an unbiased estimator of $\pi(f)$ for any $N\geq 1$, with finite expected cost and finite variance, and its inefficiency is equivalent to that of IS as $N\to\infty$.
The result supports the intuition that $\tilde{F}_{u}$ should be preferred to $\hat{F}_{u}$ in practice, and that the asymptotic efficiency of $\tilde{F}_{u}$ matches that of self-normalized importance sampling as $N\to\infty$, under some conditions on $\omega$ and $f$ such as those in Proposition 5.4.
Acknowledgements.
Through the second author and CY Initiative Emergence, this project was supported by France 2030.
Appendix A Proofs
A.1 Proofs of Section 1
A.1.1 Proof of (11)
We introduce the following extended state space justification of PIMH. Let the target and proposal distributions be
$\tilde{\pi}(x_{1},\ldots,x_{N},k)$ $=\frac{1}{N}\pi(x_{k})\prod_{n\neq k}q(x_{n}),$ (41)
$\tilde{q}(x_{1},\ldots,x_{N},k)$ $=\frac{\omega(x_{k})}{\sum_{n=1}^{N}\omega(x_{n})}\prod_{n=1}^{N}q(x_{n}).$ (42)
The state space is $\mathbb{X}^{N}\times\{1,\ldots,N\}$, i.e. $k$ is an integer in $\{1,\ldots,N\}$. The ratio of target divided by proposal on the extended space takes the following form:
$\tilde{\omega}(x_{1},\ldots,x_{N},k)=\frac{\tilde{\pi}(x_{1},\ldots,x_{N},k)}{\tilde{q}(x_{1},\ldots,x_{N},k)}=\frac{1}{N}\sum_{n=1}^{N}\omega(x_{n}).$ (43)
In other words, the ratio simplifies to $\hat{Z}(\mathbf{x})$, thus IMH with proposal $\tilde{q}$ and target $\tilde{q}$ is equivalent to Algorithm 3.
Draws from the proposal $\tilde{q}(x_{1},\ldots,x_{N},k)$ can be obtained by first sampling $x_{1},\ldots,x_{N}$ i.i.d. from $q$, then computing the importance weights $\omega(x_{n})$ for $n=1,\ldots,N$, and finally sampling $k$ from a Categorical distribution with probabilities given by $\omega(x_{n})/\sum_{m=1}^{M}\omega(x_{m})$. Under the extended target $\tilde{\pi}$, we have the following marginals and conditionals.
  • The marginal distribution of $k$ is the Uniform distribution on $\{1,\ldots,N\}$.
  • The marginal distribution of $x_{1},\ldots,x_{N}$ is obtained as
    $\tilde{\pi}(x_{1},\ldots,x_{N})=\sum_{n=1}^{N}\frac{1}{N}\pi(x_{n})\prod_{m\neq n}q(x_{m}).$ (44)
  • The conditional distribution of $k$ given $x_{1},\ldots,x_{N}$ can be obtained by dividing (41) by (44),
    $\tilde{\pi}(k|x_{1},\ldots,x_{N})=\frac{\omega(x_{k})}{\sum_{n=1}^{N}\omega(x_{n})}.$ (45)
  • The conditional distribution of $x_{1},\ldots,x_{N}$ given $k$ can be obtained by dividing (41) by $N^{-1}$, and is given by
    $\tilde{\pi}(x_{1},\ldots,x_{N}|k)=\pi(x_{k})\prod_{n\neq k}q(x_{n}),$ (46)
    and in particular $\tilde{\pi}(x_{k}|k)$ is $\pi(x_{k})$.
Note that $f(x_{k})$ can be written as a more explicit function of $(x_{1},\ldots,x_{N},k)$, as $\sum_{n=1}^{N}f(x_{n})\mathds{1}(n=k)$. We use the above properties to obtain the following equalities, as in Andrieu et al. (2010), Appendix B.5, either conditioning on $k$ in the first line, or on $x_{1},\ldots,x_{N}$ in the second line:
$\mathbb{E}_{\bar{\pi}}[\sum_{n=1}^{N}f(x_{n})\mathds{1}(n=k)]$ $=\mathbb{E}_{\bar{\pi}}[\mathbb{E}_{\bar{\pi}}[\sum_{n=1}^{N}f(x_{n})\mathds{1}(n=k)|k]]=\mathbb{E}_{\pi}[f],$ (47)
$=\mathbb{E}_{\bar{\pi}}[\mathbb{E}_{\bar{\pi}}[\sum_{n=1}^{N}f(x_{n})\mathds{1}(n=k)|x_{1},\ldots,x_{N}]]=\mathbb{E}_{\bar{\pi}}\left[\frac{\sum_{n=1}^{N}\omega(x_{n})f(x_{n})}{\sum_{n=1}^{N}\omega(x_{n})}\right].$ (48)
We obtain (11), i.e. the equality $\mathbb{E}_{\pi}[f]=\mathbb{E}_{\mathbf{x}\sim\bar{\pi}}[\hat{F}(\mathbf{x})]$, with $\hat{F}$ as in (4).
A.1.2 Proofs of Proposition 1.1 and Lemma 1.1
Proof of Proposition 1.1.
Using the result of Petrov (1975), Section III.5 (item 16, p. 60): for $X_{1},\ldots,X_{N}$ independent variables with zero mean and $p$ finite moments, $p\geq 2$, we have
$\mathbb{E}\left[\left|\sum_{n=1}^{N}X_{n}\right|^{p}\right]\leq m(p)N^{p/2-1}\sum_{n=1}^{N}\mathbb{E}[|X_{n}|^{p}],$
where $m(p)$ is a positive number depending only on $p$. As described in Ren & Liang (2001), the constant $m(p)$ satisfies $(m(p))^{1/p}\leq p-1$; in fact they provide a sharper bound, but we do not need it here. For i.i.d. variables the right-hand side becomes $m(p)N^{p/2}\mathbb{E}[|X_{1}|^{p}]$. If we consider the average instead of the sum on the left, then the right-hand side becomes $m(p)N^{-p/2}\mathbb{E}[|X_{1}|^{p}]$. Since $q(\omega)=1$ and assuming that $q(\omega^{p})<\infty$, we define $X_{n}=\omega_{n}-1$ and apply the above result to obtain
$\mathbb{E}\left[|\hat{Z}(\mathbf{x})-1|^{p}\right]\leq(p-1)^{p}N^{-p/2}q((\omega-1)^{p}).$
Next we can use the $C_{p}$-inequality, which, for $p\geq 1$, reads:
$\mathbb{E}\left[|X+Y|^{p}\right]\leq 2^{p-1}\left(\mathbb{E}\left[|X|^{p}\right]+\mathbb{E}\left[|Y|^{p}\right]\right).$
That inequality with $X=\omega$ and $Y=-1$ delivers $q((\omega-1)^{p})\leq 2^{p-1}(1+q(\omega^{p}))$. This establishes (13).
For the non-centred moment, we proceed as follows:
$\mathbb{E}\left[|\hat{Z}(\mathbf{x})-1+1|^{p}\right]=\sum_{k=0}^{p}{p\choose k}\mathbb{E}\left[|\hat{Z}(\mathbf{x})-1|^{k}\right],$
then using Hölder’s inequality, this is less than
$\sum_{k=0}^{p}{p\choose k}\mathbb{E}\left[|\hat{Z}(\mathbf{x})-1|^{p}\right]^{k/p}\leq\sum_{k=0}^{p}{p\choose k}\left((p-1)^{p}N^{-p/2}2^{p-1}(1+q(\omega^{p}))\right)^{k/p}.$
From the binomial theorem, $\sum_{k=0}^{p}{p\choose k}a^{k}=(a+1)^{p}$, we obtain
$\mathbb{E}\left[|\hat{Z}(\mathbf{x})|^{p}\right]\leq\left(1+(p-1)N^{-1/2}2^{1-1/p}(1+q(\omega^{p}))^{1/p}\right)^{p}.$
This bound goes to one as $N\to\infty$. ∎
Proof of Lemma 1.1.
Using Markov’s inequality and Proposition 1.1, we have:
$\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})\geq 1+z\right)$ $=\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})-1\geq z\right)$ (49)
$\leq\frac{\mathbb{E}_{\bar{q}}\left[\left|\hat{Z}(\mathbf{x})-1\right|^{p}\right]}{z^{p}}$ (50)
$\leq\frac{M(p)N^{-p/2}}{z^{p}}.$ (51)
The second line uses Markov’s inequality, and the third line applies Proposition 1.1. ∎
A.2 Proofs of Section 2
A.2.1 Proof of Theorem 2.1
We start with a technical result on the inverse moments of averages, which may be well-known.
Proposition A.1.
Let $r\geq 1$, $(x_{j})_{j\geq 0}$ a sequence of i.i.d. random variables with distribution $q$ on $\mathbb{X}$, and suppose that $\omega:\mathbb{X}\to(0,\infty)$ such that $q(\omega^{-\eta})<\infty$ for some $\eta>0$. Write $\omega_{j}=\omega(x_{j})$ for all $j=1,\ldots,N$. Then, for $N>\lfloor r/\eta\rfloor+1$, we have that
$\mathbb{E}\left[\left(\frac{N}{\omega_{1}+\cdots+\omega_{N}}\right)^{r}\right]\leq 2^{r}q(\omega^{-\eta})^{r/\eta}<\infty.$
Proof.
Let $\hat{W}=\frac{1}{N}(\omega_{1}+\cdots+\omega_{N})$. We will proceed by splitting the variables into blocks of size $j$ for $r/\eta\leq j\leq N$, which is possible by assumption, as follows: for $k\leq\lfloor N/j\rfloor$ we define
$\hat{W}^{j}_{k}:=\frac{1}{j}\left(\omega_{kj+1}+\cdots+\omega_{(k+1)j}\right)\qquad\text{and}\qquad\hat{W}^{j}_{\lfloor N/j\rfloor+1}:=\frac{1}{j}\left(\omega_{\lfloor N/j\rfloor j+1}+\cdots+\omega_{N}\right),$
where the final block may have fewer than $j$ elements. We lower bound $\hat{W}$ by dropping the last block if it has length strictly less than $j$,
$\hat{W}$ $\geq\frac{\hat{W}^{j}_{1}+\cdots+\hat{W}^{j}_{\left\lfloor\frac{N}{j}\right\rfloor}}{\frac{N}{j}}=\frac{\hat{W}^{j}_{1}+\cdots+\hat{W}^{j}_{\left\lfloor\frac{N}{j}\right\rfloor}}{\left\lfloor\frac{N}{j}\right\rfloor}\cdot\frac{\left\lfloor\frac{N}{j}\right\rfloor}{\frac{N}{j}}=:\frac{\left\lfloor\frac{N}{j}\right\rfloor}{\frac{N}{j}}\cdot\widetilde{W}.$
Since the mapping $m:x\mapsto 1/x^{r}$ is monotone decreasing and convex (we assumed $r\geq 1$), we have:
$\mathbb{E}\left[\hat{W}^{-r}\right]$ $=\mathbb{E}\left[m(\hat{W})\right]\leq\left(\frac{\frac{N}{j}}{\left\lfloor\frac{N}{j}\right\rfloor}\right)^{r}\mathbb{E}\left[m(\widetilde{W})\right]$
$\leq\left(1+\frac{1}{\left\lfloor\frac{N}{j}\right\rfloor}\right)^{r}\cdot\frac{1}{\left\lfloor\frac{N}{j}\right\rfloor}\sum_{k=1}^{\left\lfloor\frac{N}{j}\right\rfloor}\mathbb{E}[m(\hat{W}^{j}_{k})]$
$\leq\frac{2^{r}}{\left\lfloor\frac{N}{j}\right\rfloor}\sum_{k=1}^{\left\lfloor\frac{N}{j}\right\rfloor}\mathbb{E}\left[m\left(\hat{W}^{j}_{k}\right)\right].$
To proceed, we utilize the arithmetic-geometric mean inequality, which states that for non-negative numbers $a_{1},a_{2},\ldots,a_{j}$:
$\frac{a_{1}+a_{2}+\cdots+a_{j}}{j}\geq\left(a_{1}\cdot a_{2}\cdots a_{j}\right)^{\frac{1}{j}}.$
Applying this inequality, we obtain under the assumption that $q(\omega^{-\eta})<\infty$,
$\mathbb{E}\left[\left(\frac{j}{\omega_{1}+\dots+\omega_{j}}\right)^{r}\right]$ $\leq\mathbb{E}\left[\prod_{k=1}^{j}\left(\frac{1}{\omega_{k}}\right)^{r/j}\right]=\left(q(\omega^{-\frac{r}{j}})\right)^{j}$
$\leq q(\omega^{-\eta})^{j\frac{r}{j\eta}}=q(\omega^{-\eta})^{\frac{r}{\eta}}<\infty,$
where we have used Hölder’s inequality with the exponent $r^{\prime}=\eta j/r\geq 1$, by the choice of $j\geq r/\eta$. This yields the desired result. ∎
Proof of Theorem 2.1.
We first write the rescaled bias of normalized importance sampling as
$N\times\mathbb{E}_{\mathbf{x}\sim\bar{q}}\left[\hat{F}(\mathbf{x})-\pi(f)\right]$ $=\mathbb{E}\left[\frac{\sum_{n=1}^{N}\omega(x_{n})(f(x_{n})-\pi(f))}{\sum_{n=1}^{N}\omega(x_{n})/N}\right]$ (52)
$=N\mathbb{E}\left[\frac{\omega(x_{1})(f(x_{1})-\pi(f))}{\sum_{n=1}^{N}\omega(x_{n})/N}\right]\quad\text{by identity in distribution}$ (53)
$=N\mathbb{E}\left[\frac{\omega(x_{1})(f(x_{1})-\pi(f))}{\sum_{n=2}^{N}\omega(x_{n})/N}\right]$ (54)
$+N\mathbb{E}\left[\omega(x_{1})(f(x_{1})-\pi(f))\left\{\frac{1}{\sum_{n=1}^{N}\omega(x_{n})/N}-\frac{1}{\sum_{n=2}^{N}\omega(x_{n})/N}\right\}\right].$ (55)
By independence and $\mathbb{E}[\omega(x_{1})f(x_{1})]=\pi(f)$, the first expectation is zero. For the second term,
$\frac{1}{\sum_{n=1}^{N}\omega(x_{n})/N}-\frac{1}{\sum_{n=2}^{N}\omega(x_{n})/N}=\frac{-\omega(x_{1})/N}{(\sum_{n=1}^{N}\omega(x_{n})/N)(\sum_{n=2}^{N}\omega(x_{n})/N)}.$ (56)
Thus, we can write
$N\times\mathbb{E}_{\mathbf{x}\sim\bar{q}}\left[\hat{F}(\mathbf{x})-\pi(f)\right]$ $=-N\mathbb{E}\left[\frac{\omega(x_{1})^{2}(f(x_{1})-\pi(f))/N}{(\sum_{n=1}^{N}\omega(x_{n})/N)(\sum_{n=2}^{N}\omega(x_{n})/N)}\right],$ (57)
and we further re-use (56) so that only $x_{j}$’s with $j\neq 1$ appear in the denominator of the leading term:
$N\times\mathbb{E}_{\mathbf{x}\sim\bar{q}}\left[\hat{F}(\mathbf{x})-\pi(f)\right]$ $=-\mathbb{E}\left[\frac{\omega(x_{1})^{2}(f(x_{1})-\pi(f))}{(\sum_{n=2}^{N}\omega(x_{n})/N)^{2}}\right]$ (58)
$-\mathbb{E}\left[\frac{-\omega(x_{1})}{(\sum_{n=1}^{N}\omega(x_{n})/N)(\sum_{n=2}^{N}\omega(x_{n})/N)}\times\frac{\omega(x_{1})^{2}(f(x_{1})-\pi(f))/N}{\sum_{n=2}^{N}\omega(x_{n})/N}\right].$ (59)
Having different $x_{j}$’s in the numerator and denominator, and using their independence, the leading term in (58) is $-q(\omega^{2}\cdot(f-\pi(f)))\mathbb{E}[(\sum_{n=2}^{N}\omega(x_{n})/N)^{-2}]$.
Let $T_{N}=N^{-1}\sum_{n=2}^{N}\omega(x_{n})$. By the strong law of large numbers, $T_{N}^{-2}\xrightarrow{a.s.}1$ as $N\to\infty$. To strengthen this to convergence in $L^{1}$ of $T_{N}^{-2}$ to $1$, we use uniform integrability, e.g. Billingsley (1999), Theorem 3.5. A criterion for uniform integrability is (3.18) in Billingsley (1999), which is satisfied here since $\sup_{N}\mathbb{E}[T_{N}^{-3}]<\infty$ using $q(\omega^{-\eta})<\infty$ and Proposition A.1 with $r=3$, thus requiring $N>\lfloor 3/\eta\rfloor+1$.
It remains to show that the term in (59) goes to zero as $N\to\infty$. First we use the positivity of $\omega$ and the independence of $x_{j}$’s to get
$\mathbb{E}\left[\frac{\omega(x_{1})^{3}(f(x_{1})-\pi(f))/N}{(\sum_{n=1}^{N}\omega(x_{n})/N)(\sum_{n=2}^{N}\omega(x_{n})/N)^{2}}\right]$ $\leq\mathbb{E}\left[\left|\frac{\omega(x_{1})^{3}(f(x_{1})-\pi(f))/N}{(\sum_{n=2}^{N}\omega(x_{n})/N)^{3}}\right|\right]$ (60)
$=\frac{1}{N}\cdot\mathbb{E}\left[\left|\omega(x_{1})^{3}(f(x_{1})-\pi(f))\right|\right]\mathbb{E}\left[(\sum_{n=2}^{N}\omega(x_{n})/N)^{-3}\right].$ (61)
The first expectation is finite by assumption. Using Proposition A.1, $\mathbb{E}\left[(\sum_{n=2}^{N}\omega(x_{n})/N)^{-3}\right]\leq(N/(N-1))^{3}2^{3}q(\omega^{-\eta})^{3/\eta}$ when $N>\lfloor 3/\eta\rfloor+1$. Thus, this term in (59) behaves as a constant divided by $N$. ∎
A.2.2 Proof of Theorem 2.3
Proof of Theorem 2.3.
We write the IS estimator: $(\sum_{n=1}^{N}f(x_{n})\omega(x_{n}))/(\sum_{n=1}^{N}\omega(x_{n}))$, where $x_{1},\ldots,x_{n}$ are i.i.d. from $q$. We write the average weight: $q^{N}(\omega):=\sum_{n=1}^{N}\omega(x_{n})/N$.
First, it is enough to consider the case where the test function $f$ is non-negative. Indeed, for a general function $f$ we write $f=f_{+}-f_{-}$ where $f_{+}(x):=\max\{f(x),0\}$ and $f_{-}(x):=-\min\{f(x),0\}$. Then
$\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-\pi(f)\right\rvert=\left\lvert\frac{q^{N}(f_{+}\omega)-q^{N}(f_{-}\omega)}{q^{N}(\omega)}-(\pi(f_{+})-\pi(f_{-}))\right\rvert\leq\left\lvert\frac{q^{N}(f_{+}\omega)}{q^{N}(\omega)}-\pi(f_{+})\right\rvert+\left\lvert\frac{q^{N}(f_{-}\omega)}{q^{N}(\omega)}-\pi(f_{-})\right\rvert.$
Using $(a+b)^{s}\leq 2^{s-1}(a^{s}+b^{s})$, and applying the result for non-negative functions $f_{+}$ and $f_{-}$ separately, we obtain the result for general $f$. Thus, we now assume that $f$ takes non-negative values.
We write the absolute error between the IS estimator with the target $\pi(f)=q(f\omega)$ in two different ways. The first is:
$\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-q(f\omega)\right\rvert\leq\max_{1\leq i\leq N}f(x_{i})+q(f\omega).$ (62)
The second is:
$\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-q(f\omega)\right\rvert\leq\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-\frac{q(f\omega)}{q^{N}(\omega)}\right\rvert+q(f\omega)\left\lvert\frac{1}{q^{N}(\omega)}-1\right\rvert.$
Now we consider two cases: 1) $\lvert q^{N}(\omega)-1\rvert>0.5$, 2) $\lvert q^{N}(\omega)-1\rvert\leq 0.5$. We will separately bound the expected error under the two cases using the two inequalities above.
We start with the first case, and we assume $r<\infty$. First, we use $(a+b)^{s}\leq 2^{s-1}(a^{s}+b^{s})$ to write
$\mathbb{E}\left[\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-q(f\omega)\right\rvert^{s}\mathds{1}(\lvert q^{N}(\omega)-1\rvert>0.5)\right]$ $\leq\mathbb{E}\left[\left(\max_{1\leq i\leq N}f(x_{i})+q(f\omega)\right)^{s}\mathds{1}(\lvert q^{N}(\omega)-1\rvert>0.5)\right]$
$\leq 2^{s-1}\mathbb{E}\left[\left(\max_{1\leq i\leq N}f(x_{i})\right)^{s}\mathds{1}(\lvert q^{N}(\omega)-1\rvert>0.5)\right]\break$
$+2^{s-1}\left(q(f\omega)\right)^{s}\mathbb{P}[\lvert q^{N}(\omega)-1\rvert>0.5].$
The second term leads to a bound in $N^{-s/2}$ using Markov’s inequality as in Lemma 1.1, since $q(\omega^{s})<\infty$ under the assumptions. The first term is dealt with first using Hölder’s inequality with exponents $r/s$ and $(1-s/r)^{-1}$,
$\mathbb{E}\left[\left(\max_{1\leq i\leq N}f(x_{i})\right)^{s}\mathds{1}(\lvert q^{N}(\omega)-1\rvert>0.5)\right]$
$\leq\mathbb{E}\left[\left(\max_{1\leq i\leq N}f(x_{i})\right)^{r}\right]^{s/r}\times\mathbb{P}[\lvert q^{N}(\omega)-1\rvert>0.5]^{1-s/r}$
$\leq q(f^{r})^{s/r}N^{s/r}\cdot C\cdot N^{-0.5p(1-s/r)},$
for a constant $C$. The last inequality uses the fact that $\mathbb{E}[(\max_{1\leq i\leq N}f(x_{i}))^{r}]\leq N\mathbb{E}[f(x_{1})^{r}]$, and Markov’s inequality using $q(\omega^{p})<\infty$. Given $s\leq pr/(p+r+2)$, the exponent of $N$ satisfies
$\frac{s}{r}-\frac{p(r-s)}{2r}=\frac{2s+ps-pr}{2r}\leq\frac{-s}{2},$
using $s\leq pr/(p+r+2)\Leftrightarrow-pr\leq-s(p+r+2)$. Altogether we arrive at
$\mathbb{E}\left[\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-q(f\omega)\right\rvert^{s}\mathds{1}(\lvert q^{N}(\omega)-1\rvert>0.5)\right]\leq CN^{-s/2},$
for another constant $C$.
In the case $r=\infty$, we can directly write
$\mathbb{E}\left[\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-q(f\omega)\right\rvert^{s}\mathds{1}(\lvert q^{N}(\omega)-1\rvert>0.5)\right]$ $\leq\mathbb{E}\left[\left(\max_{1\leq i\leq N}f(x_{i})+q(f\omega)\right)^{s}\mathds{1}(\lvert q^{N}(\omega)-1\rvert>0.5)\right]$
$\leq 2^{s}\lvert f\rvert_{\infty}^{s}\mathbb{P}[\lvert q^{N}(\omega)-1\rvert>0.5]$
$\leq 2^{s}C\lvert f\rvert_{\infty}^{s}N^{-0.5p}\leq CN^{-0.5s},$
using $s\leq\min\{p,r\}\leq p$ in the last line, and changing the value of $C$ between inequalities.
For the case $|q^{N}(\omega)-1|\leq 0.5$,
$\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-\pi(f)\right\rvert\mathds{1}(\lvert q^{N}(\omega)-1\rvert\leq 0.5)$ $\leq 2\lvert q^{N}(f\omega)-\pi(f)\rvert+\pi(f)\left\lvert\frac{q^{N}(\omega)-1}{q^{N}(\omega)}\right\rvert$
$\leq 2\lvert q^{N}(f\omega)-\pi(f)\rvert+2\pi(f)\left\lvert q^{N}(\omega)-1\right\rvert.$
Therefore
$\mathbb{E}\left[\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-\pi(f)\right\rvert^{s}\mathds{1}(\lvert q^{N}(\omega)-1\rvert<0.5)\right]$ $\leq C\left(\mathbb{E}[\lvert q^{N}(f\omega)-\pi(f)\rvert^{s}]+\mathbb{E}[\left\lvert q^{N}(\omega)-1\right\rvert^{s}]\right)$
$\leq CN^{-s/2},$
for some constant $C$ that changes at each line. The first term is $O(N^{-s/2})$ with a reasoning similar to that in the proof of Proposition 1.1, since $q^{N}(f\omega)$ is the sum of $N$ i.i.d. random variables with mean $q(f\omega)$ and $s$ finite moments, since $s\leq pr/(p+r+2)\leq pr/p+r$. Putting everything together gives
$\mathbb{E}\left[\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-\pi(f)\right\rvert^{s}\right]\leq CN^{-s/2}.$
A.3 Proofs of Section 3
We prove Theorem 3.1. We assume that both target and proposal distributions admit densities with respect to a measure $\lambda$. Although we will express all subsequent notations using integration, this should be interpreted as summation when the space is discrete and $\lambda$ represents the counting measure. The rejection probability at $\mathbf{x}$ is denoted by
$r(\mathbf{x})=\int_{\mathbf{z}\neq\mathbf{x}}\left(1-\min\left(1,\frac{\hat{Z}(\mathbf{z})}{\hat{Z}(\mathbf{x})}\right)\bar{q}(\mathbf{z})\right)\lambda(\mathrm{d}\mathbf{z}).$
That definition only considers the probability of moves to states different than $\mathbf{x}$ that are rejected. We will use the following fact: at every iteration, for each chain one of the following three events occurs: 1) a proposal to a different state is accepted, 2) a proposal to a different state is rejected, 3) a proposal is made to the current state (and systematically accepted). In a continuous state space with an atomless measure $\lambda$, the last event occurs with probability zero.
We first prove a lemma that describes the coupling time $\tau$.
Lemma A.1.
Assuming $\omega(\mathbf{x})\geq\omega(\mathbf{y})$, we have the following facts:
  • Let $\tau_{0}$ be the first time when the $\mathbf{x}$-chain moves to a different state. Then $\tau\leq\tau_{0}$, i.e. the chains meet at $\tau_{0}$ or earlier.
  • Let $\tau_{1}$ be the first time when a common proposal is $\mathbf{x}$. Then $\tau\leq\tau_{1}$, i.e. the chains meet at $\tau_{1}$ or earlier.
  • The meeting time satisfies $\tau=\min\{\tau_{0},\tau_{1}\}$.
Proof of Lemma A.1.
The first two observations can be proven by induction, once we recognize that the common draws coupling of Algorithm 4 implies $\omega(\mathbf{x}_{t})\geq\omega(\mathbf{y}_{t})$ for all $t\geq 0$. Regarding the last observation, for every $t<\min\{\tau_{0},\tau_{1}\}$, the $\mathbf{x}$-chain must have rejected moves to a different state than $\mathbf{x}$ at each iteration up to $t$. In that situation, the $\mathbf{x}$-chain is still at $\mathbf{x}$, while the $\mathbf{y}$-chain never proposed a move to $\mathbf{x}$ and thus $\mathbf{x}_{t}=\mathbf{x}\neq\mathbf{y}_{t}$, as claimed. ∎
Now we calculate the tail probability of $\tau$.
Lemma A.2.
For all $t\geq 1$, $|P^{t}(\mathbf{x},\cdot)-P^{t}(\mathbf{y},\cdot)|_{\text{TV}}\leq\mathbb{P}_{\mathbf{x},\mathbf{y}}(\tau>t)=\max(r(\mathbf{x}),r(\mathbf{y}))^{t}.$
Proof of Lemma A.2.
The inequality in the statement is the celebrated coupling inequality. For the equality, we assume $\omega(\mathbf{x})\geq\omega(\mathbf{y})$ without loss of generality, which implies $r(\mathbf{x})\geq r(\mathbf{y})$. By Lemma A.1, the event $\{\tau>t\}$ is equivalent to $\{\min\{\tau_{0},\tau_{1}\}>t\}$. The latter event corresponds to the event: “the $\mathbf{x}$-chain proposes to move to a different state but gets rejected at each of the first $t$ iterations”. Then its probability is $r(\mathbf{x})^{t}$, since $r(\mathbf{x})$ is the probability of a failed attempt to move to a different state. ∎
It remains to show the following lower bound.
Lemma A.3.
For all $t\geq 1$, $|P^{t}(\mathbf{x},\cdot)-P^{t}(\mathbf{y},\cdot)|_{\text{TV}}\geq\mathbb{P}_{\mathbf{x},\mathbf{y}}(\tau>t)$.
Proof of Lemma A.3.
Again, we assume $\omega(\mathbf{x})\geq\omega(\mathbf{y})$ without loss of generality. The definition of total variation distance as a supremum over measurable sets implies $|P^{t}(\mathbf{x},\cdot)-P^{t}(\mathbf{y},\cdot)|_{\text{TV}}\geq P^{t}(\mathbf{x},\{\mathbf{x}\})-P^{t}(\mathbf{y},\{\bf{x}\})$, considering the set $\{\mathbf{x}\}$.
Under the distribution of the coupled chains, we can write $P^{t}(\mathbf{x},\{\mathbf{x}\})-P^{t}(\mathbf{y},\{\bf{x}\})$ as $\mathbb{P}(\mathbf{x}_{t}=\mathbf{x})-\mathbb{P}(\mathbf{y}_{t}=\mathbf{x})$. Now we decompose each probability according to $\tau$ being greater or less than $t$, for any $t\geq 1$:
$\mathbb{P}\left(\mathbf{x}_{t}=\bf x\right)-\mathbb{P}\left(\mathbf{y}_{t}=\mathbf{x}\right)=\mathbb{P}\left(\mathbf{x}_{t}=\mathbf{x};\tau>t\right)+\mathbb{P}\left(\mathbf{x}_{t}=\mathbf{x};\tau\leq t\right)-\mathbb{P}\left(\mathbf{y}_{t}=\mathbf{x};\tau>t\right)-\mathbb{P}\left(\mathbf{y}_{t}=\mathbf{x};\tau\leq t\right).$
We simplify with the following observations.
  • Under the event $\tau>t$: we have $\mathbf{x}_{t}=\mathbf{x}$; otherwise, the $\mathbf{x}$-chain would have successfully moved to a new state jointly with the $\mathbf{y}$-chain implying $\tau\leq t$ by Lemma A.1. Therefore,
    $\mathbb{P}\left(\mathbf{x}_{t}=\mathbf{x};\tau>t\right)=\mathbb{P}\left(\tau>t\right)\mathbb{P}\left(\mathbf{x}_{t}=\mathbf{x}\mid\tau>t\right)=\mathbb{P}\left(\tau>t\right).$
    Meanwhile, under that event we have $\mathbf{y}_{t}\neq\mathbf{x}$; otherwise, the $\mathbf{y}$-chain must have proposed a move to $\mathbf{x}$ at or before time $t$, and that would have resulted in a meeting by Lemma A.1. Therefore,
    $\mathbb{P}\left(\mathbf{y}_{t}=\mathbf{x};\tau>t\right)=0.$
  • Under the event $\tau\leq t$: we have $\mathbf{x}_{t}=\mathbf{y}_{t}$, therefore $\mathbb{P}(\mathbf{x}_{t}=\mathbf{x};\tau\leq t)=\mathbb{P}(\mathbf{y}_{t}=\mathbf{x};\tau\leq t)$.
Putting these together, we conclude that $\mathbb{P}\left(\mathbf{x}_{t}=\mathbf{x}\right)-\mathbb{P}\left(\mathbf{y}_{t}=\mathbf{x}\right)=\mathbb{P}\left(\tau>t\right)$. ∎
Theorem 3.1 is obtained by combining Lemmas A.2 and A.3.
A.4 Proofs of Section 4
A.4.1 Proof of Proposition 4.1
Let $t\geq 1$. The event $\{\tau>t\}$ only occurs when Algorithm 5 enters its while loop, in which case we must have that 1) $\mathbf{x}_{1}=\mathbf{x}$, 2) $\hat{Z}(\mathbf{x})>\hat{Z}(\mathbf{y}_{0})$, and 3) the first generated Uniform variable was greater than $\hat{Z}(\mathbf{y}_{0})/\hat{Z}(\mathbf{x})$. Thus,
$\mathbb{P}(\tau>t)=\iint\mathbb{P}_{\mathbf{x},\mathbf{y}_{0}}\left(\tau>t\right)\left(1-\min\left\{1,\frac{\hat{Z}(\mathbf{y}_{0})}{\hat{Z}(\mathbf{x})}\right\}\right)\mathds{1}\left(\hat{Z}(\mathbf{y}_{0})<\hat{Z}(\mathbf{x})\right)\mathds{1}(\mathbf{y}_{0}\neq\mathbf{x})\bar{q}(\mathrm{d}\mathbf{x})\bar{q}(\mathrm{d}\mathbf{y}_{0}).$ (63)
The quantity $\mathbb{P}_{\mathbf{x},\mathbf{y}_{0}}(\tau>t)$ in the event $\hat{Z}(\mathbf{y}_{0})<\hat{Z}(\mathbf{x})$ is equal to $r(\mathbf{x})^{t}$, as in Theorem 3.1. By upper-bounding the other terms by one and integrating with respect to $\bar{q}(\mathrm{d}\mathbf{y}_{0})$, we obtain the upper bound
$\mathbb{P}(\tau>t)\leq\int(r(\mathbf{x}))^{t}\bar{q}(\mathrm{d}\mathbf{x})=\mathbb{E}_{\bar{q}}\left[(r(\mathbf{x}))^{t}\right].$ (64)
A.4.2 Proof of Proposition 4.2
We prove Proposition 4.2 by first splitting the expectation according to whether $\hat{Z}(\mathbf{x})$ is less than or greater than 2:
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\right]=\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}(\hat{Z}(\mathbf{x})\leq 2)\right]+\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}(\hat{Z}(\mathbf{x})>2)\right].$ (65)
We then proceed through a series of lemmas to bound each term. The following lemmas are used to handle the case when $\hat{Z}(\mathbf{x})>2$:
Lemma A.4.
Under Assumption 1 and $q(\omega^{p})<\infty$ for any $p>1$, the rejection probability (19) is upper bounded as follows, for any $\theta\in[0,1]$:
$r(\mathbf{x})\leq 1-\min\left\{1,\frac{\theta}{\hat{Z}(\mathbf{x})}\right\}c_{p}(\theta),\quad\text{with}\quad c_{p}(\theta)=\frac{(1-\theta)^{p/(p-1)}}{q(\omega^{p})^{1/(p-1)}}\in[0,1].$ (66)
Proof.
Let $\theta\in[0,1]$. We start with an $L^{p}$-version of Paley-Zygmund inequality, as in page 2705, equation (12) of Petrov (2007) with $r=1$. If $W$ is a non-negative random variable and $p>1$, then
$\mathbb{P}\left(W>\theta\mathbb{E}[W]\right)\geq\frac{(1-\theta)^{p/(p-1)}\left(\mathbb{E}[W]\right)^{p/(p-1)}}{\left(\mathbb{E}[W^{p}]\right)^{1/(p-1)}}.$ (67)
Indeed, for any $b>0$, Hölder’s inequality implies
$\mathbb{E}[W]$ $=\mathbb{E}[W\mathds{1}(W>b)]+\mathbb{E}[W\mathds{1}(W\leq b)]$
$\leq\mathbb{P}(W>b)^{(1-1/p)}\mathbb{E}[W^{p}]^{1/p}+b.$
Re-arranging with $b=\theta\mathbb{E}[W]$ implies (67). We apply this to $\hat{Z}(\mathbf{x})$, under Assumption 1:
$\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})>\theta\right)\geq\frac{(1-\theta)^{p/(p-1)}}{\left(\mathbb{E}_{\bar{q}}\left[\left(\hat{Z}(\mathbf{x})\right)^{p}\right]\right)^{1/(p-1)}}\geq\frac{(1-\theta)^{p/(p-1)}}{q(\omega^{p})^{1/(p-1)}}.$ (68)
The latter inequality comes from Jensen’s, since $z\mapsto z^{p}$ is convex since $p>1$:
$\mathbb{E}_{\bar{q}}\left[\left(\hat{Z}(\mathbf{x})\right)^{p}\right]\leq\mathbb{E}_{\bar{q}}\left[\frac{1}{N}\sum_{n=1}^{N}\omega(x_{n})^{p}\right]=q(\omega^{p}).$ (69)
Inequality (68) implies that
$\displaystyle\int\min\left\{1,\frac{\hat{Z}(\mathbf{x}^{\star})}{\hat{Z}(\mathbf{x})}\right\}\bar{q}(\mathrm{d}\mathbf{x}^{\star})$ =$=\int_{\{\mathbf{x}^{\star}:\hat{Z}(\mathbf{x}^{\star})\leq\theta\}}\min\left\{1,\frac{\hat{Z}(\mathbf{x}^{\star})}{\hat{Z}(\mathbf{x})}\right\}\bar{q}(\mathrm{d}\mathbf{x}^{\star})$
+$\quad+\int_{\{\mathbf{x}^{\star}:\hat{Z}(\mathbf{x}^{\star})>\theta\}}\min\left\{1,\frac{\hat{Z}(\mathbf{x}^{\star})}{\hat{Z}(\mathbf{x})}\right\}\bar{q}(\mathrm{d}\mathbf{x}^{\star})$
$\geq 0+\min\left\{1,\frac{\theta}{\hat{Z}(\mathbf{x})}\right\}\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x}^{\star})>\theta\right)$
$\geq\min\left\{1,\frac{\theta}{\hat{Z}(\mathbf{x})}\right\}\frac{(1-\theta)^{p/(p-1)}}{q(\omega^{p})^{1/(p-1)}}.$
This yields the desired result. ∎
Lemma A.5.
Under Assumptions 1-2, there exists a constant $C>0$ such that for all $t\geq 1$, $N\geq 1$,
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}(\hat{Z}(\mathbf{x})>2)\right]\leq\frac{C}{N^{p/2}t^{p}}.$ (70)
Proof.
We split the expectation into two parts:
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})>2\right)\right]$ $=\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})\in(2,1+t)\right)\right]+\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})\geq 1+t\right)\right].$ (71)
For $\{\hat{Z}(\mathbf{x})\geq 1+t\}$, we directly apply Lemma 1.1:
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}(\hat{Z}(\mathbf{x})\geq 1+t)\right]$ $\leq\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})\geq 1+t\right)$ (72)
$\leq\frac{M(p)}{N^{p/2}t^{p}}.$ (73)
For $\{\hat{Z}(\mathbf{x})\in(2,1+t)\}$, we use Lemma A.4 with $\theta=1/2$:
$r(\mathbf{x})$ $\leq 1-\frac{c_{p}(1/2)}{2\hat{Z}(\mathbf{x})}\leq\exp\left(-\frac{c_{p}(1/2)}{2\hat{Z}(\mathbf{x})}\right).$ (74)
Let $c=c_{p}(1/2)/4$. Then using the fact that $\hat{Z}(\mathbf{x})>2$ implies that $\hat{Z}(\mathbf{x})\leq 2(\hat{Z}(\mathbf{x})-1)$, we have:
$r(\mathbf{x})^{t}$ $\leq\exp\left(-\frac{2ct}{\hat{Z}(\mathbf{x})}\right)\leq\exp\left(-\frac{ct}{\hat{Z}(\mathbf{x})-1}\right).$ (75)
We introduce the sets $A_{k}=\left[t/(k+1),t/k\right]$ for $k\geq 1$, so that $\cup_{k=1}^{\infty}A_{k}=[0,t]$ which contains $[1,t]$. Using the result of Lemma 1.1, we obtain the bound:
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}(\hat{Z}(\mathbf{x})\in(2,1+t))\right]$ $\leq\mathbb{E}_{\bar{q}}\left[\exp\left(-\frac{ct}{\hat{Z}(\mathbf{x})-1}\right)\mathds{1}\left(\hat{Z}(\mathbf{x})-1\in(1,t)\right)\right]$ (76)
$\leq\sum_{k=1}^{\infty}\mathbb{E}_{\bar{q}}\left[\exp\left(-\frac{ct}{\hat{Z}(\mathbf{x})-1}\right)\mathds{1}\left(\hat{Z}(\mathbf{x})-1\in A_{k}\right)\right]$ (77)
$\leq\sum_{k=1}^{\infty}\exp(-ck)\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})\geq 1+t/(k+1)\right)$ (78)
$\leq\sum_{k=1}^{\infty}\exp(-ck)\frac{M(p)}{N^{p/2}}\left(\frac{k+1}{t}\right)^{p}.$ (79)
Let $S_{p}=\sum_{k=1}^{\infty}\exp(-ck)(k+1)^{p}$, which is finite. Then:
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})\in(2,1+t)\right)\right]$ $\leq\frac{M(p)S_{p}}{N^{p/2}t^{p}}.$ (80)
Combining the bounds for both parts, we get:
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})>2\right)\right]$ $\leq\frac{M(p)}{N^{p/2}t^{p}}+\frac{M(p)S_{p}}{N^{p/2}t^{p}}$ (81)
$\leq\frac{M(p)(1+S_{p})}{N^{p/2}t^{p}}.$ (82)
Setting the new constant $C:=M(p)\left(1+S_{p}\right)$ completes the proof. ∎
Now, we turn our attention to controlling the expectation when $\hat{Z}(\mathbf{x})\leq 2$.
Lemma A.6.
Fix $p\geq 2$ and let $\beta_{p}$ be defined as in (24). There exist constants $A_{p},B_{p}>0$, depending only on $p$ and $q(\omega^{p})$, such that for all $N\geq 1$, for all $t\geq 1$, the following holds:
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})\leq 2\right)\right]\leq\left[\frac{A_{p}}{N^{\frac{t\wedge p}{2}}}+\frac{B_{p}}{N^{p/2}}\right]\beta_{p}^{t}.$ (83)
Proof.
We abuse notation to write $r$ as a function of the value $z$ taken by $\hat{Z}(\mathbf{x})$, instead of a function of $\mathbf{x}$, in various places in this proof. First notice that $r(z)$ is increasing in $z$. We thus have that for $t\geq 1$ that
$\mathbb{E}_{\bar{q}}\left[r(\hat{Z}(\mathbf{x}))^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})\leq 2\right)\right]$
$\leq r(2)^{t-t\wedge p}\mathbb{E}_{\bar{q}}\left[r(\hat{Z}(\mathbf{x}))^{t\wedge p}\mathds{1}\left(\hat{Z}(\mathbf{x})\leq 2\right)\right].$
We first consider the second factor. We have for any $\alpha\in(0,1)$,
$\mathbb{E}_{\bar{q}}\left[r(\hat{Z}(\mathbf{x}))^{t\wedge p}\mathds{1}\left(\hat{Z}(\mathbf{x})\leq 2\right)\right]$
$\leq\mathbb{E}_{\bar{q}}\left[r(\hat{Z}(\mathbf{x}))^{t\wedge p}\mathds{1}\left(1-\alpha\leq\hat{Z}(\mathbf{x})\leq 2\right)\right]+r(2)^{t\wedge p}\bar{q}\left\{|\hat{Z}(\mathbf{x})-1|\geq\alpha\right\}.$ (84)
That is because $\{\hat{Z}(\mathbf{x})\leq 1-\alpha\}\subset\{|\hat{Z}(\mathbf{x})-1|\geq\alpha\}$, and $r(z)\leq r(2)$ for $z\leq 2$.
At this point, notice that by Lemma A.4 with $\theta=1/2$ we have that $r(2)\leq 1-c_{p}(1/2)/4=\beta_{p}$, where $\beta_{p}$ is defined in (24). Also notice that
$r(z)$ $=1-\int\min\left\{1,\frac{z^{\ast}}{z}\right\}\bar{q}(dz^{\ast})=\int_{z^{\ast}=0}^{\infty}\bar{q}(dz^{\ast})-\int_{z^{\ast}=0}^{z}\frac{z^{\ast}}{z}\bar{q}(dz^{\ast})-\int_{z^{\ast}=z}^{\infty}\bar{q}(dz^{\ast})$
$=\int_{z^{\ast}=0}^{z}\bar{q}(dz^{\ast})-\int_{z^{\ast}=0}^{z}\frac{z^{\ast}}{z}\bar{q}(dz^{\ast})=\int_{z^{\ast}=0}^{z}\left(\frac{z-z^{\ast}}{z}\right)\bar{q}(dz^{\ast}).$
Returning to our calculation regarding the first term in (84),
$\mathbb{E}_{\bar{q}}\left[r(\hat{Z}(\mathbf{x}))^{t\wedge p}\mathds{1}\left(1-\alpha\leq\hat{Z}(\mathbf{x})\leq 2\right)\right]$
$=\mathbb{E}_{\bar{q}}\left[\left(\frac{1}{\hat{Z}(\mathbf{x})}\int_{z^{\ast}=0}^{\hat{Z}(\mathbf{x})}(\hat{Z}(\mathbf{x})-z^{\ast})\bar{q}(dz^{\ast})\right)^{t\wedge p}\mathds{1}\left(1-\alpha\leq\hat{Z}(\mathbf{x})\leq 2\right)\right]$
$\leq\frac{\bar{q}\{[0,\hat{Z}(\mathbf{x})]\}^{t\wedge p}}{(1-\alpha)^{t\wedge p}}\mathbb{E}_{\bar{q}}\left[\left(\int_{z^{\ast}=0}^{\hat{Z}(\mathbf{x})}(\hat{Z}(\mathbf{x})-z^{\ast})\frac{\bar{q}(dz^{\ast})}{\bar{q}\{[0,\hat{Z}(\mathbf{x})]\}}\right)^{t\wedge p}\mathds{1}\left(1-\alpha\leq\hat{Z}(\mathbf{x})\leq 2\right)\right]$
[…]
$\leq\frac{1}{(1-\alpha)^{t\wedge p}}\mathbb{E}_{\bar{q}}\left[\int_{z^{\ast}=0}^{\hat{Z}(\mathbf{x})}|\hat{Z}(\mathbf{x})-z^{\ast}|^{t\wedge p}\bar{q}(dz^{\ast})\right]$
$\leq\frac{1}{(1-\alpha)^{t\wedge p}}\mathbb{E}_{(\mathbf{x},\mathbf{x}^{\prime})\sim\bar{q}\otimes\bar{q}}\left[|\hat{Z}(\mathbf{x})-\hat{Z}(\mathbf{x}^{\prime})|^{t\wedge p}\right]\leq\frac{1}{(1-\alpha)^{t\wedge p}}\frac{A_{p}\bar{q}(\omega^{p})}{N^{\frac{t\wedge p}{2}}},$
for a constant $A_{p}$ depending only on $p$. The first inequality comes from $\hat{Z}(\mathbf{x})^{-1}\leq(1-\alpha)^{-1}$ on the event of interest, the second inequality is from Jensen’s since the function $u\mapsto u^{t\wedge p}$ is convex, the third inequality is from $\bar{q}(A)\leq 1$ and the indicator being smaller than one, the fourth is obtained by completing the integral over all $z^{\ast}\in(0,\infty)$, and the last is from a reasoning similar to the proof of Proposition 1.1, or by direct application of Minkowski’s inequality and Proposition 1.1.
Overall, choosing $\alpha=1/2$ we have that
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t\wedge p}\mathds{1}\left(\hat{Z}(\mathbf{x})\leq 2\right)\right]$ $\leq 2^{t\wedge p}A_{p}\bar{q}(\omega^{p})N^{-t\wedge p/2}+r(2)^{t\wedge p}\bar{q}\left\{|\hat{Z}(\mathbf{x})-1|\geq\alpha\right\}$
$\leq 2^{t\wedge p}A_{p}\bar{q}(\omega^{p})N^{-t\wedge p/2}+\beta_{p}^{t\wedge p}C_{p}N^{-p/2}2^{p},$
using Markov’s inequality as in Lemma 1.1. Finally, multiplying by $r(2)^{t-t\wedge p}$ we obtain
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})\leq 2\right)\right]\leq\frac{\beta_{p}^{t}\beta_{p}^{-t\wedge p}2^{t\wedge p}A_{p}\bar{q}(\omega^{p})}{N^{\frac{t\wedge p}{2}}}+\frac{2^{p}C_{p}\beta_{p}^{t}}{N^{p/2}},$
and we note that, since $\beta_{p}\leq 1$, we have $\beta_{p}^{-t\wedge p}2^{t\wedge p}\leq\beta_{p}^{-p}2^{p}$, and thus we can define $A_{p}$ and $B_{p}$ to obtain Lemma A.6. ∎
Proof of Proposition 4.2.
We combine the bounds from Lemmas A.5 and A.6, and note that the two terms in the bound of Lemma A.6 can be bounded by $A_{p}\beta_{p}^{t}N^{-(t\wedge p)/2}$ for some constant $A_{p}$, which is not the same $A_{p}$ as in the statement of Lemma A.6. ∎
A.4.3 Proofs of Theorem 4.1 and Corollary 4.1
Proof of Theorem 4.1.
Under Assumption 1, the PIMH chain is $\bar{\pi}$-irreducible, and by construction it is aperiodic and $\bar{\pi}$-invariant, therefore $|\bar{q}P^{t}-\bar{\pi}|_{\mathrm{TV}}\to 0$ as $t\to\infty$ (Theorem 4 in Roberts & Rosenthal 2004). Thus, for any $t\geq 0$, by the triangle inequality,
$|\bar{q}P^{t}-\bar{\pi}|_{\mathrm{TV}}\leq\sum_{j=1}^{\infty}|\bar{q}P^{t+j}-\bar{q}P^{t+j-1}|_{\mathrm{TV}}.$ (85)
By the coupling representation of the TV distance, for any $t\geq 0$, $j\geq 1$,
$|\bar{q}P^{t+j}-\bar{q}P^{t+j-1}|_{\mathrm{TV}}\leq\mathbb{E}[\mathds{1}(\mathbf{x}_{t+j}\neq\mathbf{y}_{t+j-1})]=\mathbb{P}(\tau>t+j),$ (86)
where $(\mathbf{x}_{t})$ and $(\mathbf{y}_{t})$ are jointly generated by Algorithm 5. Under Assumption 2, Proposition 4.3 applies and thus the series $\sum_{j=1}^{\infty}\mathbb{P}(\tau>t+j)$ converges. Thus, by the dominated convergence theorem we may swap expectation and limit to write
$|\bar{q}P^{t}-\bar{\pi}|_{\mathrm{TV}}\leq\mathbb{E}\left[\sum_{j=1}^{\infty}\mathds{1}(\mathbf{x}_{t+j}\neq\mathbf{y}_{t+j-1})\right]=\mathbb{E}\left[\max(0,\tau-t-1)\right],$ (87)
for all $t\geq 0$. This is (27).
We may express the expectation of a non-negative variable as a series of survival probabilities:
$\mathbb{E}\left[\max\left(0,\tau-t-1\right)\right]=\sum_{s=1}^{\infty}\mathbb{P}\left(\max\left(0,\tau-t-1\right)\geq s\right).$
For any $t\geq 0,s\geq 1$, $\max(0,\tau-1-t)\geq s$ if and only if $\tau>s+t$. Under Assumption 2, Proposition 4.3 obtains
$\mathbb{P}\left(\tau>s+t\right)\leq CN^{-1/2}(s+t)^{-p}.$
The series $\sum_{s=1}^{\infty}(s+t)^{-p}$ can be bounded as follows:
$\sum_{s=1}^{\infty}(s+t)^{-p}$ $=\sum_{s=1+t}^{\infty}s^{-p}=(1+t)^{-p}+\sum_{s=t+2}^{\infty}s^{-p}$ (88)
[…]
$={(1+t)}^{-p}+\left[-\frac{x^{-p+1}}{p-1}\right]_{1+t}^{\infty}$ (90)
$={(1+t)}^{-p}+\frac{(1+t)^{-p+1}}{p-1}$ (91)
$=(1+t)^{-p+1}\left(\frac{1}{1+t}+\frac{1}{p-1}\right)$ (92)
$=(1+t)^{-p+1}\left(\frac{(1+t/p)p}{(1+t)(p-1)}\right)$ (93)
$\leq\frac{p}{(p-1)(1+t)^{p-1}},$ (94)
[…]
$|\bar{q}P^{t}-\bar{\pi}|_{\text{TV}}\leq\frac{Cp}{\sqrt{N}(p-1)(1+t)^{p-1}},$
which completes the proof. ∎
Proof of Corollary 4.1.
The proof starts with multiple applications of the triangle inequality, Theorem 3.1, $\max(a,b)\leq a+b$ for $a,b\geq 0$:
$\left|P^{t}(\mathbf{x},\cdot)-\bar{\pi}\right|_{\mathrm{TV}}$ $\leq\left|P^{t}(\mathbf{x},\cdot)-\bar{q}P^{t}\right|_{\mathrm{TV}}+\left|\bar{q}P^{t}-\bar{\pi}\right|_{\mathrm{TV}}$
$\leq\displaystyle\int\left|P^{t}(\mathbf{x},\cdot)-P^{t}(\mathbf{y},\cdot)\right|\bar{q}(\mathrm{d}\mathbf{y})+\left|\bar{q}P^{t}-\bar{\pi}\right|_{\mathrm{TV}}$
$=\displaystyle\int\max\left(r(\mathbf{x}),r(\mathbf{y})\right)^{t}\bar{q}(\mathrm{d}\mathbf{y})+\left|\bar{q}P^{t}-\bar{\pi}\right|_{\mathrm{TV}}$
$\leq\left(r(\mathbf{x})\right)^{t}+\mathbb{E}_{\bar{q}}[\left(r(\mathbf{y})\right)^{t}]+\left|\bar{q}P^{t}-\bar{\pi}\right|_{\mathrm{TV}}.$
The result then follows from Proposition 4.2 and Theorem 4.1. ∎
Proof of Proposition 4.4.
Similarly to the proof of Corollary 4.1, we start from
$|P^{t}(x,\cdot)-\pi|_{\mathrm{TV}}=|P^{t}(x,\cdot)-\pi P^{t}|_{\mathrm{TV}}\leq r(x)^{t}+\mathbb{E}_{y\sim\pi}[r(y)^{t}].$ (95)
Let $p>1$. We can apply Lemma A.4 to obtain
$r(y)\leq 1-\min\left\{1,\frac{\theta}{\omega(y)}\right\}\frac{(1-\theta)^{p/(p-1)}}{q(\omega^{p})^{1/(p-1)}},$ (96)
and we set $\theta=1/2$, and $c=\frac{(1-\theta)^{p/(p-1)}}{q(\omega^{p})^{1/(p-1)}}$. Note that $c\leq 1$ as $q(\omega^{p})\geq q(\omega)^{p}=1$. We next bound the expected rejection probability as follows
$\mathbb{E}_{y\sim\pi}\left[r(y)^{t}\right]$ $\leq\mathbb{E}_{y\sim\pi}\left[\left(1-\min\left\{\frac{0.5}{\omega(y)},1\right\}c\right)^{t}\right]$ (97)
$\leq\mathbb{E}_{y\sim\pi}[(1-0.5c)^{t}I(\omega(y)\leq 1)]+\mathbb{E}_{y\sim\pi}\left[\left(1-\frac{0.5}{\omega(y)}c\right)^{t}I(\omega(y)\in[1,t])\right]+\mathbb{P}[\omega(y)\geq t]$ (98)
$\leq(1-0.5c)^{t}+\mathbb{E}_{y\sim\pi}\left[\left(1-\frac{0.5}{\omega(y)}c\right)^{t}I(\omega(y)\in[1,t])\right]+\frac{\tilde{C}}{t^{p-1}}$ (99)
$\leq(1-0.5c)^{t}+\mathbb{E}_{y\sim\pi}\left[\exp\{-Ct/\omega(y)\}I(\omega(y)\in[1,t])\right]+\frac{\tilde{C}}{t^{p-1}}.$ (100)
The second inequality follows by splitting the weight into $\omega\leq 1,\omega\in[1,t]$ and $\omega>t$. The third inequality employs Markov’s inequality and the assumption that $p>1$. The last inequality uses $\log(1+x)\leq x$ with $x=-0.5c/\omega(y)$, $C=0.5c$. Consider the three terms on the last line. The first term decays exponentially fast with $t$, the third term decays at the rate of $t^{-(p-1)}$. It remains to bound the second term.
Define $A_{k}:=[t/(k+1),t/k]$, then clearly $\cup_{k=1}^{\infty}A_{k}=[0,t]$. We bound the second term as follows:
$\mathbb{E}_{y\sim\pi}\left[\exp\{-Ct/\omega(y)\}I(\omega(y)\in[1,t])\right]$ $\leq\mathbb{E}_{y\sim\pi}\left[\exp\{-Ct/\omega(y)\}I(\omega(y)\in[0,t])\right]$
$=\sum_{k=1}^{\infty}\mathbb{E}_{y\sim\pi}\left[\exp\{-Ct/\omega(y)\}I(\omega(y)\in A_{k})\right]$
$\leq\sum_{k=1}^{\infty}\exp\{-Ct/(t/k)\}\mathbb{P}[\omega(y)\geq t/(k+1)]$
$\leq\sum_{k=1}^{\infty}\exp\{-Ck\}\frac{C^{\prime}(k+1)^{p-1}}{t^{p-1}}$
$=\frac{C^{\prime\prime}}{t^{p-1}}\sum_{k=1}^{\infty}\exp\{-Ck\}(k+1)^{p-1}$
$\leq\frac{C^{\prime\prime\prime}}{t^{p-1}}.$
The last inequality holds as $\sum_{k=1}^{\infty}\exp\{-Ck\}(k+1)^{p-1}<\infty$ (the terms inside the summation decay exponentially fast). This concludes the proof. ∎
A.4.4 Proof of the result in Example 3
To complement the upper bound in Corollary 4.1, we present an example where $q(\omega^{p})<\infty$, and $\left|P^{t}(\mathbf{x}_{0},\cdot)-\bar{\pi}\right|_{\text{TV}}=\tilde{\Omega}(t^{-(p-1)})$ for some $\mathbf{x}_{0}$. Here $\Omega$ hides constants that may depend on $p$, and $\tilde{\Omega}$ indicates that we are disregarding polylogarithmic factors with respect to $t$. We set $N=1$ here as the focus is on the rate in $t$, and we revert to IMH notation for simplicity.
Let us consider $\pi(x):=Z_{\pi}x^{-p}$ on $[2,\infty)$, and $q(x):=Z_{q}\log^{2}(x)x^{-(p+1)}$ on $[2,\infty)$. In this case $\omega(x)=(Z_{\pi}/Z_{q})(x/\log^{2}(x))$. We can check:
$q(\omega^{p})=\pi(\omega^{p-1})=(Z_{\pi}/Z_{q})^{p}Z_{q}\int_{x=2}^{\infty}\frac{1}{\log(x)^{2(p-1)}x}\mathrm{d}x=(Z_{\pi}/Z_{q})^{p}Z_{q}\int_{\log 2}^{\infty}\frac{1}{t^{2(p-1)}}\mathrm{d}t<\infty,$
as $p\geq 2$.
Now we estimate $\mathbb{P}_{X\sim\pi}(X>s)$ and $\mathbb{P}_{X\sim q}(X>s)$ for any $s>2$ respectively. For the former:
$\mathbb{P}_{X\sim\pi}\left(X>s\right)=Z_{\pi}\int_{s}^{\infty}\frac{1}{x^{p}}\mathrm{d}x=\frac{C_{1}}{s^{p-1}}.$
For the latter,
$\mathbb{P}_{X\sim q}\left(X>s\right)$ $=Z_{q}\int_{s}^{\infty}\frac{\log^{2}(x)}{x^{p+1}}\mathrm{d}x=Z_{q}\int_{1}^{\infty}\frac{\log^{2}(su)}{s^{p+1}u^{p+1}}\cdot s\mathrm{d}u$
$\leq\frac{Z_{q}}{s^{p}}\int_{1}^{\infty}\frac{2\log^{2}(s)+2\log^{2}(u)}{u^{p+1}}\mathrm{d}u$
$\leq\frac{C_{2}\log^{2}(s)}{s^{p}}+\frac{C_{3}}{s^{p}}\leq\frac{C_{4}\log^{2}(s)}{s^{p}},$
where the first inequality follows from $(a+b)^{2}\leq 2a^{2}+2b^{2}$.
Consider an IMH chain $(X_{t})_{t\geq 0}$ targeting $\pi$ with proposal $q$ starting at $x_{0}=3$. Fix any $t\geq 100$, define $A_{t}:=(t(\log t)^{3},\infty)$. Then the probability of $A_{t}$ under $\pi$ is
$\mathbb{P}_{X\sim\pi}\left(X\in A_{t}\right)=\frac{C_{1}}{t^{p-1}(\log t)^{3(p-1)}}.$
On the other hand, $X_{t}$ is in $A_{t}$ implies at least one of the proposals made at times $1,2,\ldots,t$ falls into $A_{t}$ (note that $x_{0}\notin A_{t}$ since $100(\log(100))^{3}\approx 10^{4}$). By the union bound, we have
$\mathbb{P}\left(X_{t}\in A_{t}\right)\leq t\cdot\mathbb{P}_{Y\sim q}\left(Y\in A_{t}\right)\leq t\cdot\frac{C_{4}\log^{2}(t(\log t)^{3})}{t^{p}(\log t)^{3p}}\leq\frac{C_{4}\log^{2}(t^{2})}{t^{p-1}(\log t)^{3p}}=\frac{4C_{4}(\log t)^{2}}{t^{p-1}(\log t)^{3p}},$
where the last inequality uses $\log(t)^{3}\leq t$ when $t\geq 100$. Therefore, we have the following lower bound on the TV distance
$\left|P^{t}(x_{0},\cdot)-\pi\right|_{\text{TV}}$ $\geq\mathbb{P}_{X\sim\pi}\left(X\in A_{t}\right)-\mathbb{P}\left(X_{t}\in A_{t}\right)$
$\geq\frac{C_{1}(\log t)^{3}}{t^{p-1}(\log t)^{3p}}-\frac{4C_{4}(\log t)^{2}}{t^{p-1}(\log t)^{3p}}.$
Since $(\log t)^{2}=o((\log t)^{3})$ as $t\rightarrow\infty$, there exists $t_{0}=t_{0}(p)$ and $C_{5}>0$ such that for any $t>t_{0}$:
$\left|P^{t}(x_{0},\cdot)-\pi\right|_{\text{TV}}$ $\geq\frac{C_{5}}{t^{p-1}(\log t)^{3(p-1)}}=\tilde{\Omega}(t^{-(p-1)}).$
A.5 Proofs of Section 5
A.5.1 Proof of Proposition 5.1
Proof.
Note that $\hat{F}_{u}$ is not bounded even if $|f|_{\infty}\leq 1$, because the sum in (36) can be arbitrarily large. By Minkowski’s inequality, for any $s\geq 1$,
$\mathbb{E}\left[|\hat{F}_{u}|^{s}\right]^{1/s}\leq\mathbb{E}\left[|\hat{F}(\mathbf{x}_{0})|^{s}\right]^{1/s}+\mathbb{E}\left[\left|\sum_{t=1}^{\tau-1}\{\hat{F}(\mathbf{x}_{t})-\hat{F}(\mathbf{y}_{t-1})\}\right|^{s}\right]^{1/s}.$ (101)
Furthermore, if $|f|_{\infty}\leq 1$ then $|\hat{F}(\mathbf{x})|\leq 1$ for all $\mathbf{x}$, thus
$\mathbb{E}\left[|\hat{F}_{u}|^{s}\right]^{1/s}\leq\mathbb{E}\left[|\hat{F}(\mathbf{x}_{0})|^{s}\right]^{1/s}+2\mathbb{E}\left[\mathds{1}(\tau>1)\left|\tau-1\right|^{s}\right]^{1/s}.$ (102)
Since $\hat{F}(\mathbf{x}_{0})\leq 1$ almost surely, $\mathbb{E}[|\hat{F}(\mathbf{x}_{0})|^{s}]^{1/s}$ is finite for all $s\geq 1$. The latter expectation is smaller than $\mathbb{E}[|\tau|^{s}]^{1/s}$. Thus, $\hat{F}_{u}$ has $s$ finite moments if $\tau$ has $s$ finite moments. Note that $\hat{F}_{u}$ can have higher moments as well: for example if $f$ is constant then $\hat{F}_{u}$ is constant.
Next, in order for $\tau$ to have $s\geq 1$ moments, we can resort to Proposition 4.3. If Assumption 2 holds with $p>s$, then $\mathbb{P}(\tau>t)\leq CN^{-1/2}t^{-p}$. We can then follow the proof of Proposition 8 in Douc et al. (2024), using Tonelli’s theorem:
$\mathbb{E}\left[\tau^{s}\right]$ $=\mathbb{E}\left[\int_{0}^{\infty}\mathds{1}(u<\tau)su^{s-1}\mathrm{d}u\right]$
$=\int_{0}^{\infty}su^{s-1}\mathbb{P}(\tau>u)\mathrm{d}u$
$=\sum_{i=0}^{\infty}\mathbb{P}(\tau>i)\int_{i}^{i+1}su^{s-1}\mathrm{d}u$
$\leq\sum_{i=0}^{\infty}\mathbb{P}(\tau>i)s(i+1)^{s-1}.$
The sum is finite under the assumption $p>s$. ∎
A.5.2 Proof of Proposition 5.2
Proof of Proposition 5.2.
We now consider the PIMH chain $(\mathbf{x}_{t})_{t\geq 0}$, started from $\bar{q}$. The case $t=0$ corresponds to Theorem 2.3. Let $t\geq 1$. We can assume that $f$ is non-negative, using the same separate treatment of $f_{+}$ and $f_{-}$ as in the beginning of the proof of Theorem 2.3.
We write
$\hat{F}^{\circ}:\mathbf{x}\mapsto\hat{F}(\mathbf{x})-\pi(f)=\frac{\sum_{n=1}^{N}\omega(x_{n})\{f(x_{n})-\pi(f)\}}{\sum_{m=1}^{N}\omega(x_{m})}.$ (103)
We can write
$\mathbb{E}_{\mathbf{x}_{0}\sim\bar{q}}\left[|\hat{F}(\mathbf{x}_{t})-\pi(f)|^{s}\right]$ $=\int\bar{q}(\mathrm{d}\mathbf{x}_{0})P(\mathbf{x}_{0},\mathrm{d}\mathbf{x}_{1})\ldots P(\mathbf{x}_{t-1},\mathrm{d}\mathbf{x}_{t})|\hat{F}^{\circ}(\mathbf{x}_{t})|^{s}$
$=\int\bar{q}(\mathrm{d}\mathbf{x}_{0})P(\mathbf{x}_{0},\mathrm{d}\mathbf{x}_{1})\ldots P(\mathbf{x}_{t-1},\mathrm{d}\mathbf{x}_{t})|\hat{F}^{\circ}(\mathbf{x}_{t})|^{s}\{\mathds{1}(A_{t})+\mathds{1}(A_{t}^{c})\},$
where the event $A_{t}$ represents “there was an acceptance in the first $t$ steps”.
In the event $A_{t}^{c}$, $\mathbf{x}_{t}=\mathbf{x}_{0}$ so
$\int\bar{q}(\mathrm{d}\mathbf{x}_{0})P(\mathbf{x}_{0},\mathrm{d}\mathbf{x}_{1})\ldots P(\mathbf{x}_{t-1},\mathrm{d}\mathbf{x}_{t})|\hat{F}^{\circ}(\mathbf{x}_{t})|^{s}\cdot\mathds{1}(A_{t}^{c})$
$=\int\bar{q}(\mathrm{d}\mathbf{x}_{0})P(\mathbf{x}_{0},\mathrm{d}\mathbf{x}_{1})\ldots P(\mathbf{x}_{t-1},\mathrm{d}\mathbf{x}_{t})|\hat{F}^{\circ}(\mathbf{x}_{0})|^{s}\cdot\mathds{1}(A_{t}^{c})$
$\leq\int\bar{q}(\mathrm{d}\mathbf{x}_{0})|\hat{F}^{\circ}(\mathbf{x}_{0})|^{s},$
by bounding the indicator by one, and we can use Theorem 2.3 to obtain a bound in $N^{-s/2}$.
Now we consider the case $A_{t}$. For $1\leq j\leq t$ define the events
$A_{j,t}:=\{\mathbf{x}_{j-1}\neq\mathbf{x}_{j}=\mathbf{x}_{j+1}=\cdots=\mathbf{x}_{t}\},$
where $A_{j,t}$ is the event that there is a jump at time $j$ and no jump after that. Then $A_{j,t}\cap A_{j^{\prime},t}=\emptyset$ for $j\neq j^{\prime}$ and $A_{t}=\cup_{j=1}^{t}A_{j,t}$. We can decompose $\mathds{1}(A_{t})$ into $\sum_{j=1}^{t}\mathds{1}(A_{j,t})$ to get
$\int\bar{q}(\mathrm{d}\mathbf{x}_{0})P(\mathbf{x}_{0},\mathrm{d}\mathbf{x}_{1})\ldots P(\mathbf{x}_{t-1},\mathrm{d}\mathbf{x}_{t})|\hat{F}^{\circ}(\mathbf{x}_{t})|^{s}\mathds{1}(A_{t})$
$=\sum_{j=1}^{t}\mathbb{E}_{\mathbf{x}_{0}\sim\bar{q}}\left[|\hat{F}^{\circ}(\mathbf{x}_{t})|^{s}\mathds{1}(A_{j,t})\right]=\sum_{j=1}^{t}\mathbb{E}_{\mathbf{x}_{0}\sim\bar{q}}\left[\mathbb{E}\left\{\left.|\hat{F}^{\circ}(\mathbf{x}_{j})|^{s}\mathds{1}(A_{j,t})\right|\mathbf{x}_{j-1}\right\}\right].$
Conditional on $\mathbf{x}_{j-1}$,
$\int P(\mathbf{x}_{j-1},d\mathbf{x}_{j})P(\mathbf{x}_{j},d\mathbf{x}_{j+1})\cdots P(\mathbf{x}_{t-1},d\mathbf{x}_{t})|\hat{F}^{\circ}(\mathbf{x}_{j})|^{s}\mathds{1}\{\mathbf{x}_{j-1}\neq\mathbf{x}_{j}=\cdots=\mathbf{x}_{t}\}$
$=\int P(\mathbf{x}_{j-1},d\mathbf{x}_{j})|\hat{F}^{\circ}(\mathbf{x}_{j})|^{s}\mathds{1}\{\mathbf{x}_{j-1}\neq\mathbf{x}_{j}\}\int P(\mathbf{x}_{j},d\mathbf{x}_{j+1})\cdots P(\mathbf{x}_{t-1},d\mathbf{x}_{t})\mathds{1}\{\mathbf{x}_{j}=\cdots=\mathbf{x}_{t}\}$
$=\int P(\mathbf{x}_{j-1},d\mathbf{x}_{j})|\hat{F}^{\circ}(\mathbf{x}_{j})|^{s}\mathds{1}\{\mathbf{x}_{j-1}\neq\mathbf{x}_{j}\}r(\mathbf{x}_{j})^{t-j}$
$=\int\bar{q}(d\zeta)\alpha(\mathbf{x}_{j-1},\zeta)|\hat{F}^{\circ}(\zeta)|^{s}r(\zeta)^{t-j}.$
We can then upper bound $\alpha$ by one, and upper bound $\sum_{j=1}^{t}r(\zeta)^{t-j}$ by $(1-r(\zeta))^{-1}$ to obtain
$\sum_{j=1}^{t}\mathbb{E}_{\mathbf{x}_{0}\sim\bar{q},\zeta\sim\bar{q}}\left[\alpha(\mathbf{x}_{j-1},\zeta)|\hat{F}^{\circ}(\zeta)|^{s}r(\zeta)^{t-j}\right]$
$\leq\mathbb{E}_{\zeta\sim\bar{q}}\left[|\hat{F}^{\circ}(\zeta)|^{s}\sum_{j=1}^{t}r(\zeta)^{t-j}\right]$
$\leq\mathbb{E}_{\zeta\sim\bar{q}}\left[|\hat{F}^{\circ}(\zeta)|^{s}\frac{1}{1-r(\zeta)}\right].$
Next, split the expectation into the cases $\hat{Z}(\zeta)>2$ and $\hat{Z}(\zeta)\leq 2$:
$\mathbb{E}_{\zeta\sim\bar{q}}\left[|\hat{F}^{\circ}(\zeta)|^{s}\frac{1}{1-r(\zeta)}\right]$ $=\mathbb{E}_{\zeta\sim\bar{q}}\left[|\hat{F}^{\circ}(\zeta)|^{s}\frac{1}{1-r(\zeta)}\mathds{1}(\hat{Z}(\zeta)\leq 2)\right]+\mathbb{E}_{\zeta\sim\bar{q}}\left[|\hat{F}^{\circ}(\zeta)|^{s}\frac{1}{1-r(\zeta)}\mathds{1}(\hat{Z}(\zeta)>2)\right].$
When $\hat{Z}(\zeta)\leq 2$, since $r$ is increasing with $\hat{Z}$, we have $r(\zeta)\leq r(2)$ and thus $(1-r(\zeta))^{-1}\leq(1-r(2))^{-1}$. This yields:
$\mathbb{E}_{\zeta\sim\bar{q}}\left[|\hat{F}^{\circ}(\zeta)|^{s}\frac{1}{1-r(\zeta)}\mathds{1}(\hat{Z}(\zeta)\leq 2)\right]$ $\leq\frac{1}{1-r(2)}\int\bar{q}(\mathrm{d}\zeta)|\hat{F}^{\circ}(\zeta)|^{s}\mathds{1}(\hat{Z}(\zeta)\leq 2)$
$\leq\frac{1}{1-r(2)}\mathbb{E}_{\mathbf{x}_{0}\sim\bar{q}}\left[|\hat{F}^{\circ}(\mathbf{x}_{0})|^{s}\right].$
We obtain a bound in $N^{-s/2}$ using Theorem 2.3.
When $\hat{Z}(\zeta)>2$, from Lemma A.4, we have $r(\zeta)\leq 1-c_{p}(1/2)/(2\hat{Z}(\zeta))$. Thus $1-r(\zeta)\geq c_{p}(1/2)/(2\hat{Z}(\zeta))$, and $(1-r(\zeta))^{-1}\leq(2/c_{p}(1/2))\hat{Z}(\zeta)$. This yields:
$\mathbb{E}_{\zeta\sim\bar{q}}\left[|\hat{F}^{\circ}(\zeta)|^{s}\frac{1}{1-r(\zeta)}\mathds{1}(\hat{Z}(\zeta)>2)\right]\leq\frac{2}{c_{p}(1/2)}\int\bar{q}(\mathrm{d}\zeta)|\hat{F}^{\circ}(\zeta)|^{s}\hat{Z}(\zeta)\mathds{1}(\hat{Z}(\zeta)>2).$
Since we assume $f\geq 0$, we can use the inequality (62):
$|\hat{F}^{\circ}(\zeta)|^{s}\leq\left(\max_{1\leq i\leq N}f(\zeta_{i})+q(\omega f)\right)^{s},$
from which we obtain
$\mathbb{E}_{\zeta\sim\bar{q}}$ $\left[|\hat{F}^{\circ}(\zeta)|^{s}\frac{1}{1-r(\zeta)}\mathds{1}(\hat{Z}(\zeta)>2)\right]$
$\leq\frac{2}{c_{p}(1/2)}\mathbb{E}_{\zeta\sim\bar{q}}\left[\left(\max_{1\leq i\leq N}f(\zeta_{i})+q(\omega f)\right)^{s}\cdot\hat{Z}(\zeta)\mathds{1}(\hat{Z}(\zeta)>2)\right]$
$\leq\frac{2^{s}}{c_{p}(1/2)}\left(\mathbb{E}_{\zeta\sim\bar{q}}\left[\left(\max_{1\leq i\leq N}f(\zeta_{i})\right)^{s}\cdot\hat{Z}(\zeta)\mathds{1}(\hat{Z}(\zeta)>2)\right]+q(\omega f)^{s}\cdot\mathbb{E}_{\zeta\sim\bar{q}}\left[\hat{Z}(\zeta)\mathds{1}(\hat{Z}(\zeta)>2)\right]\right).$
Using the facts that $\hat{Z}(\zeta)\leq 2(\hat{Z}(\zeta)-1)$ when $\hat{Z}(\zeta)>2$ and $\mathds{1}(\hat{Z}(\zeta)\geq 2)\leq|\hat{Z}(\zeta)-1|^{p-1}$, we obtain via Proposition 1.1:
$q(\omega f)^{s}\cdot\mathbb{E}_{\zeta\sim\bar{q}}\left[\hat{Z}(\zeta)\mathds{1}(\hat{Z}(\zeta)>2)\right]$ $\leq 2q(\omega f)^{s}\mathbb{E}_{\zeta\sim\bar{q}}\left[|\hat{Z}(\zeta)-1|^{p}\right]$
$\leq\frac{M(p)}{N^{p/2}}q(\omega f)^{s}.$
For the remaining term, using Hölder’s inequality yields:
$\mathbb{E}_{\zeta\sim\bar{q}}$ $\left[\left(\max_{1\leq i\leq N}f(\zeta_{i})\right)^{s}\cdot\hat{Z}(\zeta)\mathds{1}(\hat{Z}(\zeta)>2)\right]$
$\leq 2\mathbb{E}_{\zeta\sim\bar{q}}\left[\left(\max_{1\leq i\leq N}f(\zeta_{i})\right)^{r}\right]^{s/r}\cdot\mathbb{E}_{\zeta\sim\bar{q}}\left[\left|\hat{Z}(\zeta)-1\right|^{\frac{r}{r-s}}\mathds{1}(\hat{Z}(\zeta)>2)\right]^{1-s/r}.$
Under the assumptions, with $s\leq\frac{pr}{p+r+2}$, we have:
$\frac{r}{r-s}\leq\frac{p+r+2}{r+2}=1+\frac{p}{r+2}\leq p,$
where the inequality holds since $r\geq 2$ by assumption. This gives us:
$\mathbb{E}_{\zeta\sim\bar{q}}\left[\left|\hat{Z}(\zeta)-1\right|^{\frac{r}{r-s}}\mathds{1}(\hat{Z}(\zeta)>2)\right]^{1-s/r}$ $\leq\left(\frac{M(p)}{N^{p/2}}\right)^{1-s/r}.$
Finally we use the fact that $\max\{a_{1},\ldots,a_{n}\}\leq a_{1}+\cdots+a_{n}$ for non-negative $a_{i}$ to derive
$\mathbb{E}_{\zeta\sim\bar{q}}\left[\left(\max_{1\leq i\leq N}f(\zeta_{i})\right)^{r}\right]$
$=\mathbb{E}_{\zeta\sim\bar{q}}\left[\max_{1\leq i\leq N}f(\zeta_{i})^{r}\right]\leq\mathbb{E}_{x\sim q}\left[f(x)^{r}N\right],$
so that
$\mathbb{E}_{\zeta\sim\bar{q}}\left[\left(\max_{1\leq i\leq N}f(\zeta_{i})\right)^{s}\cdot\hat{Z}(\zeta)\mathds{1}(\hat{Z}(\zeta)>2)\right]$
$\leq 2\left(q(f^{r})N\right)^{s/r}\cdot\left(\frac{M(p)}{N^{p/2}}\right)^{1-s/r}.$
We end up with an exponent of $N$ equal to $s/r-(p/2)(r-s)/r$, which under the assumptions is less than $-s/2$, as detailed in the proof of Theorem 2.3. Therefore, we obtain an upper bound in $N^{-s/2}$ on all terms.
Remark A.1.
Under Assumption 1, PIMH converges in total variation. Thus, $(\mathbf{x}_{t})$ converges weakly to $\bar{\pi}$. We consider the transformation $\mathbf{x}\mapsto|\hat{F}^{\circ}(\mathbf{x})|^{q}$ and Fatou’s lemma as in Theorem 3.4 of Billingsley (1999), to obtain
$\mathbb{E}_{\bar{\pi}}[|\hat{F}^{\circ}(\mathbf{x})|^{q}]\leq\lim\inf_{t}\mathbb{E}[|\hat{F}^{\circ}(\mathbf{x}_{t})|^{q}].$
Thus, the bound of Proposition 5.2, valid for all $t\geq 0$, applies also to the $s$-th moment of $\hat{F}(\mathbf{x})-\pi(f)$ under $\bar{\pi}$.
A.5.3 Proof of Proposition 5.3
Proof of Proposition 5.3.
We start as in the proof of Proposition 5.1 in Appendix A.5.1, and employ Theorem 2.3 for the moments of the error of IS with unbounded functions. Regarding the bias cancellation term,
BC $=\sum_{t=1}^{\infty}\Delta_{t}\mathds{1}(\tau>t),$ (104)
we use Minkowski with exponent $s\geq 1$:
$\mathbb{E}\left[|\text{BC}|^{s}\right]^{1/s}$ $\leq\sum_{t=1}^{\infty}\mathbb{E}\left[|\Delta_{t}|^{s}\mathds{1}(\tau>t)\right]^{1/s}.$ (105)
Next, for each time $t$, using Hölder’s inequality with an arbitrary $\kappa>1$,
$\mathbb{E}\left[|\Delta_{t}|^{s}\mathds{1}(\tau>t)\right]\leq\mathbb{E}\left[|\Delta_{t}|^{s\kappa}\right]^{1/\kappa}\mathbb{P}(\tau>t)^{(\kappa-1)/\kappa}.$ (106)
For the sum over $t$ in (105) to be finite, and using Proposition 4.3 to bound $\mathbb{P}(\tau>t)$, we have the condition on $\kappa$ and $s$,
$-\frac{p(\kappa-1)}{s\kappa}<-1\quad\Leftrightarrow\quad\kappa>p/(p-s).$
To establish the finiteness of $\mathbb{E}\left[|\Delta_{t}|^{s\kappa}\right]$ we can resort to Proposition 5.2 if $s\kappa$ satisfies the condition
$s\kappa\leq\frac{pr}{p+r+2}.$
We can find such $\kappa$ if
$\frac{ps}{p-s}<\frac{pr}{p+r+2}.$
A.5.4 Proof of Proposition 5.4
Proof of Proposition 5.4.
We follow the proof of Proposition 5.3, with $s=2$. We thus have a exponent $\kappa>1$ that must satisfy $\kappa>p/(p-2)$, and $2\kappa\leq pr/(p+r+2)$. We choose any number $\kappa$ strictly between $p/(p-2)$ and $pr/(2p+2r+4)$, which is possible by assumption, since
$1<\frac{p}{p-2}<\frac{pr}{2p+2r+4}\Leftrightarrow 2p+4r+4<rp.$
For that $\kappa$, we can apply Proposition 5.2 to bound $\mathbb{E}[|\Delta_{t}|^{2\kappa}]^{1/\kappa}$ by a constant times $N^{-1}$. Meanwhile, the sum $\sum_{t=1}^{\infty}\mathbb{P}(\tau>t)^{(\kappa-1)/(\kappa s)}$ is finite using Proposition 4.3, and is of the form $CN^{-a}$ for some positive $a$, namely $a=(\kappa-1)/(2\kappa s)$. Thus $\mathbb{E}{|\text{BC}|^{2}}$ can be bounded by a constant times $N^{-1-a}$ for some positive $a$, and finds itself negligible in front of the MSE of IS as $N\to\infty$. ∎

### v2

@@
 The Radon–Nikodym derivative of the target distribution relative to the proposal is called the weight function.
-Under the weak assumption that the weight is unbounded but has a number of finite moments under the proposal distribution, we obtain new results on the approximation error of importance sampling and of the particle independent Metropolis–Hastings algorithm (PIMH), which includes IMH as a special case.
-For IMH and PIMH, we show that the common random numbers coupling is maximal.
-Using that coupling we derive bounds on the total variation distance of a PIMH chain to the target distribution.
-The bounds are sharp with respect to the number of particles and the number of iterations.
-Our results allow a formal comparison of the finite-time biases of importance sampling and IMH.
-We further consider bias removal techniques using couplings of PIMH, and provide conditions under which the resulting unbiased estimators have finite moments.
-We compare the asymptotic efficiency of regular and unbiased importance sampling estimators as the number of particles goes to infinity.
+Under the assumption that the weight is unbounded but has finite moments under the proposal distribution, we study the approximation error of importance sampling and of the particle independent Metropolis–Hastings algorithm (PIMH), which includes IMH as a special case.
+For the chains generated by such algorithms, we show that the common random numbers coupling is maximal.
+Using that coupling we derive bounds on the total variation distance of a PIMH chain to its target distribution.
+Our results allow a formal comparison of the finite-time biases of importance sampling and IMH, and we find the latter to be have a smaller bias.
+We further consider bias removal techniques using couplings, and provide conditions under which the resulting unbiased estimators have finite moments.
+These unbiased estimators provide an alternative to self-normalized importance sampling, implementable in the same settings.
+We compare their asymptotic efficiency as the number of particles goes to infinity, and consider their use in robust mean estimation techniques.
 1 Introduction
 1.1 Context and contributions
+Two questions.
+Before recalling the context of Monte Carlo methods in which our work is situated, we consider the following two basic questions.
+  1. 1.
+    Let $\pi$ be a target distribution of interest on a measurable space $(\mathbb{X},\mathcal{X})$.
+The user can sample from a probability distribution $q$ on the same space, with $\pi$ absolutely continuous with respect to $q$, and evaluate the Radon–Nikodym derivative $\omega(x)=\pi(x)/q(x)$.
+Among $N$ independent draws $x_{1},\ldots,x_{N}$ from $q$, the user would like to select $x_{I}$ for some index $I\in[N]=\{1,\ldots,N\}$ such that $x_{I}$ is as close as possible to $\pi$ e.g.
+in total variation distance.
+What is the best selection strategy?
+  2. 2.
+    The user can sample i.i.d.
+pairs $(\omega_{n},f_{n})$ on $(\mathbb{R}_{+},\mathbb{R})$.
+Assume that the interest is in the limit of $\hat{I}_{N}=\sum_{n=1}^{N}\omega_{n}f_{n}/\sum_{n=1}^{N}\omega_{n}$ as $N\to\infty$, equal to $I=\mathbb{E}[\omega_{1}f_{1}]/\mathbb{E}[\omega_{1}]$ in $\mathbb{R}$.
+The ratio estimator $\hat{I}_{N}$ is biased for $I$: $\mathbb{E}[\hat{I}_{N}]\neq I$.
+Is it possible to generate, in finite time, an unbiased estimator of $I$?
+These questions relate to two building blocks of Monte Carlo methods, namely importance sampling (IS) and independent Metropolis–Rosenbluth–Teller–Hastings (IMH).
+We will see that, in some generality, IMH is preferable to IS to address the first question.
+Under some assumptions on $(\omega_{n},f_{n})$, we will see that a method that combines IS and IMH delivers an unbiased estimator of $I$, thus answering positively the second question.
 Monte Carlo with global proposals.
 Monte Carlo methods aim to approximate a target distribution $\pi$ on a measurable space $(\mathbb{X},\mathcal{X})$, for example $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$.
-These techniques are crucial when analytical computation of expectations under $\pi$ is infeasible due to high dimensionality or complex distribution forms.
+These techniques are crucial when analytical computation of expectations under $\pi$ is infeasible.
 The goal is to evaluate integrals of functions $f:\mathbb{X}\longrightarrow\mathbb{R}$ with respect to $\pi$:
 […]
-Two primary approaches are Markov Chain Monte Carlo (MCMC) methods, that construct a Markov chain with $\pi$ as its stationary distribution, and Importance Sampling (IS) methods, where the target distribution is approximated by weighted samples.
-Among MCMC methods, the independent Metropolis–Hastings (IMH) algorithm is a specialized form of the Metropolis–Rosenbluth–Teller–Hastings (MRTH) algorithm (Metropolis et al., 1953; Hastings, 1970), in which proposals are drawn from a distribution $q$ independently of the current state of the chain.
+Two primary approaches are Markov chain Monte Carlo (MCMC) methods, that construct a Markov chain with $\pi$ as its stationary distribution, and importance sampling methods, where the target distribution is approximated by weighted samples.
+Among MCMC methods, the IMH algorithm is a specialized form of the Metropolis–Rosenbluth–Teller–Hastings (MRTH) algorithm (Metropolis et al., 1953; Hastings, 1970), in which proposals are drawn from a distribution $q$ independently of the current state of the chain.
 The same proposal $q$ can be employed to generate draws in importance sampling, a procedure that can be traced back to Kahn (1949), as noted in Andral (2022).
-We assume that we can draw from $q$ and that its density, also denoted by $q$, can be evaluated pointwise.
 Therefore, IMH and IS propose two ways of correcting for the discrepancy between a proposal $q$ and a target $\pi$, and their comparison is a natural and fundamental question.
 Our contributions
-concern the performance of importance sampling and independent Metropolis–Hastings, and their comparison.
-Our key assumption is that the Radon–Nykodym derivative $\omega$ of $\pi$ with respect to $q$, termed the weight, has $p$ finite moments under $q$.
+concern the performance of importance sampling and independent Metropolis–Hastings.
+Our key assumption is that the Radon–Nikodym derivative $\omega$ of $\pi$ with respect to $q$, termed the weight, has $p$ finite moments under $q$.
 We first show that the bias of self-normalized importance sampling is of order $N^{-1}$, and we obtain new bounds on the moments of the error in importance sampling in Section 2.
 We then consider IMH, and show that the common random numbers coupling is optimal in Section 3.
-Using this coupling, in Section 4 we show that the total variation distance between IMH at iteration $t$ and $\pi$ decays as $t^{p-1}$.
+Using that coupling, in Section 4 we show that the total variation distance between IMH at iteration $t$ and $\pi$ decays as $t^{p-1}$.
 We obtain matching lower bounds in an example.
-We also obtain explicit dependencies in $N$ for the particle IMH algorithm (Andrieu et al., 2010).
+We further obtain explicit dependencies in $N$ for the particle IMH algorithm (Andrieu et al., 2010), a variant of IMH where $N$ proposals are sampled at each iteration.
+To establish precise bounds that account for both $t$ and $N$, we develop novel methods to analyze the average weight, $\hat{Z}=N^{-1}\sum_{n=1}^{N}\omega(x_{n})$, and thereby control its rejection probability.
+We first use the Paley–Zygmund inequality  (Petrov, 2007) (an anticoncentration inequality) to provide a lower bound on the probability of $\hat{Z}$ being small, specifically controlling its behavior when $\hat{Z}\leq 2$.
+For larger values of $\hat{Z}$, we divide the range into an intermediate section $(2,1+t)$ and a large section $[1+t,\infty)$.
+The Markov inequality directly handles the large range.
+For the intermediate range, we employ a peeling argument, breaking it down into a union of smaller intervals and applying the Markov inequality to each.
+Our multiscale analysis, combining both concentration and anticoncentration inequalities, allows us to control the $t$-th moment of the rejection probability.
+This eventually leads to a total-variation bound through the coupling inequality.
 In Section 5 we consider the bias removal technique of Glynn & Rhee (2014) applied by Middleton et al.
 (2019) to the particle IMH algorithm.
-This yields an unbiased estimator that can be implemented whenever self-normalized importance sampling or IMH can be implemented.
-We provide conditions under which the estimators have finite moments, and conditions under which their efficiency is asymptotically equivalent to that of importance sampling.
+This yields an unbiased estimator that can be implemented whenever self-normalized importance sampling, and thus relates to the second question above.
+We provide conditions under which these unbiased estimators have finite moments, and conditions under which their efficiency is asymptotically equivalent to that of importance sampling.
 1.2 Importance sampling
@@
 The asymptotic variance of IS is directly computed from the delta method (Owen, 2013; Robert & Casella, 2004; Liu, 2008), assuming $q(\omega^{2}\cdot f^{2})<\infty$ and $q(\omega^{2})<\infty$,
-$\lim_{n\to\infty}\mathbb{V}\left[\sqrt{N}(\hat{F}(x_{1},\ldots,x_{N})-\pi(f))\right]=q(\omega^{2}\cdot(f-\pi(f))^{2}).$ (5)
+$\lim_{n\to\infty}\mathbb{V}\left[\sqrt{N}(\hat{F}(x_{1},\ldots,x_{N})-\pi(f))\right]=q(\omega^{2}\cdot(f-\pi(f))^{2})=:\sigma^{2}_{\textsc{IS}}.$ (5)
 Agapiou et al.
@@
 1.3 Independent Metropolis–Hastings
-Independent Metropolis-Hastings (IMH) is an instance of the Metropolis–Rosenbluth–Teller–Hastings algorithm, where the draws are proposed from $q$ independently of the current state of the chain (Hastings, 1970, Section 2.5); see Algorithm 2.
-Thus IMH is implementable in the same settings as importance sampling.
+Independent Metropolis-Hastings (IMH) is an instance of the Metropolis–Rosenbluth–Teller–Hastings algorithm (Hastings, 1970, Section 2.5); described in Algorithm 2.
 Under Assumption 1, the IMH chain is $\pi$-irreducible, on top of being aperiodic and $\pi$-invariant by design, thus for $\pi$-almost every $x$, $|P^{t}(x,\cdot)-\pi|_{\mathrm{TV}}\to 0$ as $t\to\infty$ (Theorem 4 in Roberts & Rosenthal, 2004), where $P$ denotes the transition kernel of IMH, $P^{t}$ denotes the $t$-steps transition kernel, and $|\mu-\nu|_{\text{TV}}=\sup_{A\in\mathcal{X}}\mu(A)-\nu(A)$.
-The asymptotic variance of the ergodic average $t^{-1}\sum_{s=0}^{t-1}f(x_{t})$ generated by IMH is finite if $\pi(f^{2})<\infty$ and $q(\omega^{2}\cdot f^{2})$ (Theorem 2 in Deligiannidis & Lee, 2018), and under these conditions its variance is greater than or equal to the asymptotic variance of importance sampling in (5) (Proposition 2 in Deligiannidis & Lee, 2018).
-Thus, in terms of asymptotic variance, the comparison is clear: IS outperforms IMH.
+The asymptotic variance of the ergodic average $t^{-1}\sum_{s=0}^{t-1}f(x_{t})$ generated by IMH, denoted by $\sigma^{2}_{\textsc{IMH}}$, is finite if and only if $\pi(f^{2})<\infty$ and $q(\omega^{2}\cdot f^{2})<\infty$ (Theorem 2 in Deligiannidis & Lee, 2018).
+Furthermore, if $\sigma^{2}_{\textsc{IMH}}$ is finite then Deligiannidis & Lee (Proposition 2, 2018) provide a general comparison:
+$\sigma^{2}_{\textsc{IS}}\leq\sigma^{2}_{\textsc{IMH}},$ (6)
+where $\sigma^{2}_{\textsc{IS}}$ is the asymptotic variance of IS in (5).
+Thus, in terms of asymptotic variance, IS outperforms IMH.
 Since IMH defines a Markov transition, it can directly be used as a step within an encompassing Gibbs sampler (Skare et al., 2003), and it is commonly used within sequential Monte Carlo samplers (Chopin, 2002; South et al., 2019), and thus has its specific uses irrespective of the performance comparison with importance sampling.
@@
 Algorithm 2 IMH algorithm describing one step starting from $x$.
-When it comes to non-asymptotic behavior, for IMH there is an important distinction between two cases (Mengersen & Tweedie, 1996): either the weight is bounded, in which case the chain is geometrically ergodic and exact rates are obtained in Wang (2022), or the weight is unbounded and the convergence is polynomial at best; in the latter case, various results are provided e.g.
+When it comes to non-asymptotic behavior, for IMH there is an important distinction between two cases (Mengersen & Tweedie, 1996): either the weight is bounded, in which case the chain is geometrically ergodic and exact rates are obtained in Wang (2022), or the weight is unbounded and the convergence cannot be geometric; in the latter case, various results are provided e.g.
 in Jarner & Roberts (2002); Douc et al.
@@
 $\bar{\omega}(\mathbf{x})=\frac{\bar{\pi}(\mathbf{x})}{\bar{q}(\mathbf{x})}=\frac{1}{N}\sum_{n=1}^{N}\omega(x_{n})=\hat{Z}(\mathbf{x}).$ (10)
+Note that $\mathbb{E}_{\mathbf{x}\sim\bar{q}}[\bar{\omega}(\mathbf{x})]=q(\omega)=1$ under Assumption 1, thus $\bar{\pi}$ is properly normalized.
 Hence, IMH as in Algorithm 2, with proposal $\bar{q}$ and target $\bar{\pi}$, is equivalent to Algorithm 3.
@@
 Algorithm 3 PIMH algorithm describing one step starting from $\mathbf{x}=(x_{1},\ldots,x_{N})$.
-In order to estimate an expectation $\pi(f)$ from the PIMH output, a Rao–Blackwellization argument explained in Andrieu et al. (2010, Appendix B.5) leads to
-$\mathbb{E}_{\mathbf{x}\sim\bar{\pi}}\left[\hat{F}(\mathbf{x})\right]=\mathbb{E}_{\bar{\pi}}\left[\frac{\sum_{n=1}^{N}\omega(x_{n})f(x_{n})}{\sum_{n=1}^{N}\omega(x_{n})}\right]=\pi(f),$ (11)
-with $\hat{F}$ as in (4).
-For completeness we restate the arguments of Andrieu et al.
-(2010) justifying (11) in Appendix A.1.
-Thus, we can compute $\hat{F}(\mathbf{x})=\sum_{n=1}^{N}\omega(x_{n})f(x_{n})/\sum_{n=1}^{N}\omega(x_{n})$ at each iteration of the chain $(\mathbf{x}_{t})_{t\geq 0}$, and the ergodic average $T^{-1}\sum_{t=0}^{T-1}\hat{F}(\mathbf{x}_{t})$ may converge to $\pi(f)$.
-This allows connections between PIMH and IS, the latter being equal to $\hat{F}(\mathbf{x})$ with $\mathbf{x}\sim\bar{q}$.
+In order to estimate an expectation $\pi(f)$ from the PIMH output, note that
+[…]
+[…]
+[…]
+Thus, we can evaluate $\hat{F}:\mathbf{x}\mapsto\sum_{n=1}^{N}\omega(x_{n})f(x_{n})/\sum_{n=1}^{N}\omega(x_{n})$ at each state of the chain $(\mathbf{x}_{t})_{t\geq 0}$, and the ergodic average $T^{-1}\sum_{t=0}^{T-1}\hat{F}(\mathbf{x}_{t})$ may converge to $\pi(f)$.
+With this notation, the IS estimator (4) is $\hat{F}(\mathbf{x})$ with $\mathbf{x}\sim\bar{q}$.
 1.4 Moment conditions on the weight
@@
 $\mathbb{E}_{\bar{q}}[|\hat{Z}(\mathbf{x})-1|^{p}]$ $\leq\left(\frac{2^{1-1/p}(p-1)(1+q(\omega^{p}))^{1/p}}{\sqrt{N}}\right)^{p}=:M(p)N^{-p/2}.$ (13)
-We will repeatedly use the following consequence of Markov’s inequality and Proposition 1.1.
-Lemma 1.1.
-Under Assumptions 1-2, with $p\geq 2$, there exists a constant $M(p)>0$ such that, for any $z>0$ and $N\geq 1$:
-$\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})\geq 1+z\right)\leq\frac{M(p)}{N^{p/2}z^{p}}.$ (14)
+Furthermore, for any $t>0$ and $N\geq 1$:
+$\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})\geq 1+t\right)\leq\frac{M(p)}{N^{p/2}t^{p}}.$ (14)
 Remark 1.1.
@@
 Theorem 2.1.
-Let $f$ with $\pi(|f|)<\infty$.
 Assume that $x_{1},\ldots,x_{n}$ are i.i.d.
-from $q$, let $\omega:x\mapsto\pi(x)/q(x)$, and let $\hat{F}(\mathbf{x})=\sum_{n=1}^{N}\omega(x_{n})f(x_{n})/\sum_{m=1}^{N}\omega(x_{m})$.
-Assume that $q(|f|\cdot\omega^{2})<\infty$, $q(|f|\cdot\omega^{3})<\infty$ and $q(\omega^{-\eta})<\infty$ for some $\eta>0$.
+from $q$, let $\omega:x\mapsto\pi(x)/q(x)$, and let $\hat{F}(\mathbf{x})=\sum_{n=1}^{N}\omega(x_{n})f(x_{n})/\sum_{m=1}^{N}\omega(x_{m})$, for some test function $f$.
+Assume that $q(|f-\pi(f)|\omega)<\infty$, $q(|f-\pi(f)|\cdot\omega^{3})<\infty$ and $q(\omega^{-\eta})<\infty$ for some $\eta>0$.
 Then
@@
     i.e. our theorem holds with $s=2$ under their assumptions.
+Example 3 (Example 1 continued).
+We revisit the Exponential example to assess the asymptotic bias and variance of IS.
+For each value of $p$, we define the rate of the proposal as $k=p/(p-1)$ to ensure the existence of moments of $\omega$ under $q$ of order up to, but not including $p$.
+We consider the bounded test function $f(x)=\sin(x)$.
+The value of $\pi(f)$, the asymptotic bias (15) and the asymptotic variance (5) of IS can be computed analytically for any $k$, as detailed in Appendix B.
+Figure 1 shows how the biases and variances of IS, when rescaled by $N$, converge to the exact asymptotic values as $N$ increases, for $p=3$ and $p=5$.
+Figure 1: Left: asymptotic absolute bias of IS for different values of $N$ and $p$, with theoretical asymptotic bias (dashed lines).
+Right: asymptotic variance of IS for different values of $N$ and $p$, with theoretical asymptotic variance (dashed lines).
 3 Optimality of coupling IMH with common draws
@@
 Let $\tau=\inf\{t\geq 1:\mathbf{x}_{t}=\mathbf{y}_{t}\}$, and let $r(\mathbf{x})$ be defined as in (19).
-Then, for all $t\geq 1$,
+Then, under Assumption 1, for all $t\geq 1$,
 $|P^{t}(\mathbf{x},\cdot)-P^{t}(\mathbf{y},\cdot)|_{\text{TV}}=\mathbb{P}_{\mathbf{x},\mathbf{y}}(\tau>t)=\max(r(\mathbf{x}),r(\mathbf{y}))^{t}.$ (22)
@@
 The construction is described in Algorithm 5.
-The generated chains $(\mathbf{x}_{t})_{t\geq 0}$ and $(\mathbf{y}_{t})_{t\geq 0}$ have the same marginal distribution, that of an IMH chain started from $\bar{q}$.
+Note the redefinition of the meeting time $\tau$, which now corresponds to $\inf\{t\geq 1:\mathbf{x}_{t}=\mathbf{y}_{t-1}\}$ The generated chains $(\mathbf{x}_{t})_{t\geq 0}$ and $(\mathbf{y}_{t})_{t\geq 0}$ have the same marginal distribution, that of an IMH chain started from $\bar{q}$.
   1. 1.
@@
     1. (a)
+      Sample $(\mathbf{x}_{t+1},\mathbf{y}_{t})\sim\bar{P}((\mathbf{x}_{t},\mathbf{y}_{t-1}),\cdot)$, the common draws coupling of PIMH in Algorithm 4.
+    2. (b)
+      If $\mathbf{x}_{t+1}=\mathbf{y}_{t}$, set $\tau=t+1$.
+    3. (c)
       Set $t=t+1$.
-    2. (b)
-      Sample $(\mathbf{x}_{1+t},\mathbf{y}_{t})\sim\bar{P}((\mathbf{x}_{t},\mathbf{y}_{t-1}),\cdot)$, the common draws coupling of PIMH in Algorithm 4.
-    3. (c)
-      If $\mathbf{x}_{1+t}=\mathbf{y}_{t}$, set $\tau=1+t$.
   6. 6.
@@
 $\beta_{p}:=1-\frac{1}{2^{\frac{3p-2}{p-1}}q(\omega^{p})^{\frac{1}{p-1}}}.$ (24)
-Under Assumptions 1-2, there exist constants $A_{p},C_{p}>0$, depending only on $p$ and $q(\omega^{p})$, such that for all $N\geq 1$, for all $t\geq 1$, the following holds:
+Under Assumptions 1-2, $\beta_{p}\in(0,1)$ and there exist finite constants $A_{p},C_{p}>0$, depending only on $p$ and $q(\omega^{p})$, such that for all $N\geq 1$, for all $t\geq 1$, the following holds:
 $\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\right]\leq\frac{A_{p}}{N^{(t\wedge p)/2}}\beta_{p}^{t}+\frac{C_{p}}{t^{p}N^{p/2}}.$ (25)
@@
 Proposition 4.3.
-Consider $\tau$ generated by Algorithm 5. Under Assumptions 1-2, there exists $C$ such that for all $N\geq 1$ and all $t\geq 1$, if $p\geq 2$ in Assumption 2,
+Consider $\tau$ generated by Algorithm 5. Under Assumptions 1-2, there exists a finite $C>0$ such that for all $N\geq 1$ and all $t\geq 1$, if $p\geq 2$ in Assumption 2,
 $\mathbb{P}(\tau>t)\leq\frac{C}{\sqrt{N}t^{p}}.$ (26)
+As a consequence, we have $\mathbb{E}[\tau]\leq 1+C^{\prime}/\sqrt{N}$ with $C^{\prime}=C\sum_{t\geq 1}t^{-p}$.
 That bound retains the slowest rates in $N$ and $t$ from the previous result.
@@
 In turn this means that the rejection probability is at least $cN^{-1/2}$.
+Example 4 (Example 1 continued).
+We run coupled lagged PIMH chains (Algorithm 5) for different values of $N$ to generate $\tau$ and compute the empirical average.
+We obtain Figure 2.
+The results illustrate Proposition 4.3, which establishes that the scaled expected meeting $\sqrt{N}(\mathbb{E}[\tau]-1)$ become bounded as $N\to\infty$.
+The figure shows that the value of $N$ for which the asymptotic behaviour is reached is larger for smaller values of $p$.
+Figure 2: : Left: Average meeting time for different values of $N$ and $p$, where $q(\omega^{p-\epsilon})<\infty$ for all $\epsilon>0$ but $q(\omega^{p})=\infty$.
+Right: Average meeting time minus one scaled by $N^{1/2}$ for different values of $N$ and $p$.
+In the limit $N\to\infty$, the scaled meeting times should stabilise.
 4.2 Polynomial convergence rates
@@
 $\left|\bar{q}P^{t}-\bar{\pi}\right|_{\text{TV}}\leq\mathbb{E}\left[\max\left(0,\tau-1-t\right)\right].$ (27)
-Furthermore, if Assumption 2 also holds, there exists a constant $C$, independent of $t$ and $N$, such that for all $N\geq 1$ and $t\geq 0$,
+Furthermore, still under Assumption 2, there exists a constant $C$, independent of $t$ and $N$, such that for all $N\geq 1$ and $t\geq 0$,
 $\left|\bar{q}P^{t}-\bar{\pi}\right|_{\text{TV}}\leq\frac{C}{\sqrt{N}(1+t)^{p-1}}.$ (28)
@@
 where the first inequality is Cauchy–Schwarz and the second uses Proposition 1.1 under Assumption 2.
-Furthermore, in the large $N$ asymptotics we expect $1-\hat{Z}(\mathbf{x})$ to behave as a Normal distribution with mean zero and standard deviation $\sqrt{q(\omega^{2})-1}/\sqrt{N}$, so that its expectation should indeed behave as $\sqrt{(2/\pi)(q(\omega^{2})-1)}/\sqrt{N}$.
+Furthermore, in the large $N$ asymptotics we expect $1-\hat{Z}(\mathbf{x})$ to behave as a Normal distribution with mean zero and standard deviation $\sqrt{q(\omega^{2})-1}/\sqrt{N}$, so that the expectation of its absolute value should indeed behave as $\sqrt{(2/\pi)(q(\omega^{2})-1)}/\sqrt{N}$.
 We can also state a bound for the convergence of the chain started at any initial point $\mathbf{x}\in\mathbb{X}$.
@@
 Under Assumptions 1-2, there exists a constant $\tilde{C}$, independent of $t$ and $N$, such that for all $N\geq 1$, $t\geq 1$, and any starting point $\mathbf{x}\in\mathbb{X}^{N}$,
-$\left|P^{t}(\mathbf{x},\cdot)-\bar{\pi}\right|_{\text{TV}}\leq(r(\mathbf{x}))^{t}+\frac{\tilde{C}}{\sqrt{N}(1+t)^{p-1}}.$ (30)
+$\left|P^{t}(\mathbf{x},\cdot)-\bar{\pi}\right|_{\text{TV}}\leq(r(\mathbf{x}))^{t}+\frac{\tilde{C}}{\sqrt{N}\;t^{p-1}}.$ (30)
 Theorem 4.1 and Corollary 4.1 provide explicit bounds on the convergence rate of the PIMH algorithm.
@@
 Consider IMH under Assumption 1, and assume $q(\omega^{p})<\infty$ for $p>1$. There exists a constant $D$ independent of $t$ such that for all $t\geq 1$, and any starting point $x\in\mathbb{X}$,
-$\left|P^{t}(x,\cdot)-\pi\right|_{\text{TV}}\leq(r(x))^{t}+\frac{D}{(1+t)^{p-1}}.$ (31)
-The purpose of the following example is to demonstrate that the rate $t^{-(p-1)}$ in Corollary 4.1 and Proposition 4.4 cannot be improved beyond polylogarithmic factors.
+$\left|P^{t}(x,\cdot)-\pi\right|_{\text{TV}}\leq(r(x))^{t}+\frac{D}{t^{p-1}}.$ (31)
+Remark 4.3.
+The weight $\omega$ can be unbounded while having infinitely many moments under $q$, i.e.
+$q(\omega^{p})<\infty$ for all $p\geq 1$.
+For example, this happens when $\pi$ is Gamma$(2,1)$ and $q$ is Exponential$(1)$, leading to $\omega(x)=x$.
+In that case Mengersen & Tweedie (1996) prove that the IMH chain cannot be geometrically ergodic, while Proposition 4.4 holds with any $p>1$.
+Indeed the actual decay of $|P^{t}(x,\cdot)-\pi|_{\textsc{TV}}$ could be between geometric and polynomial in $t$, for example of the form $\exp(-t^{1/2})$.
+The purpose of the following example is to demonstrate that the rate $t^{-(p-1)}$ in Corollary 4.1 and Proposition 4.4 cannot be improved beyond polylogarithmic factors, without further assumptions.
 The proof is provided in Appendix A.4.4.
@@
 4.4 Comparing the biases of IS and IMH
-We can now compare the bias of IMH and IS under the assumption $q(\omega^{p})<\infty$ for $p\geq 2$.
-Suppose that the goal is to obtain a single sample close to $\pi$ in total variation distance, with a budget of $N$ samples from $q$ and $N$ evaluations of $\omega$.
-  • One approach is sampling-importance resampling (SIR), which refers to the following procedure.
-First, run Algorithm 1.
-Then, draw $k\sim\text{Categorical}(\omega(x_{1}),\ldots,\omega(x_{N}))$ and return $x_{k}$.
+In response to the first question in Section 1.1, we interpret the results on the bias of IMH and IS as follows.
+  • One approach to the first question is sampling-importance resampling (SIR), which refers to the following procedure.
+First, draw $x_{1},\ldots,x_{N}$ independently from $q$.
+Compute the normalized weights $\tilde{\omega}(x_{n})=\omega(x_{n})/\sum_{m=1}^{N}\omega(x_{m})$ for $n=1,\ldots,N$.
+Then, draw $k\sim\text{Categorical}(\tilde{\omega}(x_{1}),\ldots,\tilde{\omega}(x_{N}))$ and return $x_{k}$.
 For a test function $f$ with $|f|_{\infty}\leq 1$, under the conditions of Theorem 2.1 the marginal distribution $\mu_{N}^{\text{SIR}}$ of $x_{k}$ satifies
-    $\mu_{N}^{\text{SIR}}(f)-\pi(f)\sim_{N\to\infty}-q(\omega^{2}\cdot(f-\pi(f)))N^{-1}.$ (34)
-    Skare et al. (2003) provide a similar result in the case of bounded weights, and propose a modification of SIR to reduce this bias to $N^{-2}$.
-  • On the other hand, Theorem 4.1 suggests that IMH (with one particle) after $N$ iterations provides a sample from a distribution $qP^{N-1}$, for which, under the condition $q(\omega^{p})<\infty$,
+    $\mu_{N}^{\text{SIR}}(f)-\pi(f)\sim_{N\to\infty}-q(\omega^{2}\cdot(f-\pi(f)))N^{-1}$ (34)
+    Skare et al.
+(2003) obtains a similar result by considering the difference between the probability density function of the selected sample $x_{k}$ relative to the target $\pi$, but their approach requires stronger conditions on the weight function.
+They also propose a simple modification that results in a smaller bias in $N^{-2}$.
+  • On the other hand, Proposition 4.4 suggests that IMH, with one proposal per iteration, after $N$ iterations, provides a sample from a distribution $qP^{N-1}$, for which, under the condition $q(\omega^{p})<\infty$ with $p>1$, for a finite constant $C$,
     $\sup_{f:|f|_{\infty}\leq 1}\left\{qP^{N-1}(f)-\pi(f)\right\}\leq CN^{-(p-1)}.$ (35)
-    Thus, the bias may be reduced faster when increasing the computing budget in IMH rather than in SIR, as soon as $p>2$; for the modified SIR of Skare et al. (2003), IMH is faster as soon as $p>3$.
+Thus, the terminal sample after $N$ iterations of IMH is closer to $\pi$ in total variation than the sample obtained from SIR, as soon as Assumption 2 holds with $p>2$ and as $N\to\infty$.
+Compared to the modified SIR of Skare et al.
+(2003), IMH is still preferable as soon as Assumption 2 holds with $p>3$.
 In the case of bounded weights, MCMC methods such as PIMH or particle Gibbs (Andrieu et al., 2010) are geometrically ergodic (e.g.
@@
 (2022).
+This is in contrast to the comparison of asymptotic variances, which is at the advantage of IS as recalled in (6).
 5 Bias removal for self-normalized importance sampling
@@
 Here we consider the removal of the bias, and the associated cost.
-For this purpose, Middleton et al.
+There are multiple reasons to pursue bias removal for importance sampling.
+For example, gradient estimators in stochastic optimization, or estimators in the expectation step of the Expectation-Maximization (EM) algorithm, may be obtained by self-normalized importance sampling (e.g.
+Naesseth et al., 2020; Dhaka et al., 2021; Batardière et al., 2025).
+The resulting bias typically complicates the analysis of the convergence of encompassing optimization procedure.
+Another motivation for bias removal stems from the robust mean estimation literature, as described in Section 5.5.
+For bias removal, Middleton et al.
 (2019) employ common random numbers couplings of PIMH and the approach of Glynn & Rhee (2014).
-We pursue this strategy.
-Upon running Algorithm 5 with $N\geq 1$, with $\tau=\inf\{t\geq 1:\mathbf{x}_{t}=\mathbf{y}_{t-1}\}$, one can compute the following unbiased estimator:
+We pursue this strategy as well.
+Upon running Algorithm 5 with $N\geq 1$, with $\tau=\inf\{t\geq 1:\mathbf{x}_{t}=\mathbf{y}_{t-1}\}$, one can compute the following unbiased estimator, denoted by UIS:
 $\hat{F}_{u}=\hat{F}(\mathbf{x}_{0})+\sum_{t=1}^{\tau-1}\{\hat{F}(\mathbf{x}_{t})-\hat{F}(\mathbf{y}_{t-1})\},$ (36)
@@
 Remark 5.1.
-(36) is an instance of unbiased MCMC (Jacob et al., 2020; Atchadé & Jacob, 2024), where here MCMC is PIMH, and various generic improvements could be considered, such as increasing the lag between the chains, or introducing a burn-in parameter.
-However, in the particular case of PIMH, the number of particles $N$ is a key parameter and here we focus on the regime $N\to\infty$, in which case $\hat{F}_{u}$ naturally compares with $\hat{F}(\mathbf{x}_{0})$, which is the regular IS estimator.
-Hence we call (36) the unbiased self-normalized importance sampling estimator (UIS), but it could also be called unbiased PIMH.
+(36) is an instance of unbiased MCMC (Jacob et al., 2020; Atchadé & Jacob, 2024), and various generic improvements could be considered, such as increasing the lag between the chains, or introducing a burn-in parameter.
+However, in the particular case of PIMH, the number of particles $N$ is a key parameter and here we focus on the regime $N\to\infty$, in which case $\hat{F}_{u}$ naturally compares with $\hat{F}(\mathbf{x}_{0})$, the regular IS estimator.
+Hence we view (36) as unbiased self-normalized importance sampling (UIS).
 5.2 Moments of unbiased self-normalized importance sampling
@@
 Using Minkowski’s inequality, the moments of the error of $\hat{F}_{u}$ can be bounded by the moments of the error of the IS estimator $\hat{F}(\mathbf{x}_{0})$, as in Theorem 2.3, and the moments of BC.
-A first result is that, for bounded test functions $f$, $\hat{F}_{u}$ has as many moments as the meeting time $\tau$, which is nearly $p$ under Assumption 2. The proof is in Appendix A.5.1.
+A first result is that, for bounded test functions $f$, $\hat{F}_{u}$ has as many moments as the meeting time $\tau$, which is up to $p$ (non-included) under Assumption 2.
+The proof is in Appendix A.5.1.
 Proposition 5.1.
 Assume that $|f|_{\infty}\leq 1$ and let $s\geq 1$.
-If Assumptions 1-2 hold with $p\geq 2$ and $p>s$, then the meeting time $\tau$ has $s$ finite moments, and the unbiased self-normalized importance sampling (UIS) estimator $\hat{F}_{u}$ in (36) has $s$ finite moments.
+If Assumptions 1-2 hold with $p\geq 2$ and $p>s$, then the meeting time $\tau$ has $s$ finite moments, and the unbiased self-normalized importance sampling (UIS) estimator $\hat{F}_{u}$ in (36) has $s$ finite moments for any $N\geq 1$.
 To deal with unbounded test functions, we need to control the moments of the terms $\Delta_{t}$.
@@
 When $r=\infty$, the statement holds for $f$ such that $|f|_{\infty}<\infty$ and all $s<p$ such that $p>sp/(p-s)$.
-Two finite moments ($s=2$ in the above results) are particularly useful.
-Indeed, a confidence interval for $\pi(f)$ can be constructed from independent copies of $\hat{F}_{u}$ using the Central Limit Theorem as long as $\mathbb{V}[\hat{F}_{u}]<\infty$.
-Beyond the empirical average, robust mean estimation strategies could be envisioned.
-The typical assumption in robust mean estimation is that users have access to a random variable $X$ with mean $\mu$ of interest, and a finite variance $\sigma^{2}$ that may be unknown.
-The goal is to aggregate independent copies of $X$ into an estimator of $\mu$ with lighter tails than the empirical average.
-Such results can be found in Devroye et al.
-(2016); Lugosi & Mendelson (2019b, a); Lecué & Lerasle (2020); Minsker & Ndaoud (2021).
-These techniques are directly applicable with $X=\hat{F}_{u}$ and $\mu=\pi(f)$, whereas the application of robust mean estimation to self-normalized importance sampling is not a priori straightforward, and was explored in Dau (2022) in the case of median-of-means.
+Remark 5.2.
+In relation to the second question in Section 1.1, the construction of UIS in (36) can be generalized to the case where $(\omega_{n},f_{n})$ jointly follow a distribution $p$ on $\mathbb{R}_{+}\times\mathbb{R}$, rather than being deterministic functions $\omega$ and $f$ of a common random variable $x\sim q$.
+This answers the question positively, under moment conditions on $p$.
+These general results on moment bounds, particularly the case of finite second moments ($s=2$), motivate the exploration of robust estimation strategies detailed in Section 5.5.
 5.3 The asymptotic price of bias removal
-We consider the price of debiasing self-normalized importance sampling in terms of inefficiency, defined by the variance multiplied by the average cost.
+We consider the price of debiasing self-normalized importance sampling in terms of inefficiency, here defined by the mean squared error multiplied by the average cost (see e.g.
+Glynn & Whitt, 1992).
 We start by comparing the mean squared errors of unbiased and regular IS.
@@
 If we can bound $\mathbb{E}[\text{BC}^{2}]$ by a term that decreases faster than $N^{-1}$, then the MSE of UIS would be asymptotically equivalent to that of IS.
-Intuitively, the bias cancellation term goes to zero for two reasons: first because $\tau$ goes to one as $N\to\infty$, and the bias cancellation term equals zero in the event $\{\tau=1\}$.
-Secondly, each term $\hat{F}(\mathbf{x}_{t})-\hat{F}(\mathbf{y}_{t-1})$ goes to 00 as $N\to\infty$, under the conditions of Proposition 5.2.
+Intuitively, the bias cancellation term BC in (38) goes to zero for two reasons: first because $\tau$ goes to one as $N\to\infty$, and the bias cancellation term equals zero in the event $\{\tau=1\}$.
+Secondly, each term $\Delta_{t}=\hat{F}(\mathbf{x}_{t})-\hat{F}(\mathbf{y}_{t-1})$ goes to 00 as $N\to\infty$, under the conditions of Proposition 5.2.
 We obtain the following result, proven in Appendix A.5.4.
@@
 Then either $\tau=1$ or the algorithm enters its while loop up to the meeting time $\tau$, drawing $N$ new particles at each iterate of the loop.
-Counting the cost in terms of the number of evaluations of $\pi$, UIS has an overall cost of $\mathcal{C}=2N+N(\tau-1)$.
-If Assumption 2 holds with $p\geq 2$, using Proposition 4.3 then $\mathbb{E}[\tau]=\sum_{t\geq 0}\mathbb{P}(\tau>t)$ is finite, and furthermore, $\mathbb{E}[\mathcal{C}]\sim_{N\to\infty}2N$.
-We summarise these first observations in the next statement.
+Counting the cost in units of number of evaluations of $\pi$, UIS has an overall cost of $\mathcal{C}=2N+N(\tau-1)$.
+If Assumption 2 holds with $p\geq 2$, using Proposition 4.3 then $\mathbb{E}[\tau-1]\leq CN^{-1/2}$ for a finite constant $C$.
+Thus, as $N\to\infty$, $\mathbb{E}[\mathcal{C}]$ is equivalent to $2N$.
+We can then compare the asymptotic inefficiencies of UIS and IS as in the next statement.
 Proposition 5.5.
-The cost of the UIS estimator $\hat{F}_{u}$ in (36) is $\mathcal{C}=2N+N(\tau-1)$, and, under Assumption 2 for $p\geq 2$, there exists a constant $C$ such that $\mathbb{E}[\tau]\leq 1+C/\sqrt{N}$.
-Combining Proposition 5.4 with the above result, we see that under some conditions, the inefficiency $\mathbb{E}[\mathcal{C}]\mathbb{V}[\hat{F}_{u}]$ is equivalent to twice that of self-normalized importance sampling as $N\to\infty$.
-Indeed, the mean squared errors are equivalent but the cost of UIS behaves as $2N$ instead of $N$ for IS.
-In Algorithm 5 the $N$ particles in $\mathbf{y}_{0}$ are required to determine $\tau$ but in the event $\{\tau=1\}$ they do not participate directly in the estimator $\hat{F}_{u}$.
+Denote the cost of the UIS estimator $\hat{F}_{u}$ in (36) by $\mathcal{C}=2N+N(\tau-1)$.
+Under the conditions of Proposition 5.4, the inefficiency (expected cost multiplied by mean squared error) of UIS is twice that of IS as $N\to\infty$:
+$\lim_{N\to\infty}\mathbb{E}[\mathcal{C}]\cdot\mathbb{E}\left[(\hat{F}_{u}-\pi(f))^{2}\right]=2\times\lim_{N\to\infty}N\cdot\mathbb{E}_{\mathbf{x}\sim\bar{q}}\left[(\hat{F}(\mathbf{x})-\pi(f))^{2}\right].$
+The reason for the efficiency loss in UIS is that, in Algorithm 5 the $N$ particles in $\mathbf{y}_{0}$ are required to determine $\tau$ but in the event $\{\tau=1\}$, which is increasingly likely as $N\to\infty$, these $N$ particles do not participate directly in the estimator $\hat{F}_{u}$.
 5.4 An improved unbiased estimator
-A trick provides a remedy, and cuts the asymptotic inefficiency by a half.
-We can view $\hat{F}_{u}$ as a deterministic function of initial states $\mathbf{x}_{0}$ and $\mathbf{y}_{0}$ drawn from $\bar{q}$ independently, as well as extra variables (a sequence of proposals from $\bar{q}$, a sequence of uniform random variables) that we collectively label $\zeta$; we denote $\hat{F}_{u}$ by $\hat{F}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)$.
-Then we define the symmetrized UIS estimator:
-$\tilde{F}_{u}=\frac{1}{2}\left(\hat{F}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)+\hat{F}_{u}(\mathbf{y}_{0},\mathbf{x}_{0},\zeta)\right).$ (40)
+A simple trick provides a remedy, and cuts the asymptotic inefficiency by a half.
+We can view $\hat{F}_{u}$ as a deterministic function of initial states $\mathbf{x}_{0}$ and $\mathbf{y}_{0}$ drawn from $\bar{q}$ independently, as well as additional variables: an arbitrary long sequence of proposals from $\bar{q}$, and a sequence of uniform random variables used to accept or reject proposals, of the same length.
+Denote these two sequences by $\zeta$.
+Then $\hat{F}_{u}$ can be written as $\mathcal{A}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)$, where $\mathcal{A}_{u}$ is now a deterministic function.
+Then we define the Symmetrized UIS (SUIS) estimator:
+$\tilde{F}_{u}=\frac{1}{2}\left(\mathcal{A}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)+\mathcal{A}_{u}(\mathbf{y}_{0},\mathbf{x}_{0},\zeta)\right).$ (40)
 Computing (40) only requires simple modifications of Algorithm 5.
 Indeed, either $\hat{Z}(\mathbf{x}_{0})\geq\hat{Z}(\mathbf{y}_{0})$ or $\hat{Z}(\mathbf{x}_{0})<\hat{Z}(\mathbf{y}_{0})$.
-In the first case, we always have $\hat{F}_{u}(\mathbf{y}_{0},\mathbf{x}_{0},\zeta)=\hat{F}(\mathbf{y}_{0})$, and $\hat{F}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)$ can be computed following Algorithm 5 and (36).
-In the second case, we always have $\hat{F}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)=\hat{F}(\mathbf{x}_{0})$, and $\hat{F}_{u}(\mathbf{y}_{0},\mathbf{x}_{0},\zeta)$ can be computed following Algorithm 5 and (36) with the role of $\mathbf{x}_{0}$ and $\mathbf{y}_{0}$ swapped.
-That trick amounts to a Rao–Blackwellization over the arbitrary specification of which draws from $\bar{q}$ are used as $\mathbf{x}_{0}$ and as $\mathbf{y}_{0}$.
+In the first case, we always have $\mathcal{A}_{u}(\mathbf{y}_{0},\mathbf{x}_{0},\zeta)=\hat{F}(\mathbf{y}_{0})$, and $\mathcal{A}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)$ can be computed following Algorithm 5 and (36).
+In the second case, we always have $\mathcal{A}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)=\hat{F}(\mathbf{x}_{0})$, and $\mathcal{A}_{u}(\mathbf{y}_{0},\mathbf{x}_{0},\zeta)$ can be computed following Algorithm 5 and (36) with the role of $\mathbf{x}_{0}$ and $\mathbf{y}_{0}$ swapped.
+That trick amounts to a Rao–Blackwellization over the arbitrary specification of which draws from $\bar{q}$ are used as $\mathbf{x}_{0}$ or as $\mathbf{y}_{0}$.
 The following statement is a mild variation of the previous results and is stated without a proof.
 Proposition 5.6.
-Consider the symmetrized UIS estimator $\tilde{F}_{u}$ in (40).
+Consider the SUIS estimator $\tilde{F}_{u}$ in (40), with cost $\tilde{\mathcal{C}}$.
 Suppose that the conditions of Proposition 5.4 are satisfied.
-Then $\tilde{F}_{u}$ is an unbiased estimator of $\pi(f)$ for any $N\geq 1$, with finite expected cost and finite variance, and its inefficiency is equivalent to that of IS as $N\to\infty$.
-The result supports the intuition that $\tilde{F}_{u}$ should be preferred to $\hat{F}_{u}$ in practice, and that the asymptotic efficiency of $\tilde{F}_{u}$ matches that of self-normalized importance sampling as $N\to\infty$, under some conditions on $\omega$ and $f$ such as those in Proposition 5.4.
+Then $\tilde{F}_{u}$ is an unbiased estimator of $\pi(f)$ for any $N\geq 1$, with finite expected cost and finite variance, and its inefficiency is equivalent to that of IS as $N\to\infty$:
+$\lim_{N\to\infty}\mathbb{E}[\tilde{\mathcal{C}}]\cdot\mathbb{E}\left[(\tilde{F}_{u}-\pi(f))^{2}\right]=\lim_{N\to\infty}N\cdot\mathbb{E}_{\mathbf{x}\sim\bar{q}}\left[(\hat{F}(\mathbf{x})-\pi(f))^{2}\right].$
+The result supports the intuition that $\tilde{F}_{u}$ should be preferred to $\hat{F}_{u}$ in practice.
+Note however that the cost of the SUIS estimator is at least as large as that of the UIS estimator.
+Example 6 (Examples 1-3 continued).
+We perform experiments to assess the cost, variance and the inefficiency of the SUIS estimator with Exponential target and proposal distributions, and test function $f:x\mapsto\sin(x)$, as in Example 3.
+Figure 3 displays the variance of SUIS (left), and the product of variance times expected cost (right), as a function of $N$.
+For each value of $p$, as $N$ increases we observe that the variance resembles the asymptotic variance of IS divided by $2N$ represented with dashed lines (left).
+The inefficiency of SUIS converges to the asymptotic variance of IS represented by dashed lines (right).
+The results are based on $5\times 10^{6}$ independent repeats.
+Figure 3: : Left: Variance of the symmetrized unbiased importance sampling estimator (solid) vs.
+asymptotic variance $\sigma^{2}_{\textsc{IS}}$ divided by $2N$ (dashed) across $N$, for two different values of $p$.
+Right: Inefficiency (variance $\times$ cost) of SUIS (solid) vs.
+IS (dashed).
+5.5 Combining robust mean estimation with unbiased importance sampling
+5.5.1 Motivation
+An advantage of having eliminated the bias from IS through our SUIS approach is that we can now directly apply established robust mean estimation techniques (Lugosi & Mendelson, 2019).
+Their goal is to estimate the expectation $\mu$ of a random variable $X$ with finite variance $\sigma^{2}$ using $n$ i.i.d.
+copies $X_{1},\ldots,X_{n}$.
+For a given confidence level $\delta\in(0,1)$, we seek to construct an estimator $\hat{\mu}_{n}:=\hat{\mu}_{n}(X_{1},\ldots,X_{n},\delta)$ (potentially dependent on $\delta$) that satisfies, for the smallest possible value of $\epsilon=\epsilon(n,\delta)$, for all $n$:
+$\mathbb{P}(|\hat{\mu}_{n}-\mu|>\epsilon)\leq\delta.$ (41)
+If $\hat{\mu}_{n}$ is chosen to be the empirical average, using the Central Limit Theorem as $n\to\infty$, (41) holds asymptotically with $\epsilon=\sigma\sqrt{2\log(2/\delta)/n}$, using the inequality $\Phi^{-1}(1-\delta/2)\leq\sqrt{2\log(2/\delta)}$.
+On the other hand, the non-asymptotic bound obtained for the empirical mean with Chebyshev’s inequality gives $\epsilon=\sigma\sqrt{1/(n\delta)}$, which exhibits a poor dependence on $\delta$ that cannot be improved in general (see Section 2 in Lugosi & Mendelson, 2019).
+Remarkably, alternative, implementable estimators obtain a better dependence.
+The Median-of-Means (MoM, Nemirovskij & Yudin (1983)) estimator provides a prime example, where the $X_{1},\ldots,X_{n}$ are partitioned into $K$ blocks of size $m$ with $n=mK$, and the estimator is obtained as the median of the blockwise means.
+The following result shows that the MoM estimator achieves sub-Gaussian performance.
+Theorem 5.1 (Theorem $2$ in Lugosi & Mendelson (2019)).
+Let $X_{1},\ldots,X_{n}$ be i.i.d.
+random variables with mean $\mu\in\mathbb{R}$ and variance $\sigma^{2}\in(0,\infty)$.
+Let $\delta\in(0,1)$ and $K=\lceil 8\log(1/\delta)\rceil$ be a number of blocks, and assume that $n=mK$ for some positive integer $m$.
+Denote by $\hat{\mu}_{\textsc{MoM},n}$ the MoM estimator computed with $K$ blocks of size $m$.
+Then, with probability at most $\delta$, for all $n\geq\lceil 8\log(1/\delta)\rceil$,
+$|\hat{\mu}_{\textsc{MoM},n}-\mu|>\sigma\sqrt{\frac{32\log(1/\delta)}{n}}.$ (42)
+We provide a pseudo-code description of Median-of-Means (MoM), as well as the more efficient estimators of Minsker & Ndaoud (2021), hereafter Minsker–Ndaoud (MN), and Lee & Valiant (2022), hereafter Lee–Valiant (LV), in Appendix C.
+The combination of MoM with self-normalized importance sampling (SNIS) was explored by Dau (2022).
+The proposed estimator, termed MoM-SNIS, is the median of $K$ SNIS estimators obtained with $M$ draws each, so that the cost is equivalent to SNIS with $N=M\times K$ draws.
+The above result on MoM does not directly apply because SNIS is biased for the quantity of interest.
+Yet Dau (2022) shows the following result.
+Proposition 5.7 (Proposition 2 in Dau (2022)).
+Let $\delta\in(0,1)$ and suppose that $N\geq 8(32\sigma^{2}_{\omega}\vee 1)\log(1/\delta)$, where $\sigma^{2}_{\omega}=q(\omega^{2})-1$ is the variance of the weight, assumed to be finite.
+Assume also that $\sigma^{2}_{\textsc{IS}}=q(\omega^{2}(f-\pi(f))^{2})$, the asymptotic variance of SNIS as in (5), is finite.
+Then the MoM-SNIS estimator $\hat{F}_{\textsc{MoM-SNIS}}$ with $K=\lceil 8\log(1/\delta)\rceil$ satisfies, with probability at most $\delta$,
+$\left|\hat{F}_{\textsc{MoM-SNIS}}-\pi(f)\right|>\sigma_{\textsc{IS}}\sqrt{\frac{256\log(1/\delta)}{N}}$ (43)
+Thus MoM-SNIS achieves sub-Gaussian performance under minimal assumptions on $\omega$ and $f$, but requires a minimum number of particles $N$ to be larger than $8(32\sigma^{2}_{\omega}\vee 1)\log(1/\delta)$, where $\sigma^{2}_{\omega}$ is typically unknown to the user.
+As shown in Dau (2022, Proposition 3), it is not possible to obtain a similar result for MoM-SNIS that would hold for all $N$ larger than a threshold that would depend on $\delta$ only.
+The combination of MN and LV with IS has not yet be studied.
+In contrast, upon removing the bias of IS with SUIS, one can directly use MoM and obtain sub-Gaussian performance for all $n$ larger than $\lceil 8\log(1/\delta)\rceil$ (e.g.
+Theorem 5.1), where $n$ is a number of independent copies of SUIS.
+Indeed, using Proposition 5.1, for bounded test functions and under Assumption 2 with $p>2$, SUIS has a finite variance, for any choice of $N\geq 1$.
+We can similarly obtain guarantees from off-the-shelf results for MN (Minsker & Ndaoud, 2021) and LV (Lee & Valiant, 2022).
+Obvious disadvantages relative to MoM-SNIS (Dau, 2022) include the increased variance of SUIS relative to IS, and the random nature of the computing cost.
+5.5.2 Numerical experiments in the Exponential example
+We revisit the running example (Example 1-3-4), where $\pi$ is Exponential$(1)$ and $q$ is Exponential$(k)$.
+The weight $\omega$ has $p-\lambda$ moments for any $\lambda$ with $p=k/(k-1)$, and we consider $p\in\{2.01,2.1,3\}$.
+The cases represent various degrees of tail heaviness, with $p=2.01$ being the heaviest and $p=3$ the lightest.
+The test function $f(x)=\sin(x)$ is bounded so that SUIS has a finite variance in all cases.
+We first compare the performance of self-normalized importance sampling (SNIS), exactly normalized importance sampling (ENIS) with $N=1000$, and the MoM-SNIS method of Dau (2022), tuned with $\delta=0.05$, leading to $K=\lceil 8\log(1/\delta)\rceil=24$ and we choose $M=48$ so that $K\times M=1152$ is comparable to $N$.
+We obtain the results shown in Table 1.
+The columns describe the value of $p$, the method (ENIS, SNIS or MoM-SNIS), the cost per estimator, the mean squared error (MSE), the bias, and the $95\%$ and $99\%$ quantiles of the absolute error.
+The results are based on $10^{6}$ independent repeats of ENIS and SNIS, and $4\times 10^{4}$ independent repeats of MoM-SNIS.
+Importantly, Proposition 5.7 does not apply in the cases $p\in\{2.01,2.1\}$ because $N$ is not larger than $8(32\sigma^{2}_{\omega}\vee 1)\log(1/\delta)$ in these cases.
+Still, the results show an advantage of MoM-SNIS for $p\in\{2.01,2.1\}$, both in terms of MSE and large quantiles of the absolute error.
+This comes at the cost of a noticeable bias, compared to that of SNIS.
+Any advantage seems to disappear in the case $p=3$.
+Table 1: Performance of exactly normalized importance sampling (ENIS), self-normalized importance sampling (SNIS) and MoM-SNIS in the Exponential example.
+p method cost MSE bias q95 q99
+2.01 ENIS 1000 0.0023 0.0000 0.0721 0.1124
+2.01 SNIS 1000 0.0021 0.0014 0.0921 0.1450
+2.01 MoM-SNIS 1152 0.0016 0.0370 0.0634 0.0735
+2.10 ENIS 1000 0.0018 0.0000 0.0637 0.0964
+2.10 SNIS 1000 0.0016 0.0010 0.0768 0.1206
+2.10 MoM-SNIS 1152 0.0015 0.0349 0.0609 0.0708
+3.00 ENIS 1000 0.0004 0.0000 0.0379 0.0509
+3.00 SNIS 1000 0.0004 0.0001 0.0396 0.0528
+3.00 MoM-SNIS 1152 0.0006 0.0173 0.0456 0.0564
+We then consider Table 2 representing the performance of MoM-SUIS, LV-SUIS, MN-SUIS, with MoM-SNIS repeated from the previous table for comparison.
+All methods employ $K=\lceil 8\log(1/\delta)\rceil=24$ with $\delta=0.05$, and each SUIS estimator uses $N=4$ particles.
+The number of estimators per block is set to $M=4$ so that the cost of each method is approximately equivalent to SNIS with $N=1000$.
+Theorem 5.1 does apply for MoM-SUIS for all choices of $p$, and similarly the guarantees in Lee & Valiant (2022); Minsker & Ndaoud (2021) apply to LV-SUIS and MN-SUIS.
+However, the numerical results suggest that MoM-SNIS performs at least as well as the other methods, despite the fact that Proposition 5.7 does not apply in the cases $p\in\{2.01,2.1\}$.
+All the robust mean estimation methods under consideration lead to a noticeable bias, which is expected, and to an MSE than that of SNIS in the case $p=2.01$, which is more surprising.
+We expect the MSE of ENIS and SNIS to eventually be the smallest as the budget $N$ increases.
+Table 2: Performance of robust mean estimation techniques combined with SUIS in the Exponential example.
+p method cost MSE bias q95 q99
+2.01 MoM-SNIS 1152 0.0016 0.0370 0.0634 0.0735
+2.01 MoM-SUIS 959 0.0018 0.0352 0.0742 0.0910
+2.01 LV-SUIS 959 0.0018 0.0290 0.0717 0.0919
+2.01 MN-SUIS 959 0.0017 0.0368 0.0688 0.0826
+2.10 MoM-SNIS 1152 0.0015 0.0349 0.0609 0.0708
+2.10 MoM-SUIS 944 0.0017 0.0343 0.0717 0.0870
+2.10 LV-SUIS 944 0.0016 0.0260 0.0683 0.0873
+2.10 MN-SUIS 944 0.0017 0.0372 0.0678 0.0802
+3.00 MoM-SNIS 1152 0.0006 0.0173 0.0456 0.0564
+3.00 MoM-SUIS 868 0.0009 0.0219 0.0556 0.0684
+3.00 LV-SUIS 868 0.0006 0.0101 0.0473 0.0608
+3.00 MN-SUIS 868 0.0011 0.0293 0.0562 0.0666
+Finally, we visualise the quantiles of the absolute error for SNIS, MoM-SNIS, MoM-SUIS, LV-SUIS, and MN-SUIS, focusing on $p\in\{2.01,3\}$, in Figure 4.
+The algorithmic settings are identical to those described above.
+We are particularly interested in the highest quantile levels.
+In the case $p=2.01$, the most robust method appears to be MoM-SNIS followed by MN-SUIS, whereas in the case $p=3$, regular SNIS appears to perform best up the $99.9\%$ quantile.
+Thus robust mean estimation methods appear useful only in the hard cases with heavy-tailed importance weights.
+Figure 4: : quantiles of the absolute error $|\hat{F}-\pi(f)|$ for SNIS, MoM-SNIS, MoM-SUIS, MN-SUIS and LV-SUIS for $p=2.01$ (top) and $p=3$ (bottom).
 Acknowledgements.
 Through the second author and CY Initiative Emergence, this project was supported by France 2030.
+The fourth author acknowledges support from the National Science Foundation through grant DMS–2210849.
+The authors are thankful to Hai Dang Dau, Mohamed Ndaoud, Christian P.
+Robert and Yanbo Tang for helpful discussions.
 Appendix A Proofs
 A.1 Proofs of Section 1
-A.1.1 Proof of (11)
-We introduce the following extended state space justification of PIMH. Let the target and proposal distributions be
-$\tilde{\pi}(x_{1},\ldots,x_{N},k)$ $=\frac{1}{N}\pi(x_{k})\prod_{n\neq k}q(x_{n}),$ (41)
-$\tilde{q}(x_{1},\ldots,x_{N},k)$ $=\frac{\omega(x_{k})}{\sum_{n=1}^{N}\omega(x_{n})}\prod_{n=1}^{N}q(x_{n}).$ (42)
-The state space is $\mathbb{X}^{N}\times\{1,\ldots,N\}$, i.e. $k$ is an integer in $\{1,\ldots,N\}$. The ratio of target divided by proposal on the extended space takes the following form:
-$\tilde{\omega}(x_{1},\ldots,x_{N},k)=\frac{\tilde{\pi}(x_{1},\ldots,x_{N},k)}{\tilde{q}(x_{1},\ldots,x_{N},k)}=\frac{1}{N}\sum_{n=1}^{N}\omega(x_{n}).$ (43)
-In other words, the ratio simplifies to $\hat{Z}(\mathbf{x})$, thus IMH with proposal $\tilde{q}$ and target $\tilde{q}$ is equivalent to Algorithm 3.
-Draws from the proposal $\tilde{q}(x_{1},\ldots,x_{N},k)$ can be obtained by first sampling $x_{1},\ldots,x_{N}$ i.i.d.
-from $q$, then computing the importance weights $\omega(x_{n})$ for $n=1,\ldots,N$, and finally sampling $k$ from a Categorical distribution with probabilities given by $\omega(x_{n})/\sum_{m=1}^{M}\omega(x_{m})$.
-Under the extended target $\tilde{\pi}$, we have the following marginals and conditionals.
-  • The marginal distribution of $k$ is the Uniform distribution on $\{1,\ldots,N\}$.
-  • The marginal distribution of $x_{1},\ldots,x_{N}$ is obtained as
-    $\tilde{\pi}(x_{1},\ldots,x_{N})=\sum_{n=1}^{N}\frac{1}{N}\pi(x_{n})\prod_{m\neq n}q(x_{m}).$ (44)
-  • The conditional distribution of $k$ given $x_{1},\ldots,x_{N}$ can be obtained by dividing (41) by (44),
-    $\tilde{\pi}(k|x_{1},\ldots,x_{N})=\frac{\omega(x_{k})}{\sum_{n=1}^{N}\omega(x_{n})}.$ (45)
-  • The conditional distribution of $x_{1},\ldots,x_{N}$ given $k$ can be obtained by dividing (41) by $N^{-1}$, and is given by
-    $\tilde{\pi}(x_{1},\ldots,x_{N}|k)=\pi(x_{k})\prod_{n\neq k}q(x_{n}),$ (46)
-    and in particular $\tilde{\pi}(x_{k}|k)$ is $\pi(x_{k})$.
-Note that $f(x_{k})$ can be written as a more explicit function of $(x_{1},\ldots,x_{N},k)$, as $\sum_{n=1}^{N}f(x_{n})\mathds{1}(n=k)$.
-We use the above properties to obtain the following equalities, as in Andrieu et al.
-(2010), Appendix B.5, either conditioning on $k$ in the first line, or on $x_{1},\ldots,x_{N}$ in the second line:
-$\mathbb{E}_{\bar{\pi}}[\sum_{n=1}^{N}f(x_{n})\mathds{1}(n=k)]$ $=\mathbb{E}_{\bar{\pi}}[\mathbb{E}_{\bar{\pi}}[\sum_{n=1}^{N}f(x_{n})\mathds{1}(n=k)|k]]=\mathbb{E}_{\pi}[f],$ (47)
-$=\mathbb{E}_{\bar{\pi}}[\mathbb{E}_{\bar{\pi}}[\sum_{n=1}^{N}f(x_{n})\mathds{1}(n=k)|x_{1},\ldots,x_{N}]]=\mathbb{E}_{\bar{\pi}}\left[\frac{\sum_{n=1}^{N}\omega(x_{n})f(x_{n})}{\sum_{n=1}^{N}\omega(x_{n})}\right].$ (48)
-We obtain (11), i.e. the equality $\mathbb{E}_{\pi}[f]=\mathbb{E}_{\mathbf{x}\sim\bar{\pi}}[\hat{F}(\mathbf{x})]$, with $\hat{F}$ as in (4).
-A.1.2 Proofs of Proposition 1.1 and Lemma 1.1
 Proof of Proposition 1.1.
@@
 If we consider the average instead of the sum on the left, then the right-hand side becomes $m(p)N^{-p/2}\mathbb{E}[|X_{1}|^{p}]$.
-Since $q(\omega)=1$ and assuming that $q(\omega^{p})<\infty$, we define $X_{n}=\omega_{n}-1$ and apply the above result to obtain
+Since $q(\omega)=1$ and assuming that $q(\omega^{p})<\infty$, we define $X_{n}=\omega(x_{n})-1$ and apply the above result to obtain
 $\mathbb{E}\left[|\hat{Z}(\mathbf{x})-1|^{p}\right]\leq(p-1)^{p}N^{-p/2}q((\omega-1)^{p}).$
@@
 That inequality with $X=\omega$ and $Y=-1$ delivers $q((\omega-1)^{p})\leq 2^{p-1}(1+q(\omega^{p}))$. This establishes (13).
-For the non-centred moment, we proceed as follows:
+For the non-centred moment (12), we proceed as follows:
 $\mathbb{E}\left[|\hat{Z}(\mathbf{x})-1+1|^{p}\right]=\sum_{k=0}^{p}{p\choose k}\mathbb{E}\left[|\hat{Z}(\mathbf{x})-1|^{k}\right],$
@@
 $\mathbb{E}\left[|\hat{Z}(\mathbf{x})|^{p}\right]\leq\left(1+(p-1)N^{-1/2}2^{1-1/p}(1+q(\omega^{p}))^{1/p}\right)^{p}.$
-This bound goes to one as $N\to\infty$. ∎
-Proof of Lemma 1.1.
-Using Markov’s inequality and Proposition 1.1, we have:
+This bound gives (12), and goes to one as $N\to\infty$.
+To prove (14), using Markov’s inequality and (13), we have:
 $\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})\geq 1+z\right)$ $=\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})-1\geq z\right)$ (44)
@@
 $\leq\frac{M(p)N^{-p/2}}{z^{p}}.$ (46)
-The second line uses Markov’s inequality, and the third line applies Proposition 1.1. ∎
 A.2 Proofs of Section 2
@@
 Having different $x_{j}$’s in the numerator and denominator, and using their independence, the leading term in (53) is $-q(\omega^{2}\cdot(f-\pi(f)))\mathbb{E}[(\sum_{n=2}^{N}\omega(x_{n})/N)^{-2}]$.
+Note that $q(|f-\pi(f)|\omega^{2})<\infty$ under the assumptions of Theorem 2.1, since $\omega^{2}<\max(\omega,\omega^{3})$ and both $q(|f-\pi(f)|\omega)<\infty$ and $q(|f-\pi(f)|\cdot\omega^{3})<\infty$.
 Let $T_{N}=N^{-1}\sum_{n=2}^{N}\omega(x_{n})$.
@@
 $+2^{s-1}\left(q(f\omega)\right)^{s}\mathbb{P}[\lvert q^{N}(\omega)-1\rvert>0.5].$
-The second term leads to a bound in $N^{-s/2}$ using Markov’s inequality as in Lemma 1.1, since $q(\omega^{s})<\infty$ under the assumptions.
+The second term leads to a bound in $N^{-s/2}$ using Markov’s inequality as in Proposition 1.1, since $q(\omega^{s})<\infty$ under the assumptions.
 The first term is dealt with first using Hölder’s inequality with exponents $r/s$ and $(1-s/r)^{-1}$,
@@
 In a continuous state space with an atomless measure $\lambda$, the last event occurs with probability zero.
+We assume Assumption 1 throughout so that $\mathbb{P}(\hat{Z}(\mathbf{x})=0)=0$ under $q$, and the states $\mathbf{x}$, $\mathbf{y}$ in this section are such that $\hat{Z}(\mathbf{x})>0$, $\hat{Z}(\mathbf{y})>0$, otherwise $r(\mathbf{x}),r(\mathbf{y})$ would not be well-defined.
 We first prove a lemma that describes the coupling time $\tau$.
@@
 $\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\right]=\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}(\hat{Z}(\mathbf{x})\leq 2)\right]+\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}(\hat{Z}(\mathbf{x})>2)\right].$ (60)
-We then proceed through a series of lemmas to bound each term. The following lemmas are used to handle the case when $\hat{Z}(\mathbf{x})>2$:
+We then proceed through a series of lemmas to bound each term. The following lemmas are used to handle the case where $\hat{Z}(\mathbf{x})>2$:
 Lemma A.4.
@@
 Proof.
-Let $\theta\in[0,1]$. We start with an $L^{p}$-version of Paley-Zygmund inequality, as in page 2705, equation (12) of Petrov (2007) with $r=1$. If $W$ is a non-negative random variable and $p>1$, then
+Let $\theta\in[0,1]$.
+We start with a $L^{p}$-version of the Paley-Zygmund inequality, as on page 2705, equation (12) of Petrov (2007) with $r=1$.
+If $W$ is a non-negative random variable and $p>1$, then
 $\mathbb{P}\left(W>\theta\mathbb{E}[W]\right)\geq\frac{(1-\theta)^{p/(p-1)}\left(\mathbb{E}[W]\right)^{p/(p-1)}}{\left(\mathbb{E}[W^{p}]\right)^{1/(p-1)}}.$ (62)
@@
 $\leq\mathbb{P}(W>b)^{(1-1/p)}\mathbb{E}[W^{p}]^{1/p}+b.$
-Re-arranging with $b=\theta\mathbb{E}[W]$ implies (67). We apply this to $\hat{Z}(\mathbf{x})$, under Assumption 1:
+Rearranging with $b=\theta\mathbb{E}[W]$ implies (62). We apply this to $\hat{Z}(\mathbf{x})$, under Assumption 1:
 $\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})>\theta\right)\geq\frac{(1-\theta)^{p/(p-1)}}{\left(\mathbb{E}_{\bar{q}}\left[\left(\hat{Z}(\mathbf{x})\right)^{p}\right]\right)^{1/(p-1)}}\geq\frac{(1-\theta)^{p/(p-1)}}{q(\omega^{p})^{1/(p-1)}}.$ (63)
@@
 $\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})>2\right)\right]$ $=\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})\in(2,1+t)\right)\right]+\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})\geq 1+t\right)\right].$ (66)
-For $\{\hat{Z}(\mathbf{x})\geq 1+t\}$, we directly apply Lemma 1.1:
+For $\{\hat{Z}(\mathbf{x})\geq 1+t\}$, we directly apply Proposition 1.1:
 $\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}(\hat{Z}(\mathbf{x})\geq 1+t)\right]$ $\leq\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})\geq 1+t\right)$ (67)
@@
 $r(\mathbf{x})^{t}$ $\leq\exp\left(-\frac{2ct}{\hat{Z}(\mathbf{x})}\right)\leq\exp\left(-\frac{ct}{\hat{Z}(\mathbf{x})-1}\right).$ (70)
-We introduce the sets $A_{k}=\left[t/(k+1),t/k\right]$ for $k\geq 1$, so that $\cup_{k=1}^{\infty}A_{k}=[0,t]$ which contains $[1,t]$. Using the result of Lemma 1.1, we obtain the bound:
+We introduce the sets $A_{k}=\left[t/(k+1),t/k\right]$ for $k\geq 1$, so that $\cup_{k=1}^{\infty}A_{k}=[0,t]$ which contains $[1,t]$. Using the result of Proposition 1.1, we obtain the bound:
 $\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}(\hat{Z}(\mathbf{x})\in(2,1+t))\right]$ $\leq\mathbb{E}_{\bar{q}}\left[\exp\left(-\frac{ct}{\hat{Z}(\mathbf{x})-1}\right)\mathds{1}\left(\hat{Z}(\mathbf{x})-1\in(1,t)\right)\right]$ (71)
@@
 $\leq 2^{t\wedge p}A_{p}\bar{q}(\omega^{p})N^{-t\wedge p/2}+\beta_{p}^{t\wedge p}C_{p}N^{-p/2}2^{p},$
-using Markov’s inequality as in Lemma 1.1. Finally, multiplying by $r(2)^{t-t\wedge p}$ we obtain
+using Markov’s inequality as in Proposition 1.1. Finally, multiplying by $r(2)^{t-t\wedge p}$ we obtain
 $\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})\leq 2\right)\right]\leq\frac{\beta_{p}^{t}\beta_{p}^{-t\wedge p}2^{t\wedge p}A_{p}\bar{q}(\omega^{p})}{N^{\frac{t\wedge p}{2}}}+\frac{2^{p}C_{p}\beta_{p}^{t}}{N^{p/2}},$
@@
 Thus, $\hat{F}_{u}$ has $s$ finite moments if $\tau$ has $s$ finite moments.
-Note that $\hat{F}_{u}$ can have higher moments as well: for example if $f$ is constant then $\hat{F}_{u}$ is constant.
+Note that $\hat{F}_{u}$ can have higher moments as well: for example, if $f$ is constant, then $\hat{F}_{u}$ is constant.
 Next, in order for $\tau$ to have $s\geq 1$ moments, we can resort to Proposition 4.3.
@@
 Meanwhile, the sum $\sum_{t=1}^{\infty}\mathbb{P}(\tau>t)^{(\kappa-1)/(\kappa s)}$ is finite using Proposition 4.3, and is of the form $CN^{-a}$ for some positive $a$, namely $a=(\kappa-1)/(2\kappa s)$.
-Thus $\mathbb{E}{|\text{BC}|^{2}}$ can be bounded by a constant times $N^{-1-a}$ for some positive $a$, and finds itself negligible in front of the MSE of IS as $N\to\infty$.
+Thus, $\mathbb{E}{|\text{BC}|^{2}}$ can be bounded by a constant times $N^{-1-a}$ for some positive $a$, and finds itself negligible in front of the MSE of IS as $N\to\infty$.
 ∎
+Appendix B Calculations in the Exponential example
+Let $\pi$ be Exponential(1) and $q$ be Exponential($k$) as in Example 1. Then $\omega(x)=k^{-1}e^{-(1-k)x}$, and the $p$-th moment of $\omega$ under $q$ is given by
+$q(\omega^{p})$ $=\int_{0}^{\infty}k^{-p}e^{-\{p(1-k)+k\}x}\mathrm{d}x,$
+which is finite if and only if $p(1-k)+k>0$, or equivalently $p<{k}/(k-1)$. Let $f:x\mapsto\sin(x)$. We compute the following quantities:
+  • Integral of interest: $I=\int f(x)\pi(x)\mathrm{d}x=\int_{0}^{\infty}\sin(x)e^{-x}\mathrm{d}x$.
+  • Asymptotic bias of IS: $\mathcal{B}=-\int(f(x)-\pi(f))\omega^{2}(x)q(x)\mathrm{d}x$.
+  • Asymptotic variance of IS: $\mathcal{V}=\int(f(x)-\pi(f))^{2}\omega^{2}(x)q(x)\mathrm{d}x$.
+B.1 Integral of interest
+We have, using integration by parts twice, each time differentiating the trigonometric function and integrating the exponential function:
+$I$ $=\int_{0}^{\infty}\sin(x)e^{-x}\mathrm{d}x$
+$=\left[-\sin(x)e^{-x}\right]_{0}^{\infty}+\int_{0}^{\infty}e^{-x}\cos(x)\mathrm{d}x$
+$=0+\left[-\cos(x)e^{-x}\right]_{0}^{\infty}-\int_{0}^{\infty}e^{-x}\sin(x)\mathrm{d}x$
+$=1-\int_{0}^{\infty}e^{-x}\sin(x)\mathrm{d}x=1-I.$
+From this we obtain $I=\frac{1}{2}$.
+B.2 Asymptotic bias and variance of IS
+We notice that we can obtain both the asymptotic bias and variance of IS from the following integrals:
+$C_{1}$ $=\int\omega^{2}(x)q(x)\mathrm{d}x=\int_{0}^{\infty}k^{-1}\exp(-(2-k)x)\mathrm{d}x=\frac{1}{k(2-k)},$
+$C_{2}$ $=\int f(x)\omega^{2}(x)q(x)\mathrm{d}x=\int_{0}^{\infty}\sin(x)k^{-1}\exp(-(2-k)x)\mathrm{d}x,$
+$C_{3}$ $=\int f(x)^{2}\omega^{2}(x)q(x)\mathrm{d}x=\int_{0}^{\infty}\sin(x)^{2}k^{-1}\exp(-(2-k)x)\mathrm{d}x$
+$=\int_{0}^{\infty}\frac{1}{2}(1-\cos(2x))k^{-1}\exp(-(2-k)x)\mathrm{d}x$
+$=\frac{1}{2}C_{1}-\frac{1}{2}\int_{0}^{\infty}\cos(2x)k^{-1}\exp(-(2-k)x)\mathrm{d}x.$
+Then we can compute $C_{2}$ and the integral in $C_{3}$ using two steps of integration by parts, just as we did for the integral of interest $I$. We obtain:
+$k\times C_{2}$ $=\int_{0}^{\infty}\sin(x)\exp(-(2-k)x)\mathrm{d}x$
+$=\left[\frac{\sin(x)\exp(-(2-k)x)}{-(2-k)}\right]_{0}^{\infty}+\int_{0}^{\infty}\frac{\cos(x)\exp(-(2-k)x)}{2-k}\mathrm{d}x$
+$=\left[0+\left[\frac{\cos(x)\exp(-(2-k)x)}{-(2-k)^{2}}\right]_{0}^{\infty}-\frac{1}{(2-k)^{2}}\int_{0}^{\infty}\exp(-(2-k)x)\sin(x)\mathrm{d}x\right]$
+$=\frac{1}{(2-k)^{2}}-\frac{1}{(2-k)^{2}}(k\times C_{2}),$
+so that
+$C_{2}=\frac{1}{k(1+(2-k)^{2})}.$
+Similarly, we have:
+$\int_{0}^{\infty}\cos(2x)\exp(-(2-k)x)\mathrm{d}x$
+$=\left[\frac{\cos(2x)\exp(-(2-k)x)}{-(2-k)}\right]_{0}^{\infty}-\int_{0}^{\infty}\frac{-2\sin(2x)\exp(-(2-k)x)}{-(2-k)}\mathrm{d}x$
+$=\frac{1}{2-k}+\left[\frac{2\sin(2x)\exp(-(2-k)x)}{(2-k)^{2}}\right]_{0}^{\infty}-\frac{4}{(2-k)^{2}}\int_{0}^{\infty}\cos(2x)\exp(-(2-k)x)\mathrm{d}x,$
+so that
+$\int_{0}^{\infty}\cos(2x)\exp(-(2-k)x)\mathrm{d}x=\frac{1}{2-k}\times\frac{1}{1+4(2-k)^{-2}}=\frac{2-k}{4+(2-k)^{2}},$
+and thus
+$C_{3}=\frac{1}{2}\times\frac{1}{k(2-k)}-\frac{1}{2}\times\frac{2-k}{k(4+(2-k)^{2})}.$
+We put everything together with:
+$\mathcal{B}$ =$=-\left\{C_{2}-I\times C_{1}\right\},$
+$\mathcal{V}$ $=C_{3}-2\times I\times C_{2}+I^{2}\times C_{1}.$
+Appendix C Robust mean estimation
+This appendix provides algorithmic descriptions of three robust mean estimators, implemented in the experiments of Section 5.5.
+C.1 Median-of-Means (MoM)
+The MoM estimator is described in Algorithm 6.
+  1. 1.
+    Input: i.i.d. samples $X_{1},\ldots,X_{n}$ with mean $\mu$ and variance $\sigma^{2}$, confidence parameter $\delta\in(0,1)$.
+  2. 2.
+    Set $K=\lceil 8\log(1/\delta)\rceil$.
+  3. 3.
+    Partition $[n]=\{1,\ldots,n\}$ into $K$ blocks $B_{1},\ldots,B_{K}$, with each $B_{k}$ of size $\lfloor n/K\rfloor\leq|B_{k}|\leq\lfloor n/K\rfloor+1$.
+  4. 4.
+    For $j\in[K]$, compute block empirical mean
+    $\bar{X}_{j}=\frac{1}{|B_{j}|}\sum_{i\in B_{j}}X_{i}.$
+  5. 5.
+    Return
+    $\text{empmed}(\bar{X}_{1},\ldots,\bar{X}_{K}).$
+Algorithm 6 Median-of-Means (MoM) estimator.
+Note that the empirical median $\text{empmed}(x_{1},\ldots,x_{k})$ is defined as $x_{i}$ where $i$ is such that:
+|$|\{j\in[k]:x_{j}\leq x_{i}\}|\geq\frac{k}{2}\quad\textrm{ and }\quad|\{j\in[k]:x_{j}\geq x_{i}\}|\geq\frac{k}{2}.$
+If multiple indices satisfy this condition, we take the smallest one.
+C.2 Minsker–Ndaoud (MN)
+The MN estimator (Minsker & Ndaoud 2021) is a weighted average of means computed in blocks, with weights inversely proportional to the blockwise variances.
+The idea is that blocks with higher variance, more likely to contain outliers, are down-weighted in the final estimate.
+The procedure is described in Algorithm 7.
+  1. 1.
+    Input: i.i.d. samples $X_{1},\ldots,X_{n}$ with mean $\mu$ and variance $\sigma^{2}$, confidence parameter $\delta\in(0,1)$, power parameter $a\in\mathbb{N}^{*}$.
+  2. 2.
+    Set $K=\lceil 8\log(1/\delta)\rceil$.
+  3. 3.
+    Partition $[n]=\{1,\ldots,n\}$ into $K$ blocks $B_{1},\ldots,B_{K}$, with each $B_{k}$ of size $\lfloor n/K\rfloor\leq|B_{k}|\leq\lfloor n/K\rfloor+1$.
+  4. 4.
+    For $j\in[K]$, compute block empirical mean
+    $\bar{X}_{j}=\frac{1}{|B_{j}|}\sum_{i\in B_{j}}X_{i}.$
+  5. 5.
+    Compute
+    $\hat{\kappa}=\text{empmed}(\bar{X}_{1},\ldots,\bar{X}_{K})\quad\textrm{ and }\quad d_{i}=(X_{i}-\hat{\kappa})^{2}\quad\textrm{ for }\quad i\in[n].$
+  6. 6.
+    Apply Algorithm 6 to $(d_{i})_{i=1}^{n}$ to obtain $\tilde{\sigma}^{2}$, a MoM-based variance estimate.
+  7. 7.
+    For $j\in[K]$:
+    1. (a)
+      Compute block standard deviation
+      $\hat{\sigma}_{j}^{2}=\frac{1}{|B_{j}|}\sum_{i\in B_{j}}(X_{i}-\bar{X}_{j})^{2}.$
+    2. (b)
+      Compute the weight:
+      $w_{j}=\frac{1}{(\hat{\sigma}_{j}^{2}+\tilde{\sigma}^{2})^{a/2}}.$
+  8. 8.
+    Return
+    $\frac{\sum_{j=1}^{K}w_{j}\bar{X}_{j}}{\sum_{k=1}^{K}w_{k}}.$
+Algorithm 7 Minsker–Ndaoud (MN) estimator.
+The regularization term $\tilde{\sigma}^{2}$ stabilizes the weights and prevents small variances from causing numerical instability.
+The power parameter $a$ controls the sensitivity of the weights to the block variances.
+In our experiments, we fix $a=2$.
+We refer to Theorem 3.1 of Minsker & Ndaoud (2021) for theoretical guarantees of this estimator similar to those recalled in Theorem 5.1 for MoM, but with different constants.
+Contrarily to MoM, MN is asymptotically efficient: its variance is asymptotically equal to that of the sample mean.
+C.3 Lee–Valiant (LV)
+The Lee–Valiant estimator (Lee & Valiant 2022) achieves optimal constants in sub-Gaussian concentration bounds, i.e.
+it improves the constants obtained for MoM in Theorem 5.1.
+It involves first a MoM estimate, and then computes the final estimate as the MoM estimate plus a weighted sum of the data centered with the MoM estimate, where weights depend on the distance to the MoM estimate.
+The description is in Algorithm 8.
+Note that we use $K=\lceil 8\log(1/\delta)\rceil$ blocks, as in MoM and MN, but Lee & Valiant (2022) choose $K=\log(1/\delta)$, assuming this is an integer.
+  1. 1.
+    Input: i.i.d. samples $X_{1},\ldots,X_{n}$ with mean $\mu$ and variance $\sigma^{2}$, confidence parameter $\delta\in(0,1)$.
+  2. 2.
+    Set $K=\lceil 8\log(1/\delta)\rceil$.
+  3. 3.
+    Partition $[n]=\{1,\ldots,n\}$ into $K$ blocks $B_{1},\ldots,B_{K}$, with each $B_{k}$ of size $\lfloor n/K\rfloor\leq|B_{k}|\leq\lfloor n/K\rfloor+1$.
+  4. 4.
+    For $j\in[K]$, compute block empirical mean
+    $\bar{X}_{j}=\frac{1}{|B_{j}|}\sum_{i\in B_{j}}X_{i}.$
+  5. 5.
+    Compute $\hat{\kappa}=\text{empmed}(\bar{X}_{1},\ldots,\bar{X}_{K})$.
+  6. 6.
+    Find the solution $\alpha$ to the equation:
+    $\sum_{i=1}^{n}\min\left(1,\alpha(X_{i}-\hat{\kappa})^{2}\right)=\frac{\log(1/\delta)}{3}.$
+  7. 7.
+    Compute the correction term:
+    $\hat{\Delta}_{n}=\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\hat{\kappa})\left(1-\min\left(1,\alpha(X_{i}-\hat{\kappa})^{2}\right)\right).$
+  8. 8.
+    Return
+    $\hat{\kappa}+\hat{\Delta}_{n}.$
+Algorithm 8 Lee–Valiant (LV) estimator.